### v1

# Sharp non-uniqueness for the Navier–Stokes equations in scaling critical spaces

Abstract.
It is known that uniqueness of mild solutions to the incompressible Navier–Stokes equations holds in the critical class $C([0,T);L^{n}(\mathbb{R}^{n}))$ for $n\geqslant 3$. In this paper, we prove that this result is sharp in the sense that uniqueness fails if $L^{n}(\mathbb{R}^{n})$ is replaced by some scaling critical spaces that are even slightly larger. We achieve this through a complete classification for every pair $(p,q)$ of whether uniqueness of mild solutions in the critical Besov class $C([0,T);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))$ holds or not. Our non-uniqueness mechanism produces infinitely many global solutions emanating even from zero initial state, whose large-time asymptotics are governed by non-trivial stationary flow. To the best of our knowledge, such non-unique solutions provide the first examples of non-dissipative unforced Navier–Stokes flow with critical regularity.
Key words and phrases:
Navier–Sotkes equations, non-uniqueness, scaling critical Besov spaces
2020 Mathematics Subject Classification:
35Q30, 35A02, 76D03
1. Introduction
Let us consider the initial value problem for the incompressible Navier–Stokes equations on $\mathbb{R}^{n}$ with $n\geqslant 2$:
[…]
where $u=u(t,x):[0,\infty)\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ is the unknown velocity field of the fluid, whereas $u_{0}=u_{0}(x):\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a given initial datum satisfying $\operatorname{div}u_{0}=0$. We denote by $\mathbb{P}={\rm Id}+\nabla(-\Delta)^{-1}\operatorname{div}$ the Helmholtz projection.
In dimensions $n\geqslant 3$, Lions–Masmoudi [Lio-Mas] and Furioli–Rieusset–Terraneo [Fur-Rie-Ter-2000] proved that mild solutions to (1.1) are unique in the critical space $C([0,T);L^{n}(\mathbb{R}^{n}))$. The aim of this paper is to prove the sharpness of the choice of $L^{n}(\mathbb{R}^{n})$ in the sense that the uniqueness fails when one replaces $L^{n}(\mathbb{R}^{n})$ by a scaling critical space that is even slightly larger; for instance, relaxing $L^{n}(\mathbb{R}^{n})=\dot{F}_{n,2}^{0}(\mathbb{R}^{n})$ to a slightly larger Lizorkin–Triebel space $\dot{F}_{n,q}^{0}(\mathbb{R}^{n})$ with any $q>2$ leads to non-uniqueness of the mild solutions in $C([0,T);\dot{F}_{n,q}^{0}(\mathbb{R}^{n}))$. To achieve this, we show that the necessary and sufficient condition for the uniqueness of mild solutions to (1.1) in the scaling critical Besov spaces is given by $1\leqslant p<n$ with $1\leqslant q\leqslant\infty$, or $p=n$ with $1\leqslant q\leqslant 2$; see Figure 1 below. Moreover, for the non-unique case $p=n$ with $2<q\leqslant\infty$, or $n<p\leqslant\infty$ with $1\leqslant q\leqslant\infty$, we prove that the zero initial datum $u_{0}=0$ generates infinitely many small global solutions $u\in C([0,\infty);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))$, which asymptotically converge to non-trivial small steady states $U\in\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ solving the unforced stationary Navier–Stokes equations:
$\begin{cases}-\Delta U+\mathbb{P}\operatorname{div}(U\otimes U)=0,\qquad&x\in\mathbb{R}^{n},\\ \operatorname{div}U=0,\qquad&x\in\mathbb{R}^{n}.\end{cases}$ (1.2)
1.1. Known results related to our work
Before we state our main results precisely, we recall previous studies related to our work and mention the place of our work. By the Duhamel principle, we may formally rewrite (1.1) as the integral equation
$u(t)=e^{t\Delta}u_{0}-\int_{0}^{t}e^{(t-\tau)\Delta}\mathbb{P}\operatorname{div}(u(\tau)\otimes u(\tau))d\tau,$ (1.3)
A solution of (1.3) is called a mild solution. It is natural to construct mild solutions in scaling critical spaces. Here, let us recall the notion of scaling criticality. If $u$ is a mild solution, then $u_{\lambda}(t,x)=\lambda u(\lambda^{2}t,\lambda x)$ also solves (1.1) for all $\lambda>0$. A Banach space $X$ is called a scaling critical space if ${\left\|u_{\lambda}(0,\cdot)\right\|}_{X}={\left\|u(0,\cdot)\right\|}_{X}$ for all $\lambda>0$. Typical examples of scaling critical spaces are as follows
$\dot{H}^{\frac{n}{2}-1}(\mathbb{R}^{n})\hookrightarrow L^{n}(\mathbb{R}^{n})\hookrightarrow L^{n,\infty}(\mathbb{R}^{n}),$ (1.4)
$\dot{B}_{p,q}^{\frac{n}{p}-1}(\mathbb{R}^{n})\hookrightarrow{\rm BMO}^{-1}(\mathbb{R}^{n})\qquad(1\leqslant p<\infty,\ 1\leqslant q\leqslant\infty).$ (1.5)
This line of research originates from the pioneering work of Fujita–Kato [Fuj-Kato-64-ARMA], where they proved well-posedness in $\dot{H}^{n/2-1}(\mathbb{R}^{n})$, and was extended to $L^{n}(\mathbb{R}^{n})$ by Kato [Kato-84] and Giga–Miyakawa [Gig-Miy-85]. Moreover, Kozono–Yamazaki [Koz-Yam-94], Cannone–Planchon [Can-Pla-96], and Planchon [Pla-98] proved well-posedness in $\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ for $1\leqslant p<\infty$ and $1\leqslant q\leqslant\infty$. In the endpoint case $p=\infty$, Koch–Tataru [Koh-Tat-01] established well-posedness for small data in ${\rm BMO}^{-1}(\mathbb{R}^{n})$, however it was shown by [Bou-Pav-08, Yon-10, Wan-15] that (1.1) is ill-posed in $\dot{B}_{\infty,q}^{-1}(\mathbb{R}^{n})$ ($1\leqslant q\leqslant\infty$). We refer to [Kozono-Shimizu-2018, Takeuchi-2025-SIMA, Yam-2000, arXiv:2509.21272v2] for the forced Navier–Stokes flow in critical spaces.
Next, we address the issue of uniqueness for the Navier–Stokes equations. This topic was brought to the forefront by Leray’s pioneering work [Leray-1934-Acta], which produced weak solutions for any $L^{2}$ initial datum but did not settle their uniqueness. It is well-known that the Leray–Hopf weak solutions are unique if they belong to the Ladyzhenskaya–Prodi–Serrin class $L^{p}(0,T;L^{q}(\mathbb{R}^{n}))$ with $n/q+2/p=1$; see [Serrin-1963, Masuda-1984, Escauriaza-Seregin-Sverak-2003, Kozono-Sohr-1996] for more detail. For the mild solutions, Lions–Masmoudi [Lio-Mas] and Furioli–Rieusset–Terraneo [Fur-Rie-Ter-2000] proved the uniqueness in the scaling critical class $C([0,T);L^{n}(\mathbb{R}^{n}))$ with $n\geqslant 3$, which corresponds to an endpoint $(p,q)=(\infty,n)$ of the Ladyzhenskaya–Prodi–Serrin class. Note that [Fur-Rie-Ter-2000] considered not only in $L^{n}(\mathbb{R}^{n})$ but also other critical spaces such as Morrey spaces. See also [Yam-2000] for the uniqueness of small mild solutions in $L^{\infty}(0,T;L^{n,\infty}(\mathbb{R}^{n}))$ and [Miu-2005] for the uniqueness in $C([0,T);bmo^{-1}(\mathbb{R}^{n}))\cap L^{\infty}_{\rm loc}((0,T);L^{\infty}(\mathbb{R}^{n}))$. The recent progress of this direction is given by Zhan [arXiv:2402.01174v2], where he considered the uniqueness of mild solutions in $C([0,T);\widetilde{L}^{n,\infty}(\mathbb{R}^{n}))$ with a new space $\widetilde{L}^{n,\infty}(\mathbb{R}^{n})$ which is larger than $L^{n}(\mathbb{R}^{n})$. In terms of the Besov spaces, the following proposition holds.
Proposition 1.1.
Let $n\geqslant 3$. Let $p$ and $q$ satisfy either of the following.
  • (U1)
    $1\leqslant p<n$ and $1\leqslant q\leqslant\infty$,
  • (U2)
    $p=n$ and $1\leqslant q\leqslant 2$.
Then, the uniqueness of mild solutions to (1.1) holds in $C([0,T);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))$, provided a smallness condition on solutions only when $q=\infty$.
The proof of Proposition 1.1 follows immediately by adopting the argument in [Fur-Rie-Ter-2000, Fuj-AIPHC, Iwabuchi-Okazaki-2025] in the case of (U1) and by the result of [Fur-Rie-Ter-2000] with the continuous embedding $\dot{B}_{n,q}^{0}(\mathbb{R}^{n})\hookrightarrow L^{n}(\mathbb{R}^{n})$ for the case of (U2). Note that Proposition 1.1 still holds if one replaces Besov spaces by Lizorkin–Triebel spaces as the continuous embedding $\dot{F}_{p,q}^{n/p-1}(\mathbb{R}^{n})\hookrightarrow L^{n}(\mathbb{R}^{n})$ for all $(p,q)$ satisfying (U1) or (U2).
Finally, we briefly review works on non-uniqueness. One major direction originates from applying Nash’s convex integration to the 3D Euler equations on the torus (see, [DeLellis-Szekelyhidi-09, Isett-18, Daneri-Szekelyhidi-2017-ARMA, Lellis-Szekelyhidi-2013-Invent, Buckmaster-Lellis-Isett-Szekelyhidi-2015-AnnMath] for instance). Using an iterative scheme based on the Euler–Reynolds system, these works constructed Euler flows violating the energy conservation law and made substantial progress toward the Onsager conjecture. Although incorporating the viscous term was long thought to be fundamentally difficult, Buckmaster–Vicol [Buc-Vic-19] introduced the concept of intermittent convex integration and proved that uniqueness of weak solutions to the Navier–Stokes equations on $\mathbb{T}^{3}$ breaks down in the supercritical class $C([0,T);H^{\beta}(\mathbb{T}^{3}))$ for some $0<\beta\ll 1/2$. Cheskidov–Luo [Che-Luo-21] showed that an endpoint Ladyzhenskaya–Prodi–Serrin class $L^{2}(0,T;L^{\infty}(\mathbb{T}^{n}))$, is optimal with respect to the time integrability exponent: they constructed non-unique solutions in $L^{p}(0,T;L^{\infty}(\mathbb{T}^{3}))$ for $p<2$. They also considered another endpoint phenomenon in two dimensions and proved non-uniqueness in $C([0,T);L^{p}(\mathbb{T}^{2}))$ for $p<2$ (see [Che-Luo-22-AnnPDE]). Furthermore, Cheskidov–Zeng–Zhang [arXiv:2503.05692] proved that every initial datum in $H^{1/2}(\mathbb{T}^{3})$ generates infinitely many Navier–Stokes flows whose kinetic energy is continuous and monotonically decreasing in time. Remarkably, Coiculescu–Palasek [Coi-Pal-25] showed that even in ${\rm BMO}^{-1}(\mathbb{T}^{3})$, where well-posedness for small data is known, a certain large initial datum generates two smooth solutions. Their argument builds on ideas developed by Palasek [Pal-25-IMRN] for the dyadic Navier–Stokes model. Pushing this direction further, Cheskidov–Dai–Palasek [arXiv:2511.09556] observed that non-uniqueness in the form of instantaneous Type I blow-up may be generated from arbitrary $C^{\infty}$ solenoidal initial data. Another strategy to prove non-uniqueness, developed by Jia–Sverák [Jia-Sve-Invent, Jia-Sve-JFA], is to exploit the instability of certain self-similar solutions. A notable result in this direction is due to Albritton–Brué–Colombo [Alb-Bru-Col-AnnMath], who proved non-uniqueness of Leray–Hopf weak solutions to the forced Navier–Stokes equations. See Aoki–Maekawa [Aok-Mae] for the corresponding result on a two-dimensional half space. More recently, Hou–Wang–Yang [2509.25116v1] obtained an analogous statement in the unforced case $f\equiv 0$ by a computer-assisted proof.
1.2. Main results
The purpose of this paper is to show non-uniqueness of mild solutions to (1.1) in $C([0,T);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))$ for all $(p,q)$ satisfying neither (U1) nor (U2). Now, the precise statement of our main theorem reads as follows.
Theorem 1.2.
Let $n\geqslant 3$. Let $p$ and $q$ satisfy either of the following.
  • (N1)
    $p=n$ and $2<q\leqslant\infty$
  • (N2)
    $n<p\leqslant\infty$ and $1\leqslant q\leqslant\infty$.
Then, there exist constants $\varepsilon_{0}=\varepsilon_{0}(n,p,q)>0$ and $\eta=\eta(n,p,q)\geqslant 0$ such that for any $0<\varepsilon\leqslant\varepsilon_{0}$ and solenoidal initial datum $u_{0}\in\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ with ${\left\|u_{0}\right\|}_{\dot{B}_{p,q}^{n/p-1}}\leqslant\eta$, (1.1) possesses a global mild solution $u_{\varepsilon}\in C([0,\infty);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))$ satisfying
$\lim_{t\to\infty}{\left\|u_{\varepsilon}(t)-U_{\varepsilon}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}=0,$ (1.6)
where $U_{\varepsilon}\in\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ solves (1.2) and admits the estimate
$C^{-1}\varepsilon\leqslant{\left\|U_{\varepsilon}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\leqslant C\varepsilon$ (1.7)
with some positive constant $C=C(n,p,q)$ independent of $u_{0}$ and $\varepsilon$.
Remark 1.3.
Let us mention some remarks on our result.
  1. 1.
    Chemin [Che-92] and Planchon [Pla-98] established well-posedness for (1.1) in critical Besov spaces for all $1\leqslant p<\infty$ and $1\leqslant q\leqslant\infty$. This does not contradict our result, since their solution space is based on
    $C([0,T);\dot{B}_{p,q}^{\frac{n}{p}-1}(\mathbb{R}^{n}))\cap\widetilde{L^{1}}(0,T;\dot{B}_{p,q}^{\frac{n}{p}+1}(\mathbb{R}^{n})).$ (1.8)
    whereas our non-unique solutions do not satisfy the auxiliary dissipative condition $\widetilde{L^{1}}(0,T;\dot{B}_{p,q}^{n/p+1}(\mathbb{R}^{n}))$.
  2. 2.
    The large-time behavior of global solutions to (1.1) with small initial data has been studied extensively; see [Pla-98, Nakasato-25-JDE, Gallagher-Iftimie-Planchon-2002, Gallagher-Iftimie-Planchon-2003, Takeuchi-2025, Watanabe-2025, Benameu-2015-JMAA] for instance. However, all of these results exploit the viscous dissipation and address only solutions that decay and converge asymptotically to zero. In contrast, the non-decaying global solutions $u_{\varepsilon}$ constructed in this paper are the first examples [^1] In the framework of weak solutions, the existence of non-dissipative solutions is widely known; see [Buc-Vic-19]. of non-dissipative unforced Navier–Stokes flows in critical spaces where the Stokes semigroup is strongly continuous. [^2] If this is not the case, solutions may not decay even at the linear level.
  3. 3.
    For any $(p,q)$ satisfying (N1) or (N2), the continuous embeddings
    $\dot{B}_{n,2+\delta}^{0}(\mathbb{R}^{n})\hookrightarrow\dot{F}_{n,2+\delta}^{0}(\mathbb{R}^{n})\hookrightarrow\dot{F}_{p,q}^{\frac{n}{p}-1}(\mathbb{R}^{n})$ (1.9)
    hold with some $\delta=\delta(p,q)\in(0,1)$. Thus, non-uniqueness for $u_{0}=0$ remains valid after replacing the Besov spaces by the corresponding Lizorkin–Triebel spaces. In particular, we see that enlarging $L^{n}(\mathbb{R}^{n})=\dot{F}_{n,2}^{0}(\mathbb{R}^{n})$ in [Lio-Mas, Fur-Rie-Ter-2000] even slightly to $\dot{F}_{n,q}^{0}(\mathbb{R}^{n})$ with $q>2$ instantly leads to non-uniqueness. Moreover, we obtain the breakdown of unconditional uniqueness for small solutions in $L^{\infty}(0,T;{\rm BMO}^{-1}(\mathbb{R}^{n}))$ since ${\rm BMO}^{-1}(\mathbb{R}^{n})=\dot{F}_{\infty,2}^{-1}(\mathbb{R}^{n})$. This is different from the phenomenon in [Coi-Pal-25], which concerns non-uniqueness of smooth solutions for large data in ${\rm BMO}^{-1}(\mathbb{T}^{3})$. See Figure 1.
  4. 4.
    On the results for the stationary Navier–Stokes equations in critical spaces, Chen [Che-93] proved existence and uniqueness of small solutions in $L^{n}(\mathbb{R}^{n})$, and Kaneko–Kozono–Shimizu [Kan-Koz-Shi-19] established unique solvability in $\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ for small external forces under condition (U1). Thus, for $(p,q)$ satisfying (U1) or (U2), small steady solutions are unique in $\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$. Consequently, Theorem 1.2 asserts sharp non-uniqueness of small solutions in $\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$. For the ill-posedness in the sense of discontinuity of the solution map, see [Tsu-19-ARMA, Li-Yu-Zhu-25, Tan-Tsu-Zha-25, Fuj-24].
  5. 5.
    Concerning non-uniqueness for steady flows, Luo [Luo-ARMA-2019] constructed nontrivial solutions belonging to $L^{2}(\mathbb{T}^{n})$ with $n\geqslant 4$. See [Rieusset-25-JFA] for two-dimensional case in $H^{-1}(\mathbb{T}^{2})$. Note that both spaces are supercritical and the three-dimensional case has been open. Our result appears to be the first to establish non-uniqueness of steady solutions in scaling critical spaces including three-dimensional case. Meanwhile, our method cannot be implemented in two dimensions. This obstruction essentially reduces to the well-known difficulty of the solvability for the two-dimensional stationary Navier–Stokes equations on $\mathbb{R}^{2}$.
  6. 6.
    As will be explained later, our construction of solutions relies on the structure provided by the classical Nash lemma, whereas we do not employ any of the periodic building blocks commonly used in convex-integration approaches to fluid equations, such as Beltrami or Mikado flows.
[…]
Figure 1. Let $X_{p,q}$ be either $\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ or $\dot{F}_{p,q}^{n/p-1}(\mathbb{R}^{n})$. The blue region indicates the range of $(1/q,1/p)$ for which mild solutions to (1.1) are unique in $C([0,T);X_{p,q})$ and small stationary solutions in $X_{p,q}$ are unique, whereas the red region indicates the range where uniqueness fails for both non-stationary and stationary solutions. Note that the continuous embedding $X_{p_{1},q_{1}}\hookrightarrow X_{p_{2},q_{2}}$ for $1\leqslant p_{1}\leqslant p_{2}\leqslant\infty$ and $1\leqslant q_{1}\leqslant q_{2}\leqslant\infty$ implies that the underlying function space becomes larger as a point in the figure moves leftward or downward.
1.3. Outline of the proof
Let us explain the strategy of the proof of Theorem 1.2. The strategy is that we first build a dominant part of the solution and then solve the remaining perturbative problem by a fixed point argument in a function space where well-posedness is available. More precisely, we construct an almost frequency-localized bulding blocks $\{V_{j}\}_{j\geqslant 0}$ of stationary flows, each essentially supported on the $j$th frequency shell, and design the principal profile
$V=\sum_{j=0}^{\infty}V_{j}\neq 0\qquad{\rm in}\quad\dot{B}_{p,q}^{\frac{n}{p}-1}(\mathbb{R}^{n})$ (1.10)
so that the associated residual forcing
$F:=-\Delta V+\mathbb{P}\operatorname{div}(V\otimes V),$ (1.11)
belongs to well-posedness class $\dot{B}_{r,1}^{n/r-3}(\mathbb{R}^{n})$ ($r<n$) and admits the following smallness hierarchy:
$\|F\|_{\dot{B}_{r,1}^{\frac{n}{r}-3}}\ll\|V\|_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\ll 1.$ (1.12)
Once (1.12) is achieved, the perturbation [^3] The perturbed system is given in (5.1) below. $W$ driven by $F$ can be constructed in $\dot{B}_{r,1}^{n/r-1}(\mathbb{R}^{n})$ by well-posedness theory, and we see that
${\left\|W\right\|}_{\dot{B}_{r,1}^{n/r-1}}\lesssim{\left\|F\right\|}_{\dot{B}_{r,1}^{n/r-3}}\ll{\left\|V\right\|}_{\dot{B}_{p,q}^{n/p-1}}.$ (1.13)
Hence, the desired steady state is obtained in the form $U=V+W$. Note that ${\left\|U\right\|}_{\dot{B}_{p,q}^{n/p-1}}\sim{\left\|V\right\|}_{\dot{B}_{p,q}^{n/p-1}}\neq 0$ holds. Furthermore, to construct a non-stationary solution $u$ whose asymptotic profile is $U$, we decompose $u=U+w$. Then, we may follow the standard stability argument to construct a small dissipative perturbed flow $w\in C([0,\infty);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))\cap\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{n/p-1+2/\theta}(\mathbb{R}^{n}))$ with some $2<\theta<\infty$. Then, the auxilliary dissipative space enables us to show $\|w(t)\|_{\dot{B}_{p,q}^{n/p-1}}\to 0$ as $t\to\infty$.
We now explain how to construct the leading profile $V$ satisfying (1.12). A natural first attempt, motivated by existing works such as [Coi-Pal-25], would be to choose each $V_{j}$ as a superposition of plane waves concentrated on a single highly lacunary frequency scale $N_{j}=A^{\gamma b^{j}}$ ($A\gg 1$), for instance, $V_{j}(x)=\sum_{k\in\mathcal{K}}f_{k,j}(x)\cos(N_{j}a_{k}\cdot x)k$, where $\mathcal{K}\subset\mathbb{R}^{n}$ is a finite set, $f_{k,j}$ are smooth amplitudes, and $a_{k}$ are constant vectors with $a_{k}\cdot k=0$. However, such a single-scale construction is too rigid to capture the delicate dependence on the Besov summability index $q$, which is essential for observing the subtle non-uniqueness phenomena addressed in (N2). Instead, we distribute the $j$th building block over many internal scales. To this end, for $\lambda\gg 1$ (which corresponds to $\varepsilon^{-1/\alpha}$ with some $\alpha$) and a fixed large number $\mu\in\mathbb{N}$, we introduce a tetration-like sequence [^4] This is defined by $\lambda_{0}=\lambda$ and $\lambda_{j}=\lambda^{2^{\lambda_{j-1}^{4\mu}}}$; see (4.9). $\{\lambda_{j}\}_{j\geqslant 0}$ and define the $j$th level frequency index set by
$\Lambda_{j}\coloneqq\Bigl\{\ell\in\mathbb{N}\ ;\ \mu\log_{\lambda}\lambda_{j-1}\leqslant\ell\leqslant\log_{\lambda}\lambda_{j}\Bigr\},$
and take $V_{j}$ in the schematic form
$V_{j}(x)=\sum_{k\in\mathcal{K}}f_{k,j}(x)\sum_{\ell\in\Lambda_{j}}\frac{1}{\sqrt{h_{j}\ell}}\cos\left(\lambda^{\ell}a_{k}\cdot x\right)k,\qquad h_{j}:=\sum_{\ell\in\Lambda_{j}}\frac{1}{\ell}.$ (1.14)
Here, the choice of $1/\sqrt{h_{j}\ell}$ is inspired by [Tsu-19-ARMA]. With this multi-scale design, the oscillation calculus in Section 3 below yields
${\left\|V_{j}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\sim\sum_{k\in\mathcal{K}}\frac{{\left\|f_{k,j}\right\|}_{L^{p}}}{\sqrt{h_{j}}}\left(\sum_{j\in\Lambda_{j}}\frac{\lambda^{(\frac{n}{p}-1)q\ell}}{\ell^{\frac{q}{2}}}\right)^{\frac{1}{q}},$ (1.15)
which may shrink as $j\to\infty$ if and only if $(p,q)$ satisfies (N1) or (N2). Thus, the series (1.10) converges and we may define $V$. For the analysis of $F$, we decompose $F$ as
$F=\sum_{j=1}^{\infty}\Bigl(-\Delta V_{j-1}+\mathbb{P}\operatorname{div}(V_{j}\otimes V_{j})\Bigr)+\sum_{j_{1}\neq j_{2}}\mathbb{P}\operatorname{div}(V_{j_{1}}\otimes V_{j_{2}})=:F_{1}+F_{2}.$ (1.16)
The cross-shell interaction term $F_{2}$ is favorable, as it consists of products of flows concentrated on distinct frequency shells. By contrast, $F_{1}$ exhibits a more delicate structure. Indeed, within the nonlinear interaction $\operatorname{div}(V_{j}\otimes V_{j})$, resonances among modes inside the frequency shell $\Lambda_{j}$ generate an inverse-cascade transfer toward low frequencies, which exhibit the worst regularity. By selecting the directional set $\mathcal{K}$ and the amplitudes $f_{k,j}$ appropriately through Nash’s geometric lemma, one can arrange for this low-frequency contribution to cancel $\Delta V_{j-1}$. Consequently, the forcing term $F$ attains improved regularity and smaller magnitude relative to $V$.
1.4. Organization of this paper
This paper is organized as follows. In Section 2, we summarize the notations used in this paper and elementary facts on Besov and Chemin–Lerner spaces. In Section 3, we provide key Besov estimates for oscillating functions, which play a key role in constructing the leading part of the non-unique stationary solutions. In Section 4, we construct the principal part of the non-trivial steady state and calculate Besov norms of it and the perturbative force. In Section 5, we make use of special flow constructed in Section 4 to complete the proof of Theorem 1.2.
2. Preliminaries
Throughout this paper, we denote by $C$ and $c$ the constants, which may differ in each line. In particular, C=C(,,)C=C(*,...,*) denotes the constant which depends only on the quantities appearing in parentheses. Let $\mathbb{N}$ be the set of all positive integers and put $\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$. For $\mathbb{M}\subset\mathbb{Z}$, we define $2^{\mathbb{M}}:=\{2^{m}\ ;\ m\in\mathbb{M}\}$. The identity matrix is denoted by $\rm Id$}. For a metric space $(X,d)$, $x_{0}\in X$, and a positive number $r>0$, set $B_{X}(x_{0},r):=\{x\in X\ ;\ d(x,x_{0})\leqslant r\}$. Let $\mathscr{S}(\mathbb{R}^{n})$ be the set of all Schwartz functions on $\mathbb{R}^{n}$, and we denote by $\mathscr{S}^{\prime}(\mathbb{R}^{n})$ the set of all tempered distributions on $\mathbb{R}^{n}$.
Next, we recall the definitions of Besov and Chemin–Lerner spaces. Let $\varphi_{0}\in\mathscr{S}(\mathbb{R}^{n})$ satisfy
$\operatorname{supp}\widehat{\varphi_{0}}\subset\left\{\xi\in\mathbb{R}^{n}\ ;\ 2^{-1}\leqslant|\xi|\leqslant 2\right\},\quad 0\leqslant\widehat{\varphi_{0}}(\xi)\leqslant 1,$ (2.1)
and
$\sum_{j\in\mathbb{Z}}\widehat{\varphi_{j}}(\xi)=1\qquad{\rm for\ all\ }\xi\in\mathbb{R}^{n}\setminus\{0\},$ (2.2)
where we have set $\varphi_{j}(x):=2^{nj}\varphi_{0}(2^{j}x)$. Using this $\{\varphi_{j}\}_{j\in\mathbb{Z}}$, we may define the Littlewood–Paley decomposition as
$\Delta_{j}f:=\mathscr{F}^{-1}\left[\widehat{\varphi_{j}}(\xi)\widehat{f}(\xi)\right],\qquad f\in\mathscr{S}^{\prime}(\mathbb{R}^{n}),\quad j\in\mathbb{Z}.$ (2.3)
For $1\leqslant p,q\leqslant\infty$ and $s\in\mathbb{R}$, the Besov space $\dot{B}_{p,q}^{s}(\mathbb{R}^{n})$ is defined as
$\dot{B}_{p,q}^{s}(\mathbb{R}^{n}):={}$ $\left\{f\in\mathscr{S}^{\prime}(\mathbb{R}^{n})/\mathscr{P}(\mathbb{R}^{n})\ ;\ {\left\|f\right\|}_{\dot{B}_{p,q}^{s}}<\infty\right\},$ (2.4)
${\left\|f\right\|}_{\dot{B}_{p,q}^{s}}:={}$ ${\left\|\left\{2^{sj}{\left\|\Delta_{j}f\right\|}_{L^{p}}\right\}_{j\in\mathbb{Z}}\right\|}_{\ell^{q}(\mathbb{Z})},$ (2.5)
where $\mathscr{P}(\mathbb{R}^{n})$ is the set of all polynomials on $\mathbb{R}^{n}$. It is well-known that if $s<n/p$ or $(s,q)=(n/p,1)$, then it holds
$\dot{B}_{p,q}^{s}(\mathbb{R}^{n})\sim\left\{f\in\mathscr{S}^{\prime}(\mathbb{R}^{n})\ ;\ {\left\|f\right\|}_{\dot{B}_{p,q}^{s}}<\infty,\quad f=\sum_{j\in\mathbb{Z}}\Delta_{j}f\quad{\rm in\ }\mathscr{S}^{\prime}(\mathbb{R}^{n})\right\}.$ (2.6)
Let us prepare a bilinear estimate in Besov spaces.
Lemma 2.1.
Let $n\geqslant 3$. Then, for any $1\leqslant p,q,r\leqslant\infty$ satisfying
$\frac{n}{p}+\frac{n}{r}>2,\qquad 0\leqslant\frac{n}{r}-\frac{n}{p}<1,$ (2.7)
there exists a positive constant $C=C(n,p,q,r)$ such that
${\left\|(-\Delta)^{-1}\mathbb{P}\operatorname{div}(V\otimes W)\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\leqslant C{\left\|V\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}{\left\|W\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}$ (2.8)
for all $V\in\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ and $W\in\dot{B}_{r,1}^{n/r-1}(\mathbb{R}^{n})$.
We omit the precise proof of Lemma 2.1 as it is immediately obtained by [Abidi-Paicu-2007]*Corollary 2.5. See [Bah-Che-Dan-11, Saw-18] for more information on Besov spaces.
Next, we recall the Chemin–Lerner spaces which were first introduced in [Che-Ler-95]:
$\widetilde{L^{\theta}}(I;\dot{B}_{p,q}^{s}(\mathbb{R}^{n})):={}$ $\left\{F:I\to\mathscr{S}^{\prime}(\mathbb{R}^{n})/\mathscr{P}(\mathbb{R}^{n})\ ;\ {\left\|F\right\|}_{\widetilde{L^{\theta}}(I;\dot{B}_{p,q}^{s})}<\infty\right\},$ (2.9)
${\left\|F\right\|}_{\widetilde{L^{\theta}}(I;\dot{B}_{p,q}^{s})}:={}$ ${\left\|\left\{2^{sj}{\left\|\Delta_{j}F\right\|}_{L^{\theta}(I;L^{p})}\right\}_{j\in\mathbb{Z}}\right\|}_{\ell^{q}(\mathbb{Z})}$ (2.10)
for $1\leqslant p,q,\theta\leqslant\infty$, $s\in\mathbb{R}$, and an interval $I\subset\mathbb{R}$. We also use the notation
$\widetilde{C}(I;\dot{B}_{p,q}^{s}(\mathbb{R}^{n})):=C(I;\dot{B}_{p,q}^{s}(\mathbb{R}^{n}))\cap\widetilde{L^{\infty}}(I;\dot{B}_{p,q}^{s}(\mathbb{R}^{n})).$ (2.11)
The Chemin–Lerner spaces are well-known to be particularly convenient when exploiting the maximal regularity of the heat kernel, and we take advantage of this framework through the following lemma, which will be used in our stability analysis for non-stationary flows around the steady state.
Lemma 2.2 ([Fuj-AIPHC]).
Let $n\geqslant 2$. Let $1\leqslant p,q,\theta\leqslant\infty$ satisfy
$2<\theta<\infty,\qquad\min\left\{n,\frac{2n}{p}\right\}-2+\frac{2}{\theta}>0.$ (2.12)
Then, there exists a positive constant $C=C(n,p,q,\theta)$ such that
${\left\|\int_{t_{0}}^{t}e^{(t-\tau)\Delta}\mathbb{P}\operatorname{div}(v(\tau)\otimes w(\tau))d\tau\right\|}_{\widetilde{L^{\infty}}(I;\dot{B}_{p,q}^{\frac{n}{p}-1})\cap\widetilde{L^{\theta}}(I;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}$ (2.13)
$\quad\leqslant C{\left\|v\right\|}_{\widetilde{L^{\infty}}(I;\dot{B}_{p,q}^{\frac{n}{p}-1})}{\left\|w\right\|}_{\widetilde{L^{\theta}}(I;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}$ (2.14)
for all $I=(t_{0},t_{1})\subset\mathbb{R}$, $v\in\widetilde{L^{\infty}}(I;\dot{B}_{p,q}^{\frac{n}{p}-1}(\mathbb{R}^{n}))$, and $w\in\widetilde{L^{\theta}}(I;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}}(\mathbb{R}^{n}))$.
3. Highly oscillating functions in Besov spaces
In the proof of our main result, we frequently use the sharp Besov estimates for finite superpositions of functions obtained by modulating a smooth amplitude with a plane wave of the form (3.4) below. If each Fourier transform $\widehat{f_{k}}$ of the amplitudes were compactly supported in a low frequency region, then the Fourier support of each $g_{\lambda,\ell,k}$ is concentrated around $\xi\sim\lambda^{\ell}a_{k}$, and we have
${\left\|g_{\lambda,\ell,k}\right\|}_{\dot{B}_{p,q}^{s}}\sim|b_{\ell}|(\lambda^{\ell}|a_{k}|)^{s}{\left\|f_{k}\right\|}_{L^{p}}$ (3.1)
For sufficiently large $\lambda$, these supports are disjoint, which implies
$\|F_{\lambda,L}\|_{\dot{B}^{s}_{p,q}}\sim\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}\sum_{k=1}^{K}|a_{k}|^{s}{\left\|f_{k}\right\|}_{L^{p}}.$ (3.2)
However, in our construction of the non-unique steady Navier–Stokes flow, the amplitudes $f_{1},\dots,f_{K}$ necessarily belong to $C_{c}^{\infty}(\mathbb{R}^{n})$, so their Fourier transforms decay rapidly but are not compactly supported. Therefore, it is necessary to proceed with due care so that the Fourier tails of each $f_{k}$ may be treated as error terms. The following proposition realizes this.
Proposition 3.1.
Let $1\leqslant p,q\leqslant\infty$ and $s\in\mathbb{R}$ satisfy
$s>\frac{n}{p}-n.$ (3.3)
Let $K,L\in\mathbb{N}$. Let $a_{1},\dots,a_{K}\subset\mathbb{R}^{n}\setminus\{0\}$ satisfy $|a_{k}|\in 2^{\mathbb{N}}$ $(k=1,\dots,K)$ and assume that $a_{k}$, $a_{k^{\prime}}$ do not point in the same direction unless $k=k^{\prime}$. For $\lambda\in 2^{\mathbb{N}}$, $b_{1},\dots,b_{L}\in\mathbb{R}$, and $f_{1},\dots,f_{K}\in\mathscr{S}(\mathbb{R}^{n})$, set
$F_{\lambda,L}(x):=\sum_{k=1}^{K}\sum_{\ell=1}^{L}g_{\lambda,\ell,k}(x),\qquad g_{\lambda,\ell,k}(x):=b_{\ell}f_{k}(x)e^{i\lambda^{\ell}a_{k}\cdot x}.$ (3.4)
Then, for any $M\in\mathbb{N}$ with $M>s$, there exists a positive constant $C$, depending only on $n$, $p$, $q$, $s$, $K$, and $M$ such that
${\left\|F_{\lambda,L}\right\|}_{\dot{B}_{p,q}^{s}}\leqslant{}$ $C\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}\sum_{k=1}^{K}|a_{k}|^{s}{\left\|f_{k}\right\|}_{L^{p}}$ (3.5)
$+C\sum_{k=1}^{K}\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s+n(1-\frac{1}{p})-M}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}\cap L^{1}}$ (3.6)
and
${\left\|F_{\lambda,L}\right\|}_{\dot{B}_{p,q}^{s}}\geqslant{}$ $C^{-1}\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}\max_{1\leqslant k\leqslant K}|a_{k}|^{s}{\left\|f_{k}\right\|}_{L^{p}}$ (3.7)
$-C\max_{1\leqslant k\leqslant K}\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s-1}|b_{\ell}|{\left\|\nabla f_{k}\right\|}_{L^{p}}$ (3.8)
$-C\max_{1\leqslant k\leqslant K}\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s+n(1-\frac{1}{p})-M}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}\cap L^{1}}$ (3.9)
for all $L\in\mathbb{N}$ and $\lambda\in 2^{\mathbb{N}}$ with $\lambda\geqslant 2^{100}$.
Remark 3.2.
To obtain (3.6) with $q=1$, we may remove the directional assumption on $a_{1},\dots,a_{K}$. Indeed, applying (3.6) with $K=L=1$ to each $f_{k}(x)e^{i\lambda^{\ell}a_{k}\cdot x}$ gives
${\left\|F_{\lambda,L}\right\|}_{\dot{B}_{p,1}^{s}}\leqslant{}$ $\sum_{k=1}^{K}\sum_{\ell=1}^{L}|b_{\ell}|{\left\|f_{k}(x)e^{i\lambda^{\ell}a_{k}\cdot x}\right\|}_{\dot{B}_{p,1}^{s}}$ (3.10)
$\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s}|b_{\ell}|{\left\|f_{k}\right\|}_{L^{p}}$ (3.11)
$+C\sum_{k=1}^{K}\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s+n(1-\frac{1}{p})-M}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}\cap L^{1}}.$ (3.12)
To prove Proposition 3.1, we introduce operators that localize frequencies in the angular direction. For any $a,a^{\prime}\in\mathbb{R}^{n}\setminus\{0\}$, we define the angle between $a$ and $a^{\prime}$ by
$\vartheta(a,a^{\prime}):=\operatorname{Arccos}\left(\frac{a\cdot a^{\prime}}{|a||a^{\prime}|}\right).$ (3.13)
For $a_{1},\dots,a_{K}$ given in Proposition 3.1, we put
$\vartheta_{*}:=\min_{k\neq k^{\prime}}\vartheta\left(a_{k},a_{k}^{\prime}\right)\in(0,\pi].$ (3.14)
Here, we set $\vartheta_{*}:=1$ if $K=1$. Let $\widetilde{\phi}_{1},\dots,\widetilde{\phi}_{K}\in C_{c}^{\infty}(\mathbb{S}^{n-1})$ satisfy $\operatorname{supp}\widetilde{\phi}_{k}\subset\{\sigma\in\mathbb{S}^{n-1}\ ;\ \vartheta(\sigma,a_{k})<\vartheta_{*}/3\}$ and $\widetilde{\phi}_{k}(\sigma)=1$ for all $\sigma\in\mathbb{S}^{n-1}$ with $\vartheta(\sigma,a_{k})\leqslant\vartheta_{*}/6$. We also set $\widetilde{\phi}_{0}:=1-\sum_{k=1}^{K}\widetilde{\phi}_{k}$. Using them, we define
$P_{k}f:=\mathscr{F}^{-1}\left[\phi_{k}(\xi)\widehat{f}(\xi)\right],\qquad\phi_{k}(\xi):=\widetilde{\phi}_{k}\left(\xi/|\xi|\right)$ (3.15)
for $k=0,\dots,K$ and $f\in\mathscr{S}^{\prime}(\mathbb{R}^{n})$ with $0\notin\operatorname{supp}\widehat{f}$. We remark that there holds
$P_{m}\Delta_{j}g_{\lambda,\ell,k}$ $=e^{i\lambda^{\ell}a_{k}\cdot x}\mathscr{F}^{-1}\left[\widehat{\chi_{m,j}}(\xi+\lambda^{\ell}a_{k})\widehat{f_{k}}(\xi)\right]$ (3.16)
$=e^{i\lambda^{\ell}a_{k}\cdot x}\left(\chi_{m,j,\ell,k}*f_{k}\right)(x),$ (3.17)
where we have defined
$\chi_{m,j}:=\mathscr{F}^{-1}[\phi_{m}\widehat{\varphi_{j}}],\qquad\chi_{m,j,\ell,k}(x):=e^{-i\lambda^{\ell}a_{k}\cdot x}\chi_{m,j}(x).$ (3.18)
For these kernels, there holds ${\left\|\chi_{m,j,\ell,k}\right\|}_{L^{\rho}}={\left\|\chi_{m,j}\right\|}_{L^{\rho}}=2^{n(1-\frac{1}{\rho})j}{\left\|\chi_{m,0}\right\|}_{L^{\rho}}$ holds for $1\leqslant\rho\leqslant\infty$. Moreover, we prepare some notations; for every $k\in\{1,\dots,K\}$ and $\ell\in\{1,\dots,L\}$, set
$j_{\ell,k}:=\log_{2}\left(\lambda^{\ell}|a_{k}|\right),$ (3.19)
$R_{\ell,k}:=\{(m,j)\in\{0,\dots,K\}\times\mathbb{Z}\ ;\ (m,j)\neq(k,j_{\ell,k}),(k,j_{\ell,k}\pm 1)\}.$ (3.20)
For the proof of Proposition 3.1, we prepare a lemma.
Lemma 3.3.
There exists a constant $c_{*}=c_{*}(a_{1},\dots,a_{K})\in 2^{\mathbb{Z}}$ such that for any $\ell\in\{1,\dots,L\}$, $k\in\{1,\dots,K\}$, $(m,j)\in R_{\ell,k}$, and $\lambda\in 2^{\mathbb{N}}$, there holds
$\operatorname{supp}\left[\widehat{\chi_{m,j}}(\cdot+\lambda^{\ell}a_{k})\right]\subset\left\{\xi\in\mathbb{R}^{n}\ ;\ |\xi|\geqslant c_{*}\max\{2^{j},\lambda^{\ell}|a_{k}|\}\right\}.$ (3.21)
Proof.
Let $\xi\in\mathbb{R}^{n}$ satisfy $\widehat{\chi_{m,j}}(\xi+\lambda^{\ell}a_{k})\neq 0$. We first consider the case of $m=k$. Then, by $2^{j-1}\leqslant|\xi+\lambda^{\ell}a_{k}|\leqslant 2^{j+1}$ and $|\lambda^{\ell}a_{k}|=2^{j_{\ell,k}}$, it holds
$|\xi|$ $=|\xi+\lambda^{\ell}a_{k}-\lambda^{\ell}a_{k}|$ (3.22)
$\geqslant\begin{cases}|\xi+\lambda^{\ell}a_{k}|-|\lambda^{\ell}a_{k}|\geqslant 2^{j-1}-2^{j-2}&(j\geqslant j_{\ell,k}+2),\\ |\lambda^{\ell}a_{k}|-|\xi+\lambda^{\ell}a_{k}|\geqslant 2^{j_{\ell,k}}-2^{j_{\ell,k}-1}&(j\leqslant j_{\ell,k}-2),\end{cases}$ (3.23)
$\geqslant\frac{1}{4}\max\left\{2^{j},|\lambda^{\ell}a_{k}|\right\}.$ (3.24)
For the case of $m\neq k$, we see that
$|\xi|^{2}={}$ $|\xi+\lambda^{\ell}a_{k}|^{2}+|\lambda^{\ell}a_{k}|^{2}-2|\xi+\lambda^{\ell}a_{k}||\lambda^{\ell}a_{k}|\cos\vartheta(\xi+\lambda^{\ell}a_{k},\lambda^{\ell}a_{k})$ (3.25)
$={}$ $\frac{1}{2}\left(|\xi+\lambda^{\ell}a_{k}|-|\lambda^{\ell}a_{k}|\cos\vartheta(\xi+\lambda^{\ell}a_{k},\lambda^{\ell}a_{k})\right)^{2}$ (3.26)
$+\frac{1}{2}\left(|\lambda^{\ell}a_{k}|-|\xi+\lambda^{\ell}a_{k}|\cos\vartheta(\xi+\lambda^{\ell}a_{k},\lambda^{\ell}a_{k})\right)^{2}$ (3.27)
$+\frac{1}{2}\sin^{2}\vartheta(\xi+\lambda^{\ell}a_{k},\lambda^{\ell}a_{k})\left(|\lambda^{\ell}a_{k}|^{2}+|\xi+\lambda^{\ell}a_{k}|^{2}\right)$ (3.28)
$\geqslant{}$ $\frac{1}{2}\sin^{2}(\vartheta_{*}/6)\left(|\lambda^{\ell}a_{k}|^{2}+|\xi+\lambda^{\ell}a_{k}|^{2}\right),$ (3.29)
which yields $|\xi|\geqslant c\sin(\vartheta_{*}/6)\max\left\{2^{j},|\lambda^{\ell}a_{k}|\right\}$. Thus, we complete the proof. ∎
Let us now provide the proof of Proposition 3.1
Proof of Proposition 3.1.
We split the proof into two steps.
Step 1. The upper bound estimate. We may decompose $F_{\lambda,L}$ as
$F_{\lambda,L}={}\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{k^{\prime}=0}^{K}\sum_{j\in\mathbb{Z}}P_{k^{\prime}}\Delta_{j}g_{\lambda,\ell,k}={}F_{\lambda,L}^{({\rm p})}+F_{\lambda,L}^{({\rm r})},$ (3.30)
where the principal part $F_{\lambda,L}^{({\rm p})}$ and the remainder part $F_{\lambda,L}^{({\rm r})}$ are defined by
$F_{\lambda,L}^{({\rm p})}:=\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{|j-j_{\ell,k}|\leqslant 2}P_{k}\Delta_{j}g_{\lambda,\ell,k},$ (3.31)
$\begin{aligned} F_{\lambda,L}^{({\rm r})}:={}&\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{\begin{subarray}{c}0\leqslant k^{\prime}\leqslant K\\ k^{\prime}\neq k\end{subarray}}\sum_{|j-j_{\ell,k}|\leqslant 2}P_{k^{\prime}}\Delta_{j}g_{\lambda,\ell,k}\\ &+\sum_{k^{\prime}=0}^{K}\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{|j-j_{\ell,k}|\geqslant 3}P_{k^{\prime}}\Delta_{j}g_{\lambda,\ell,k}.\end{aligned}$ (3.32)
For the estimate of the principal part, we see that
${\left\|F_{\lambda,L}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{s}}$ $\leqslant\sum_{k=1}^{K}\left\{\sum_{j^{\prime}\in\mathbb{Z}}\left(2^{sj}{\left\|\sum_{\ell=1}^{L}\sum_{|j-j_{\ell,k}|\leqslant 2}P_{k}\Delta_{j}\Delta_{j^{\prime}}g_{\lambda,\ell,k}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}$ (3.33)
$=\sum_{k=1}^{K}\left\{\sum_{\ell=1}^{L}\sum_{|j^{\prime}-j_{\ell,k}|\leqslant 3}\left(2^{sj}{\left\|\sum_{|j-j_{\ell,k}|\leqslant 2}P_{k}\Delta_{j}\Delta_{j^{\prime}}g_{\lambda,\ell,k}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}$ (3.34)
$\leqslant C\sum_{k=1}^{K}\left\{\sum_{\ell=1}^{L}\sum_{|j^{\prime}-j_{\ell,k}|\leqslant 3}\left(2^{sj}{\left\|g_{\lambda,\ell,k}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}$ (3.35)
$=C\sum_{k=1}^{K}\left\{\sum_{\ell=1}^{L}\sum_{|j^{\prime}-j_{\ell,k}|\leqslant 3}\left(2^{sj}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}{\left\|f_{k}\right\|}_{L^{p}}$ (3.36)
$\leqslant C\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}\sum_{k=1}^{K}|a_{k}|^{s}{\left\|f_{k}\right\|}_{L^{p}}.$ (3.37)
For the remainder terms, there holds
${\left\|F_{\lambda,L}^{({\rm r})}\right\|}_{\dot{B}_{p,1}^{s}}$ $\leqslant\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{\begin{subarray}{c}0\leqslant k^{\prime}\leqslant K\\ k^{\prime}\neq k\end{subarray}}\sum_{|j-j_{\ell,k}|\leqslant 2}\sum_{|j^{\prime}-j|\leqslant 1}2^{sj^{\prime}}{\left\|P_{k^{\prime}}\Delta_{j}\Delta_{j^{\prime}}g_{\lambda,\ell,k}\right\|}_{L^{p}}$ (3.38)
$\qquad+\sum_{k=1}^{K}\sum_{k^{\prime}=0}^{K}\sum_{\ell=1}^{L}\sum_{|j-j_{\ell,k}|\geqslant 3}\sum_{|j^{\prime}-j|\leqslant 1}2^{sj^{\prime}}{\left\|P_{k^{\prime}}\Delta_{j}\Delta_{j^{\prime}}g_{\lambda,\ell,k}\right\|}_{L^{p}}$ (3.39)
$\leqslant C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{(k^{\prime},j)\in R_{\ell,k}}2^{sj}{\left\|P_{k^{\prime}}\Delta_{j}g_{\lambda,\ell,k}\right\|}_{L^{p}}.$ (3.40)
Here, Lemma 3.3 yields
${\left\|P_{m}\Delta_{j}g_{\lambda,\ell,k}\right\|}_{L^{p}}$ $\leqslant|b_{\ell}|\sum_{j^{\prime}\geqslant\max\{j,j_{\ell,k}\}+j_{*}}{\left\|\Delta_{j^{\prime}}\chi_{m,j,\ell,k}*f_{k}\right\|}_{L^{p}}$ (3.41)
$\leqslant C|b_{\ell}|\sum_{j^{\prime}\geqslant\max\{j,j_{\ell,k}\}+j_{*}}2^{-Mj^{\prime}}{\left\|\chi_{m,j,\ell,k}*\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.42)
$=C|b_{\ell}|\left(\max\{2^{j},\lambda^{\ell}|a_{k}|\}\right)^{-M}{\left\|\chi_{m,j,\ell,k}*\nabla^{M}f_{k}\right\|}_{L^{p}},$ (3.43)
where we have put $j_{*}=\log_{2}(2c_{*})$. Then, we obtain
${\left\|F_{\lambda,L}^{({\rm r})}\right\|}_{\dot{B}_{p,1}^{s}}\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{(k^{\prime},j)\in R_{\ell,k}}|b_{\ell}|2^{sj}\left(\max\{2^{j},\lambda^{\ell}|a_{k}|\}\right)^{-M}{\left\|\chi_{k^{\prime},j,\ell,k}*\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.44)
$\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{k^{\prime}=0}^{K}\sum_{j\geqslant j_{\ell,k}+1}|b_{\ell}|2^{(s-M)j}{\left\|\chi_{k^{\prime},j,\ell,k}\right\|}_{L^{1}}{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.45)
$+C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{k^{\prime}=0}^{K}\sum_{j\leqslant j_{\ell,k}}|b_{\ell}|2^{sj}(\lambda^{\ell}|a_{k}|)^{-M}{\left\|\chi_{k^{\prime},j,\ell,k}\right\|}_{L^{p}}{\left\|\nabla^{M}f_{k}\right\|}_{L^{1}}$ (3.46)
$\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{j\geqslant j_{\ell,k}+1}|b_{\ell}|2^{(s-M)j}{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.47)
$+C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{j\leqslant j_{\ell,k}}|b_{\ell}|2^{\{s+n(1-\frac{1}{p})\}j}(\lambda^{\ell}|a_{k}|)^{-M}{\left\|\nabla^{M}f_{k}\right\|}_{L^{1}}$ (3.48)
$\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s-M}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.49)
$+C\sum_{k=1}^{K}\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s+n(1-\frac{1}{p})-M}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{1}}.$ (3.50)
Combining (3.37) and (3.50), we obtain the desired estimate.
Step 2. The lower bound estimate. Since $P_{0},P_{1},\dots,P_{K}$ are bounded in homogeneous Besov spaces, there holds
${\left\|F_{\lambda,L}\right\|}_{\dot{B}_{p,q}^{s}}\geqslant c\max_{1\leqslant k\leqslant K}{\left\|P_{k}F_{\lambda,L}\right\|}_{\dot{B}_{p,q}^{s}}.$ (3.51)
Fix a $k\in\{1,\dots,K\}$ and decompose $P_{k}F_{\lambda,L}$ as
$P_{k}F_{\lambda,L}=F^{({\rm p})}_{\lambda,k,L}+F^{({\rm r})}_{\lambda,k,L}$; (3.52)
$\quad F^{({\rm p})}_{\lambda,k,L}:=\sum_{\ell=1}^{L}\sum_{|j-j_{\ell,k}|\leqslant 2}P_{k}\Delta_{j}g_{\lambda,\ell,k},$ (3.53)
$\quad F^{({\rm r})}_{\lambda,k,L}:=\sum_{\ell=1}^{L}\sum_{|j-j_{\ell,k}|\geqslant 3}P_{k}\Delta_{j}g_{\lambda,\ell,k}+\sum_{\begin{subarray}{c}1\leqslant k^{\prime}\leqslant K\\ k^{\prime}\neq k\end{subarray}}\sum_{\ell=1}^{L}P_{k}g_{\lambda,\ell,k^{\prime}}.$ (3.54)
For the lower bound estimate of the principal part $F^{({\rm p})}_{\lambda,k,L}$, we see that
${\left\|F_{\lambda,k,L}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{s}}$ $=\sum_{k=1}^{K}\left\{\sum_{\ell=1}^{L}\sum_{|j^{\prime}-j_{\ell,k}|\leqslant 3}\left(2^{sj^{\prime}}{\left\|\sum_{|j-j_{\ell,k}|\leqslant 2}P_{k}\Delta_{j}\Delta_{j^{\prime}}g_{\lambda,\ell,k}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}$ (3.55)
$\geqslant\sum_{k=1}^{K}\left\{\sum_{\ell=1}^{L}\left(2^{sj_{\ell,k}}{\left\|\sum_{|j-j_{\ell,k}|\leqslant 2}P_{k}\Delta_{j}\Delta_{j_{\ell,k}}g_{\lambda,\ell,k}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}$ (3.56)
$=\sum_{k=1}^{K}\left\{\sum_{\ell=1}^{L}\left(2^{sj_{\ell,k}}{\left\|P_{k}\Delta_{j_{\ell,k}}g_{\lambda,\ell,k}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}.$ (3.57)
Here, there holds
$P_{k}\Delta_{j_{\ell,k}}g_{\lambda,\ell,k}(x)$ $=b_{\ell}\int_{\mathbb{R}^{n}}\chi_{k,j_{\ell,k}}(y)f_{k}(x-y)e^{i\lambda^{\ell}a_{k}\cdot(x-y)}dy$ (3.58)
$=b_{\ell}e^{i\lambda^{\ell}a_{k}\cdot x}f_{k}(x)\int_{\mathbb{R}^{n}}\chi_{k,j_{\ell,k}}(y)e^{-i\lambda^{\ell}a_{k}\cdot y}dy$ (3.59)
$\qquad+b_{\ell}e^{i\lambda^{\ell}a_{k}\cdot x}\int_{\mathbb{R}^{n}}(f_{k}(x-y)-f_{k}(x))\chi_{k,j_{\ell,k}}(y)e^{-i\lambda^{\ell}a_{k}\cdot y}dy$ (3.60)
$=:b_{\ell}e^{i\lambda^{\ell}a_{k}\cdot x}f_{k}(x)+R_{\lambda,\ell,k}(x),$ (3.61)
where we have used $\widehat{\chi_{k,j_{\ell,k}}}(\lambda^{\ell}a_{k})=1$. The remainder term $R_{\lambda,\ell,k}$ is bounded as
${\left\|R_{\lambda,\ell,k}\right\|}_{L^{p}}$ $\leqslant|b_{\ell}|{\left\|\nabla f_{k}\right\|}_{L^{p}}\int_{\mathbb{R}^{n}}|y||\chi_{k,j_{\ell,k}}(y)|dy$ (3.62)
$\leqslant C|b_{\ell}|2^{-j_{\ell,k}}{\left\|\nabla f_{k}\right\|}_{L^{p}}$ (3.63)
$\leqslant C|b_{\ell}|(\lambda^{\ell}|a_{k}|)^{-1}{\left\|\nabla f_{k}\right\|}_{L^{p}}.$ (3.64)
Thus, we see that
${\left\|P_{k}\Delta_{j_{\ell,k}}g_{\lambda,\ell,k}\right\|}_{L^{p}}$ $\geqslant|b_{\ell}|{\left\|f_{k}\right\|}_{L^{p}}-C(\lambda^{\ell}|a_{k}|)^{-1}|b_{\ell}|{\left\|\nabla f_{k}\right\|}_{L^{p}},$ (3.65)
which implies
${\left\|F_{\lambda,k,L}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{s}}\geqslant{}$ $c\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}\max_{1\leqslant k\leqslant K}|a_{k}|^{s}{\left\|f_{k}\right\|}_{L^{p}}$ (3.66)
$-C\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s-1}|b_{\ell}|{\left\|\nabla f_{k}\right\|}_{L^{p}}.$ (3.67)
On the estimate of $F^{({\rm r})}_{\lambda,k,L}$, the similar strategy as for the estimate of $F^{({\rm r})}_{\lambda,k,L}$ yields
${\left\|F^{({\rm r})}_{\lambda,k,L}\right\|}_{\dot{B}_{p,q}^{s}}\leqslant{}$ $\sum_{\ell=1}^{L}\sum_{|j-j_{\ell,k}|\geqslant 3}\sum_{|j^{\prime}-j|\leqslant 1}2^{sj^{\prime}}{\left\|P_{k}\Delta_{j}\Delta_{j^{\prime}}g_{\lambda,\ell,k}\right\|}_{L^{p}}$ (3.68)
$+\sum_{\begin{subarray}{c}1\leqslant k^{\prime}\leqslant K\\ k^{\prime}\neq k\end{subarray}}\sum_{\ell=1}^{L}\sum_{j\in\mathbb{Z}}2^{sj}{\left\|P_{k}\Delta_{j}g_{\lambda,\ell,k}\right\|}_{L^{p}}$ (3.69)
$\leqslant{}$ $C\max_{1\leqslant k\leqslant K}\sum_{\ell=1}^{L}\sum_{(k^{\prime},j)\in R_{\ell,k}}2^{sj}{\left\|P_{k^{\prime}}\Delta_{j}g_{\lambda,\ell,k}\right\|}_{L^{p}}$ (3.70)
$\leqslant{}$ $C\max_{1\leqslant k\leqslant K}\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s-M}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.71)
$+C\max_{1\leqslant k\leqslant K}\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s+n(1-\frac{1}{p})-M}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}\cap L^{1}}.$ (3.72)
Gathering (3.67) and (3.72), we complete the proof. ∎
4. Principal part of solutions and perturbed forces
In this section, we construct the principal part of the non-unique steady flow outlined in Section 1.3, and we prove that the perturbative forcing term admits better regularity and smallness than the principal part. The central aim of this section is to establish the following theorem.
Theorem 4.1.
For $n\geqslant 3$, there exists a positive constant $\lambda_{*}>0$ such that for every $\lambda\in 2^{\mathbb{N}}$ with $\lambda\geqslant\lambda_{*}$, there exists a divergence-free vector field $V_{\lambda}\in\mathscr{S}^{\prime}(\mathbb{R}^{n})$ such that the following (i) and (ii) holds:
  1. (i)
    For any $p$ and $q$ satisfying (N1) or (N2) in Theorem 1.2, $V_{\lambda}$ belongs to $\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ and it holds
    $C^{-1}\lambda^{-\alpha(p)}\leqslant{\left\|V_{\lambda}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\leqslant C\lambda^{-\alpha(p)}$ (4.1)
    with $\alpha(p):=(3-n/p)(2n+4)-3$ and a positive constant $C=C(n,p,q)$ independent of $\lambda$.
  2. (ii)
    For any $n/2<r<n$, the perturbative force
    $F_{\lambda}:=-\Delta V_{\lambda}+\mathbb{P}\operatorname{div}(V_{\lambda}\otimes V_{\lambda})$ (4.2)
    belongs to $\dot{B}_{r,1}^{n/r-3}(\mathbb{R}^{n})$ and satisfies
    ${\left\|F_{\lambda}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-3}}\leqslant C\lambda^{-\beta}$ (4.3)
    with $\beta:=6n+10$ and a positive constant $C=C(n,r)$ independent of $\lambda$.
Remark 4.2.
Note that $\alpha(p)\leqslant\alpha(\infty)=6n+9<\beta$, which implies $F_{\lambda}$ is much smaller than $V_{\lambda}$.
4.1. Notations and key lemmas
Before proving Theorem 4.1, we collect several auxiliary results and notations. We begin by recalling the well-known geometric lemma known as the Nash lemma.
Lemma 4.3 ([Nash-54]).
Let $n\geqslant 2$. There exist a positive constant $r_{0}$, a bounded closed interval $\mathcal{I}\subset(0,\infty)$, a finite subset $\mathcal{K}$ of $\mathbb{Z}^{n}\setminus\{0\}$, and $C^{\infty}$ functions $\Gamma_{k}:B_{\mathcal{S}_{n}}({\rm Id},r_{0})\to\mathcal{I}$ $(k\in\mathcal{K})$ such that
$M=\sum_{k\in\mathcal{K}}\Gamma_{k}(M)^{2}k\otimes k,\qquad M\in B_{\mathcal{S}_{n}}({\rm Id},r_{0}).$ (4.4)
Here, $\mathcal{S}_{n}$ stands for the set of all real symmetric $n\times n$ matrices.
Proof.
For the readers’ convenience, we provide an outline of an elementary proof. Let us define $\mathcal{K}$ by
$\mathcal{K}:=\left\{\mathbbm{e}_{i}\ ;\ i=1,\dots,n\right\}$ $\cup\left\{\mathbbm{e}_{i}+\mathbbm{e}_{j}\ ;\ i,j=1,\dots,n,\ i\neq j\right\}$ (4.5)
$\cup\left\{\mathbbm{e}_{i}-\mathbbm{e}_{j}\ ;\ i,j=1,\dots,n,\ i\neq j\right\},$ (4.6)
where $\mathbbm{e}_{1},\dots,\mathbbm{e}_{n}\in\mathbb{R}^{n}$ stands for the $i$-th standard basis vector. We define $\{\Gamma_{k}\}_{k\in\mathcal{K}}$ by
$\Gamma_{k}(M):=\begin{cases}\left(M_{ii}-\dfrac{1}{2}\right)^{\frac{1}{2}},&(k=\mathbbm{e}_{i},\ i=1,\dots,n),\\[7.0pt] \left(\dfrac{1}{4(n-1)}\pm\dfrac{M_{ij}}{2}\right)^{\frac{1}{2}},&(k=\mathbbm{e}_{i}\pm\mathbbm{e}_{j},\ i,j=1,\dots,n,\ i\neq j),\end{cases}$ (4.7)
for all $M=(M_{ij})\in\mathcal{S}_{n}$ with $|M-{\rm Id}|\leqslant r_{0}:=(4(n-1))^{-1}$. Then, we may easily check that (4.4) holds true, and complete the proof. ∎
Here, we introduce some notation that will be used throughout this section.
  1. 1.
    Let $\{a_{k}\}_{k\in\mathcal{K}}\subset\mathbb{R}^{n}\setminus\{0\}$ fulfill $k\cdot a_{k}=0$ and $|a_{k}|\in 2^{\mathbb{N}}$ for all $k\in\mathcal{K}$, and also satisfy that $a_{k}$ and $a_{k^{\prime}}$ are not parallel unless $k=k^{\prime}$. For instance, we may choose
    $a_{k}:=\begin{cases}\mathbbm{e}_{i+1},&(k=\mathbbm{e}_{i},\ i=1,\dots,n-1),\\ \mathbbm{e}_{1},&(k=\mathbbm{e}_{n}),\\ \sqrt{2}\left(\mathbbm{e}_{i}\mp\mathbbm{e}_{j}\right),&(k=\mathbbm{e}_{i}\pm\mathbbm{e}_{j},\ i,j=1,\dots,n,\ i\neq j).\end{cases}$ (4.8)
  2. 2.
    Let $\mu:=2n+4$. For $\lambda\in 2^{\mathbb{N}}$, we define $\{\lambda_{j}\}_{j\in\mathbb{N}_{0}}$ inductively by
    $\lambda_{0}=\lambda,\qquad\lambda_{j}=\exp_{\lambda}\!\left(\exp_{2}\!\left(\lambda_{j-1}^{4\mu}\right)\right),$ (4.9)
    where we have used the notation $\exp_{a}(b)\coloneqq a^{b}$ for $a>0$ and $b\in\mathbb{R}$.
    We note that, throughout this section, $\lambda$ may be replaced by a larger dyadic number whenever necessary, without further comment. We emphasize that the required lower bound on $\lambda$ is chosen independently of $j$. Moreover, almost all quantities introduced below depend on the parameter $\lambda$; however, we do not record this dependence explicitly in the notation so as to avoid unnecessarily complicating it.
  3. 3.
    For $j\in\mathbb{N}$, set
    $\Lambda_{j}:=\left\{\ell\in\mathbb{N}\ ;\ \mu\log_{\lambda}\lambda_{j-1}\leqslant\ell\leqslant\log_{\lambda}\lambda_{j}\right\}.$ (4.10)
    In what follows, we choose $\lambda$ sufficiently large so that $\max\Lambda_{j}<\min\Lambda_{j+1}$ holds for all $j\in\mathbb{N}_{0}$.
  4. 4.
    We define $\{h_{j}\}_{j\in\mathbb{N}}$ by
    $h_{j}:=\sum_{\ell\in\Lambda_{j}}\frac{1}{\ell}.$ (4.11)
    Note that $h_{j}\sim\log_{2}\log_{\lambda}\lambda_{j}=\lambda_{j-1}^{4\mu}$.
  5. 5.
    We choose a $\rho\in C_{c}^{\infty}(\mathbb{R}^{n};[0,\infty))$ so that
    $\operatorname{supp}\rho=B_{\mathbb{R}^{n}}(0,1),\qquad\int_{\mathbb{R}^{n}}\rho(y)dy=1.$ (4.12)
    Let $\rho_{j}$ be the mollifier of the scale $\lambda_{j}^{-2}$; more precisely, we define
    $\rho_{j}(x):=\lambda_{j}^{2n}\rho\left(\lambda_{j}^{2}x\right).$ (4.13)
  6. 6.
    Using a function $\psi\in C^{\infty}(\mathbb{R})$ satisfying $\mathbbm{1}_{(-\infty,0]}\leqslant\psi\leqslant\mathbbm{1}_{(-\infty,1)}$, we define a sequence $\{\psi_{j}\}_{j\in\mathbb{N}_{0}}\subset C_{c}^{\infty}(\mathbb{R}^{n})$ by
    $\psi_{j}(x):=\psi(|x|-\lambda^{j}).$ (4.14)
  7. 7.
    Let us define the differential operator $\mathcal{D}$ by
    $\mathcal{D}v:=-\nabla v-(\nabla v)^{\top}+2(\operatorname{div}v){\rm Id}$ (4.15)
    for smooth vector field $v:\mathbb{R}^{n}\to\mathbb{R}^{n}$. Note that $\mathcal{D}v$ is a real symmetric matrix and it holds
    $\operatorname{div}\mathcal{D}v=-\Delta v+\nabla\operatorname{div}v=-\Delta\mathbb{P}v.$ (4.16)
In order to define the elements of the desired $V_{\lambda}$, we introduce the following operators.
Definition 4.4.
For $j\in\mathbb{N}$ and $v\in C_{c}^{\infty}(\mathbb{R}^{n})$ satisfying ${\left\|\mathcal{D}v\right\|}_{L^{\infty}}\leqslant r_{0}\lambda_{j-1}^{6}$, we define
${\mathcal{V}}_{j}[v]:=\sqrt{2}\lambda_{j-1}^{3}\lambda_{j}^{-2}\sum_{k\in\mathcal{K}}{\rho}_{j}*\operatorname{div}\mathcal{D}\left({\Gamma}_{j,k}[v]{\Psi}_{j,k}k\right),$ (4.17)
where we have set
${\Gamma}_{j,k}[v](x):=\Gamma_{k}\left({\rm Id}-\lambda_{j-1}^{-6}\mathcal{D}v(x)\right),$ (4.18)
${\Psi}_{j,k}(x):=\frac{1}{\sqrt{h_{j}}}\sum_{\ell\in\Lambda_{j}}\frac{1}{\sqrt{\ell}}\psi_{j}(x)\cos\left(\lambda^{\ell}a_{k}\cdot x\right).$ (4.19)
Let us establish elementary derivative estimates for each $\mathcal{V}_{j}[v]$. Although these bounds are far from optimal, they ensure the well-definedness of Definition 4.6 below, and they provide an essential foundation for obtaining the sharp estimates later on.
Lemma 4.5.
For every $m\in\mathbb{N}_{0}$, there exists a positive constant $C=C(m)$ such that
${\left\|\nabla^{m}{\mathcal{V}}_{j}[v]\right\|}_{L^{p}}\leqslant C\lambda_{j}^{2m+3}.$ (4.20)
for all $j\in\mathbb{N}$, $1\leqslant p\leqslant\infty$, and $v\in C_{c}^{\infty}(\mathbb{R}^{n})$ with ${\left\|\mathcal{D}v\right\|}_{L^{\infty}}\leqslant r_{0}\lambda_{j-1}^{6}$. Moreover, it holds
${\left\|\mathcal{D}{\mathcal{V}}_{j}[v]\right\|}_{L^{\infty}}\leqslant r_{0}\lambda_{j}^{6}.$ (4.21)
Proof.
From the Young inequality, we see that
${\left\|\nabla^{m}{\mathcal{V}}_{j}[v]\right\|}_{L^{p}}$ $\leqslant C\lambda_{j-1}^{3}\lambda_{j}^{-2}{\left\|\nabla^{m+2}\rho_{k}\right\|}_{L^{1}}\sum_{k\in\mathcal{K}}{\left\|\Gamma_{j,k}[v]\Psi_{j,k}\right\|}_{L^{p}}$ (4.22)
$\leqslant C\lambda_{j-1}^{3}\lambda_{j}^{2m+2}\sum_{k\in\mathcal{K}}{\left\|\Psi_{j,k}\right\|}_{L^{p}}.$ (4.23)
Since $\Psi_{j,k}$ is bounded as
${\left\|\Psi_{j,k}\right\|}_{L^{p}}$ $\leqslant\frac{C}{\sqrt{h_{j}}}\lambda^{\frac{n}{p}j}\sum_{\ell=1}^{\log_{\lambda}\lambda_{j}}\frac{1}{\sqrt{\ell}}$ (4.24)
$\leqslant C\lambda^{-2\mu}_{j-1}\lambda^{\frac{n}{p}j}\sqrt{\log_{\lambda}\lambda_{j}}$ (4.25)
$\leqslant C\lambda^{-2\mu}_{j-1}\lambda_{j},$ (4.26)
we obtain (4.20). For the estimate of (4.21), (4.20) with $p=\infty$ and $m=1$ yields
${\left\|\mathcal{D}\mathcal{V}[v]\right\|}_{L^{\infty}}\leqslant C{\left\|\nabla\mathcal{V}[v]\right\|}_{L^{\infty}}\leqslant C\lambda_{j}^{5}\leqslant r_{0}\lambda_{j}^{6}$ (4.27)
provided that $\lambda$ is sufficiently large. Thus, we complete the proof. ∎
We now define, by induction, a sequence of flows that will serve as building blocks for the velocity field required in Theorem 4.1.
Definition 4.6.
Let us define a sequence $\{{V}_{j}\}_{j\in\mathbb{N}_{0}}\subset C_{c}^{\infty}(\mathbb{R}^{n})$ iteratively by
$\begin{cases}{V}_{0}:=0,&(j=0),\\ {V}_{j}:={\mathcal{V}}_{j}[{V}_{j-1}],\quad&(j=1,2,3,\dots).\end{cases}$ (4.28)
Remark 4.7.
Some remarks are in order.
  1. 1.
    Note that (4.21) guarantees that $V_{j-1}$ belongs to the domain of the operator $\mathcal{V}_{j}[\cdot]$, and hence the sequence $\{V_{j}\}_{j\in\mathbb{N}_{0}}$ is well-defined.
  2. 2.
    By the definition of $\mathcal{V}_{j}$, we see that each $V_{j}$ is solenoidal.
  3. 3.
    It follows from
    $\operatorname{supp}V_{j-1}$ $\subset\operatorname{supp}\psi_{j-1}+B_{\mathbb{R}^{n}}(0,\lambda_{j-1}^{-2})$ (4.29)
    $\subset B_{\mathbb{R}^{n}}(0,\lambda^{j-1}+1+\lambda_{j-1}^{-2})$ (4.30)
    $\subset B_{\mathbb{R}^{n}}(0,\lambda^{j})$ (4.31)
    that $\psi_{j}(x)=1$ holds for all $x\in\operatorname{supp}V_{j-1}$. This fact is to be used for the cancellation of the worst nonlinearity.
To investigate the precise structure of $\{V_{j}\}_{j\in\mathbb{N}_{0}}$, we decompose it into the principal and remainder part as
$V_{j}=V^{(\rm p)}_{j}+V^{({\rm r})}_{j}$, (4.32)
where we have set
$V^{(\rm p)}_{j}(x):={}\sqrt{2}\lambda_{j-1}^{3}\sum_{k\in\mathcal{K}}\Gamma_{j,k}[V_{j-1}](x)\Psi_{j,k}(x)k$ (4.33)
and
$V^{(\rm r)}_{j}=V^{({\rm r},1)}_{j}+V^{({\rm r},2)}_{j}$, (4.34)
$\quad V^{({\rm r},1)}_{j}(x):={}\rho_{j}*V^{(\rm p)}_{j}(x)-V^{(\rm p)}_{j}(x),$ (4.35)
$\quad V^{({\rm r},2)}_{j}(x):={}\sqrt{2}\lambda_{j-1}^{3}\lambda_{j}^{-2}\rho_{j}*\sum_{k\in\mathcal{K}}\operatorname{div}\mathcal{D}\left(\Gamma_{j,k}[V_{j-1}](x)\psi_{j}(x)k\right)\cos\left(\lambda_{j}a_{k}\cdot x\right).$ (4.36)
Lemma 4.8.
Let $1\leqslant p,q\leqslant\infty$ and $s\in\mathbb{R}$ satisfy
$\frac{n}{p}-n<s<2n.$ (4.37)
Then, there exists a positive constant $C=C(n,p,q,s)$ such that
$C^{-1}\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q)\leqslant{\left\|V^{({\rm p})}_{j}\right\|}_{\dot{B}_{p,q}^{s}}\leqslant C\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q),$ (4.38)
and
${\left\|V^{({\rm r})}_{j}\right\|}_{\dot{B}_{p,q}^{s}}\leqslant C\lambda_{j}^{-1}\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q)$ (4.39)
for all $j\in\mathbb{N}$, where the sequence $\{\sigma_{j}(s,q)\}_{j\in\mathbb{N}}$ have been defined by
$\sigma_{j}(s,q):=\left(\displaystyle\sum_{\ell\in\Lambda_{j}}\frac{\lambda^{sq\ell}}{\ell^{\frac{q}{2}}}\right)^{\frac{1}{q}}.$ (4.40)
Remark 4.9.
The assumption $s<2n$ is not optimal, while this is enough for our purpose.
Proof.
To begin with, we prepare the $L^{\infty}$ estimates for the derivatives of $\Gamma_{j,k}[V_{j-1}]$. Since $\Gamma_{j,k}[V_{j-1}]$ is constant for $j=1$, we focus on the case of $j\geqslant 2$. Making use of (4.20) for $V_{j-1}=\mathcal{V}_{j-1}[V_{j-2}]$, it holds ${\left\|\nabla^{s}V_{j-1}\right\|}_{L^{\infty}}\leqslant C\lambda_{j-1}^{2s+3}$ ($s\in\mathbb{N}_{0}$). Thus, from the Faà di Bruno formula, it follows that
${\left\|\nabla^{m}\Gamma_{j,k}[V_{j-1}]\right\|}_{L^{\infty}}$ $\leqslant C\sum_{\ell=1}^{m}\sum_{s_{1}+\dots+s_{\ell}=m}\lambda_{j-1}^{-6\ell}{\left\|\nabla^{s_{1}}\mathcal{D}V_{j-1}\right\|}_{L^{\infty}}\cdots{\left\|\nabla^{s_{\ell}}\mathcal{D}V_{j-1}\right\|}_{L^{\infty}}$ (4.41)
$\leqslant C\sum_{\ell=1}^{m}\sum_{s_{1}+\dots+s_{\ell}=m}\lambda_{j-1}^{-6\ell}\cdot{\lambda_{j-1}^{2(s_{1}+1)+3}}\cdots{\lambda_{j-1}^{2(s_{\ell}+1)+3}}$ (4.42)
$=C\lambda_{j-1}^{2m}\sum_{\ell=1}^{m}\lambda_{j-1}^{-\ell}$ (4.43)
$\leqslant C\lambda_{j-1}^{2m}.$ (4.44)
Let us prove (4.38). It follows from Proposition 3.1 with $M=2n$ that
${\left\|V^{({\rm p})}_{j}\right\|}_{\dot{B}_{p,q}^{s}}$ $\leqslant C\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q)\sum_{k\in\mathcal{K}}|a_{k}|^{s}{\left\|\Gamma_{j,k}[V_{j-1}]\psi_{j}\right\|}_{L^{p}}$ (4.45)
$\quad+C\lambda_{j-1}^{3-2\mu}\sum_{\ell\in\Lambda_{j}}\frac{\lambda^{(s-n)\ell}}{\sqrt{\ell}}\sum_{k\in\mathcal{K}}|a_{k}|^{s-n}{\left\|\nabla^{2n}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}\cap L^{1}}$ (4.46)
$\leqslant C\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q)\left(1+\lambda_{j-1}^{(4-\mu)n}\right)$ (4.47)
$\leqslant C\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q),$ (4.48)
where we have used
$\sum_{\ell\in\Lambda_{j}}\frac{\lambda^{(s-n)\ell}}{\sqrt{\ell}}\leqslant\left(\sum_{\ell\in\Lambda_{j}}\lambda^{-nq^{\prime}\ell}\right)^{\frac{1}{q^{\prime}}}\sigma_{j}(s,q)\leqslant\lambda_{j-1}^{-n\mu}\sigma_{j}(s,q)$ (4.49)
and
${\left\|\nabla^{2n}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}\cap L^{1}}\leqslant C\lambda_{j-1}^{4n}.$ (4.50)
For the estimate of $V_{j}^{({\rm p})}$ from below, it holds by Proposition 3.1 and similar calculations as above that
${\left\|V^{({\rm p})}_{j}\right\|}_{\dot{B}_{p,q}^{s}}$ $\geqslant C^{-1}\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q)\max_{k\in\mathcal{K}}|a_{k}|^{s}{\left\|\Gamma_{j,k}[V_{j-1}]\psi_{j}\right\|}_{L^{p}}$ (4.51)
$\quad-C\lambda_{j-1}^{3-2\mu}\sum_{\ell\in\Lambda_{j}}\frac{\lambda^{(s-1)\ell}}{\sqrt{\ell}}\max_{k\in\mathcal{K}}|a_{k}|^{s-1}{\left\|\nabla(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}}$ (4.52)
$\quad-C\lambda_{j-1}^{3-2\mu}\sum_{\ell\in\Lambda_{j}}\frac{\lambda^{(s-n)\ell}}{\sqrt{\ell}}\max_{k\in\mathcal{K}}|a_{k}|^{s-n}{\left\|\nabla^{2n}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}\cap L^{1}}$ (4.53)
$\geqslant C^{-1}\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q)\left(1-C^{2}\lambda_{j-1}^{2-\mu}-C^{2}\lambda_{j-1}^{(4-\mu)n}\right)$ (4.54)
$\geqslant C^{-1}\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q).$ (4.55)
Next, we show (4.39). It follows from $\int_{\mathbb{R}^{n}}\rho_{j}(y)dy=1$ and the mean value formula that
$\Delta_{\ell}V^{({\rm r},1)}_{j}(x)$ $=\int_{\mathbb{R}^{n}}\rho_{j}(y)\left(\Delta_{\ell}V^{({\rm p})}_{j}(x-y)-\Delta_{\ell}V^{({\rm p})}_{j}(x)\right)dy$ (4.56)
$=-\sum_{i=1}^{n}\int_{\mathbb{R}^{n}}y_{i}\rho_{j}(y)\int_{0}^{1}\partial_{x_{i}}\Delta_{\ell}V^{({\rm p})}_{j}(x-\theta y)d\theta dy,$ (4.57)
which implies
${\left\|\Delta_{\ell}V^{({\rm r},1)}_{j}\right\|}_{L^{p}}$ $\leqslant\int_{\mathbb{R}^{n}}|y|\rho_{j}(y)dy{\left\|\nabla\Delta_{\ell}V^{({\rm p})}_{j}\right\|}_{L^{p}}=C\lambda_{j}^{-2}{\left\|\nabla\Delta_{\ell}V^{({\rm p})}_{j}\right\|}_{L^{p}}.$ (4.58)
Thus, we have
${\left\|V^{({\rm r},1)}_{j}\right\|}_{\dot{B}_{p,q}^{s}}$ $\leqslant C\lambda_{j}^{-2}{\left\|V^{({\rm p})}_{j}\right\|}_{\dot{B}_{p,q}^{s+1}}$ (4.59)
$\leqslant C\lambda_{j}^{-2}\lambda_{j-1}^{3-2\mu}\sigma_{j}(s+1,q)$ (4.60)
$\leqslant C\lambda_{j}^{-1}\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q).$ (4.61)
We follow the same argument as in the proof of (4.38) to obtain
${\left\|V^{({\rm r},2)}_{j}\right\|}_{\dot{B}_{p,q}^{s}}$ $\leqslant C\lambda_{j}^{-2}\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q)\sum_{k\in\mathcal{K}}|a_{k}|^{s}{\left\|\nabla^{2}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}}$ (4.62)
$\quad+C\lambda_{j}^{-2}\lambda_{j-1}^{3-2\mu}\sum_{\ell\in\Lambda_{j}}\frac{\lambda^{(s-n)\ell}}{\sqrt{\ell}}\sum_{k\in\mathcal{K}}|a_{k}|^{s-n}{\left\|\nabla^{2n+2}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}\cap L^{1}}$ (4.63)
$\leqslant C\lambda_{j}^{-2}\lambda_{j-1}^{7-2\mu}\sigma_{j}(s,q)\left(1+\lambda_{j-1}^{(4-\mu)n}\right)$ (4.64)
$\leqslant C\lambda_{j}^{-2}\lambda_{j-1}^{7-2\mu}\sigma_{j}(s,q)$ (4.65)
$\leqslant C\lambda_{j}^{-1}\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q).$ (4.66)
This completes the proof. ∎
4.2. Proof of Theorem 4.1
Now, we are in a position to prove Theorem 4.1. We separate the proof into two steps.
Step 1. Construction and estimates of the principal stationary flow. Let us define
$\widetilde{\sigma}_{j}(s,q):={\left\|\left\{\sigma_{j}(s,q)\right\}_{j^{\prime}}\right\|}_{\ell^{q}(j^{\prime}\geqslant j)},\qquad j\in\mathbb{N}_{0}.$ (4.67)
It follows from Lemmas 4.8 that
$\sum_{j=1}^{\infty}{\left\|V_{j}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}$ $\leqslant C\sum_{j=1}^{\infty}\lambda_{j-1}^{3-2\mu}\sigma_{j}\left(\frac{n}{p}-1,q\right)$ (4.68)
$\leqslant C\lambda^{3-2\mu}\widetilde{\sigma}_{1}\left(\frac{n}{p}-1,q\right)$ (4.69)
$\leqslant C\lambda^{3-2\mu}\frac{\lambda^{(\frac{n}{p}-1)\mu}}{\sqrt{\mu}}=C\lambda^{3-(3-\frac{n}{p})\mu}.$ (4.70)
Hence, the series
$V^{({\rm p})}:=\sum_{j=1}^{\infty}V_{j}^{({\rm p})}$ (4.71)
absolutely converges in the strong topology of $\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ and we see that
${\left\|V^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\leqslant\sum_{j=1}^{\infty}{\left\|V_{j}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\leqslant C\lambda^{3-(3-\frac{n}{p})\mu}.$ (4.72)
For the estimate of $V^{({\rm p})}$ from below, it holds
${\left\|V^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\geqslant{\left\|V_{1}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}-\sum_{j=2}^{\infty}{\left\|V_{j}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}.$ (4.73)
Using (4.38), we have
${\left\|V_{1}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}$ $\geqslant C^{-1}\lambda^{3-2\mu}\sigma_{1}\left(\frac{n}{p}-1,q\right)$ (4.74)
$\geqslant C^{-1}\lambda^{3-2\mu}\frac{\lambda^{(\frac{n}{p}-1)\mu}}{\sqrt{\mu}}=C^{-1}\lambda^{3-(3-\frac{n}{p})\mu}.$ (4.75)
Similarly to (4.68), we see that
$\sum_{j=2}^{\infty}{\left\|V_{j}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}$ $\leqslant C\sum_{j=2}^{\infty}\lambda_{j-1}^{3-2\mu}\sigma_{j}\left(\frac{n}{p}-1,q\right)$ (4.76)
$\leqslant C\lambda_{1}^{3-2\mu}\widetilde{\sigma}_{2}\left(\frac{n}{p}-1,q\right)$ (4.77)
$\leqslant C\lambda_{1}^{3-(3-\frac{n}{p})\mu}.$ (4.78)
Hence, we obtain
${\left\|V^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\geqslant C^{-1}\lambda^{3-(3-\frac{n}{p})\mu}\left\{1-C^{2}\left(\frac{\lambda}{\lambda_{1}}\right)^{3-(3-\frac{n}{p})\mu}\right\}\geqslant C^{-1}\lambda^{3-(3-\frac{n}{p})\mu}.$ (4.79)
for sufficiently large $\lambda$.
Next, we consider the remainder part. By Lemma 4.8, there holds
$\sum_{j=1}^{\infty}{\left\|V_{j}^{({\rm r})}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}$ $\leqslant C\sum_{j=1}^{\infty}\lambda_{j}^{-1}\lambda_{j-1}^{3-2\mu}\sigma_{j}\left(\frac{n}{r}-1,1\right)$ (4.80)
$\leqslant C\sum_{j=1}^{\infty}\lambda_{j}^{\frac{n}{r}-2}\lambda_{j-1}^{3-2\mu}$ (4.81)
$\leqslant C\lambda_{1}^{\frac{n}{r}-2}\lambda^{3-2\mu}.$ (4.82)
Thus, due to $r>n/2$, we see that the series
$V^{({\rm r})}:=\sum_{j=1}^{\infty}V_{j}^{({\rm r})}$ (4.83)
absolutely converges in the strong topology of $\dot{B}_{r,1}^{n/r-1}(\mathbb{R}^{n})$ and
${\left\|V^{({\rm r})}\right\|}_{\dot{B}_{r,1}^{n/r-1}}$ $\leqslant\sum_{j=1}^{\infty}{\left\|V_{j}^{({\rm r})}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}$ (4.84)
$\leqslant C\lambda_{1}^{\frac{n}{r}-2}\lambda^{3-2\mu}=C\lambda^{(1-\frac{n}{p})\mu}\lambda_{1}^{\frac{n}{r}-2}\lambda^{3-(3-\frac{n}{p})\mu}.$ (4.85)
Now, we define $V:=V^{({\rm p})}+V^{({\rm r})}$. Then, we see that
$\begin{aligned} {\left\|V\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}&\leqslant{\left\|V^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}+C{\left\|V^{({\rm r})}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\\ &\leqslant C\left(1+\lambda^{(1-\frac{n}{p})\mu}\lambda_{1}^{\frac{n}{r}-2}\right)\lambda^{3-(3-\frac{n}{p})\mu}\\ &\leqslant C\lambda^{3-(3-\frac{n}{p})\mu},\end{aligned}$ (4.86)
$\begin{aligned} {\left\|V\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}&\geqslant{\left\|V^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}-C{\left\|V^{({\rm r})}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\\ &\geqslant C^{-1}\left(1-C\lambda^{(1-\frac{n}{p})\mu}\lambda_{1}^{\frac{n}{r}-2}\right)\lambda^{3-(3-\frac{n}{p})\mu}\\ &\geqslant C^{-1}\lambda^{3-(3-\frac{n}{p})\mu},\end{aligned}$ (4.87)
which implies the optimal estimate (4.1).
Step 2. Estimates of the perturbed force. In order to gain a better regularity and estimate for the force
$F=-\Delta V+\mathbb{P}\operatorname{div}(V\otimes V),$ (4.88)
we focus on its precise structure. To this end, we may decompose $F$ as
$F=F_{1}+F_{2}+F_{3};$ (4.89)
$\quad F_{1}:={}\sum_{j=1}^{\infty}\left(-\Delta V_{j-1}+\mathbb{P}\operatorname{div}\left(V^{({\rm p})}_{j}\otimes V^{({\rm p})}_{j}\right)\right)$, (4.90)
$\quad F_{2}:=\sum_{j_{1}\neq j_{2}}\mathbb{P}\operatorname{div}\left(V^{({\rm p})}_{j_{1}}\otimes V^{({\rm p})}_{j_{2}}\right)$, (4.91)
$\quad F_{3}:=\mathbb{P}\operatorname{div}\left(V^{({\rm p})}\otimes V^{({\rm r})}+V^{({\rm r})}\otimes V^{({\rm p})}+V^{({\rm r})}\otimes V^{({\rm r})}\right).$ (4.92)
We first consider the estimate for $F_{1}$, which includes the most crucial cancellation in our analysis. Using the formula $2\cos^{2}\theta=1+\cos 2\theta$ and
$\lambda_{j-1}^{6}\psi_{j}(x)^{2}\sum_{k\in\mathcal{K}}\Gamma_{k}\left({\rm Id}-\lambda_{j-1}^{-6}\mathcal{D}V_{j-1}\right)^{2}k\otimes k$ $=\lambda_{j-1}^{6}\psi_{j}(x)^{2}\left({\rm Id}-\lambda_{j-1}^{-6}\mathcal{D}V_{j-1}\right)$ (4.93)
$=\lambda_{j-1}^{6}\psi_{j}^{2}(x){\rm Id}-\mathcal{D}V_{j-1},$ (4.94)
which is implied by Lemma 4.3 and $\psi_{j}(x)=1$ for $x\in\operatorname{supp}V_{j-1}$, we have
$V^{({\rm p})}_{j}\otimes V^{({\rm p})}_{j}$={} $2\lambda_{j-1}^{6}\sum_{k_{1},k_{2}\in\mathcal{K}}\Gamma_{j,k_{1}}[V_{j-1}]\Gamma_{j,k_{2}}[V_{j-1}]\Psi_{j,k_{1}}\Psi_{j,k_{2}}k_{1}\otimes k_{2}$ (4.95)
$={}$ $\lambda_{j-1}^{6}\psi_{j}^{2}\sum_{k\in\mathcal{K}}\Gamma_{k}\left({\rm Id}-\lambda_{j-1}^{6}\mathcal{D}V_{j-1}\right)^{2}k\otimes k$ (4.96)
$+\frac{\lambda_{j-1}^{6}}{h_{j}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\Lambda_{j}}\frac{1}{\ell}\Gamma_{k}\left({\rm Id}-\lambda_{j-1}^{-6}\mathcal{D}V_{j-1}\right)^{2}\psi_{j}^{2}\cos\left(2\lambda^{\ell}a_{k}\cdot x\right)k\otimes k$ (4.97)
$+2\lambda_{j-1}^{6}\sum_{k_{1}\neq k_{2}}\Gamma_{j,k_{1}}[V_{j-1}]\Gamma_{j,k_{2}}[V_{j-1}]\Psi_{j,k_{1}}\Psi_{j,k_{2}}k_{1}\otimes k_{2}$ (4.98)
$={}$ $\lambda_{j-1}^{6}\psi_{j}^{2}{\rm Id}-\mathcal{D}V_{j-1}+F_{1,1;j}+F_{1,2;j},$ (4.99)
where we have set
$F_{1,1;j}:=\frac{\lambda_{j-1}^{6}}{h_{j}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\Lambda_{j}}\frac{1}{\ell}\Gamma_{j,k}\left[V_{j-1}\right]^{2}\psi_{j}^{2}\cos\left(2\lambda^{\ell}a_{k}\cdot x\right)k\otimes k,$ (4.100)
$F_{1,2;j}:=2\lambda_{j-1}^{6}\sum_{\begin{subarray}{c}k_{1},k_{2}\in\mathcal{K}\\ k_{1}\neq k_{2}\end{subarray}}\Gamma_{j,k_{1}}[V_{j-1}]\Gamma_{j,k_{2}}[V_{j-1}]\Psi_{j,k_{1}}\Psi_{j,k_{2}}k_{1}\otimes k_{2}.$ (4.101)
Then, from the cancellation property
$-\Delta V_{j-1}+\mathbb{P}\operatorname{div}\left(\lambda_{j-1}^{6}\psi_{j}^{2}{\rm Id}-\mathcal{D}V_{j-1}\right)=-\Delta V_{j-1}+\mathbb{P}\nabla(\lambda_{j-1}^{6}\psi_{j}^{2})+\Delta V_{j-1}=0$ (4.102)
it follows that
$F_{1}$ $=\sum_{j=1}^{\infty}\left(-\Delta V_{j-1}+\mathbb{P}\operatorname{div}\left(\lambda_{j-1}^{6}\psi_{j}^{2}{\rm Id}-\mathcal{D}V_{j-1}+F_{1,1;j}+F_{1,2;j}\right)\right)$ (4.103)
$=\sum_{j=1}^{\infty}\mathbb{P}\operatorname{div}F_{1,1;j}+\sum_{j=1}^{\infty}\mathbb{P}\operatorname{div}F_{1,2;j}$ (4.104)
$=:F_{1,1}+F_{1,2}.$ (4.105)
To proceed with the estimate of $F_{1,1}$, we begin with the estimate of each element $F_{1,1;j}$. It follows from Proposition 3.1 with $M=n$ that
${\left\|F_{1,1;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}$ $\leqslant C\lambda_{j-1}^{6-4\mu}\sum_{\ell\in\lambda_{j}}\frac{\lambda^{(\frac{n}{r}-2)\ell}}{\ell}\sum_{k\in\mathcal{K}}{\left\|\Gamma_{j,k}[V_{j-1}]\psi_{j}\right\|}_{L^{r}}$ (4.106)
$\quad+C\lambda_{j-1}^{6-4\mu}\sum_{\ell\in\lambda_{j}}\frac{\lambda^{-2\ell}}{\ell}\sum_{k\in\mathcal{K}}{\left\|\nabla^{n}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{r}\cap L^{1}}$ (4.107)
$\leqslant C\lambda_{j-1}^{6-4\mu}\cdot C\lambda_{j-1}^{(\frac{n}{r}-2)\mu}\cdot C\lambda^{\frac{n}{r}}+C\lambda_{j-1}^{6-4\mu}\cdot C\lambda_{j-1}^{-2\mu}\cdot C\lambda^{n}\lambda_{j-1}^{2n}$ (4.108)
$\leqslant C\lambda_{j-1}^{2n+6-4\mu}.$ (4.109)
For the estimate of $F_{1,2;j}$, since it is rewritten as
$F_{1,2;j}(x)=\frac{\lambda_{j-1}^{6}}{h_{j}}\sum_{\pm}\sum_{\ell_{1},\ell_{2}\in\Lambda_{j}}\sum_{\begin{subarray}{c}k_{1}\neq k_{2}\end{subarray}}$ $\frac{1}{\sqrt{\ell_{1}\ell_{2}}}\Gamma_{j,k_{1}}[V_{j-1}](x)\Gamma_{j,k_{2}}[V_{j-1}](x)$ (4.110)
$\psi_{j}(x)^{2}\cos\left((\lambda^{\ell_{1}}a_{k_{1}}\pm\lambda^{\ell_{2}}a_{k_{2}})\cdot x\right)k_{1}\otimes k_{2},$ (4.111)
we see that
${\left\|F_{1,2;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}\leqslant{}$ $C\lambda_{j-1}^{6-2\mu}\sum_{\pm}\sum_{\begin{subarray}{c}\ell_{1}\in\Lambda_{j}\end{subarray}}\sum_{\ell_{2}:\ell_{2}\geqslant\ell_{1}}\sum_{\begin{subarray}{c}k_{1}\neq k_{2}\end{subarray}}$ (4.112)
$\left(\frac{\lambda^{(\frac{n}{r}-2)\ell_{2}}}{\sqrt{\ell_{1}\ell_{2}}}|\lambda^{\ell_{1}-\ell_{2}}a_{k_{1}}\pm a_{k_{2}}|^{\frac{n}{r}-2}{\left\|\Gamma_{j,k_{1}}[V_{j-1}]\Gamma_{j,k_{2}}[V_{j-1}]\psi_{j}^{2}\right\|}_{L^{r}}\right$. (4.113)
$\left.\quad+\frac{\lambda^{-2\ell_{2}}}{\sqrt{\ell_{1}\ell_{2}}}|\lambda^{\ell_{1}-\ell_{2}}a_{k_{1}}\pm a_{k_{2}}|^{-2}{\left\|\nabla^{n}(\Gamma_{j,k_{1}}[V_{j-1}]\Gamma_{j,k_{2}}[V_{j-1}]\psi_{j}^{2})\right\|}_{L^{r}}\right)$ (4.114)
$\leqslant{}$ $C\lambda_{j-1}^{6-4\mu}\cdot C\lambda_{j-1}^{(\frac{n}{r}-2)\mu}\cdot C\lambda^{\frac{n}{r}}+C\lambda_{j-1}^{6-4\mu}\cdot C\lambda_{j-1}^{-2\mu}\cdot C\lambda^{n}\lambda_{j-1}^{2n}$ (4.115)
$\leqslant{}$ $C\lambda_{j-1}^{2n+6-4\mu},$ (4.116)
where we have used
$|\lambda^{\ell_{1}-\ell_{2}}a_{k_{1}}\pm a_{k_{2}}|^{2}={}$ $\left(\lambda^{\ell_{1}-\ell_{2}}|a_{k_{1}}|\pm|a_{k_{2}}|\cos\vartheta(a_{k_{1}},a_{k_{2}})\right)^{2}+|a_{k_{2}}|^{2}\sin^{2}\vartheta(a_{k_{1}},a_{k_{2}})$ (4.117)
$\geqslant{}$ $|a_{k_{2}}|^{2}\sin^{2}\vartheta(a_{k_{1}},a_{k_{2}})>0.$ (4.118)
Here, one may recall (3.13) for the definition of $\vartheta(\cdot,\cdot)$. Hence, we obtain
${\left\|F_{1}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-3}}$ $\leqslant C\sum_{j=1}^{\infty}\left({\left\|F_{1,1;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}+{\left\|F_{1,2;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}\right)$ (4.119)
$\leqslant C\sum_{j=1}^{\infty}\lambda_{j-1}^{2n+6-4\mu}$ (4.120)
$\leqslant{}C\lambda^{2n+6-4\mu}=C\lambda^{-\beta}.$ (4.121)
Next, we consider the estimate on $F_{2}$. Let us rewrite $F_{2}$ as
$F_{2}=\sum_{\begin{subarray}{c}j_{1},j_{2}\in\mathbb{N}\\ j_{1}\neq j_{2}\end{subarray}}\mathbb{P}\operatorname{div}F_{2;j_{1},j_{2}},$ (4.122)
where we have set
$F_{2;j_{1},j_{2}}:={}V^{({\rm p})}_{j_{1}}\otimes V^{({\rm p})}_{j_{2}}$ (4.123)
$\quad=2\lambda_{j_{1}-1}^{3}\lambda_{j_{2}-1}^{3}\sum_{k_{1},k_{2}\in\mathcal{K}}\Gamma_{j_{1},k_{1}}[V_{j_{1}-1}]\Gamma_{j_{2},k_{2}}[V_{j_{2}-1}]\Psi_{j_{1},k_{1}}\Psi_{j_{2},k_{2}}k_{1}\otimes k_{2}$ (4.124)
$\quad=\frac{\lambda_{j_{1}-1}^{3}\lambda_{j_{2}-1}^{3}}{\sqrt{h_{j_{1}}h_{j_{2}}}}\sum_{\pm}\sum_{\ell_{1}\in\Lambda_{j_{1}}}\sum_{\ell_{2}\in\Lambda_{j_{2}}}\sum_{k_{1},k_{2}\in\mathcal{K}}\frac{1}{\sqrt{\ell_{1}\ell_{2}}}\Gamma_{j_{1},k_{1}}[V_{j_{1}-1}](x)\Gamma_{j_{2},k_{2}}[V_{j_{2}-1}](x)$ (4.125)
$\qquad\qquad\qquad\qquad\qquad\qquad\psi_{j_{1}}(x)\psi_{j_{2}}(x)\cos\left((\lambda^{\ell_{1}}a_{k_{1}}\pm\lambda^{\ell_{2}}a_{k_{2}})\cdot x\right)k_{1}\otimes k_{2}.$ (4.126)
As $F_{2;j_{1},j_{2}}=F_{2;j_{2},j_{1}}$, it suffices to consider the case $j_{1}<j_{2}$. Since it holds
$|\lambda^{\ell_{1}-\ell_{2}}a_{k_{1}}\pm a_{k_{2}}|\geqslant|a_{k_{2}}|-\lambda^{\ell_{1}-\ell_{2}}|a_{k_{1}}|\geqslant\frac{1}{2}|a_{k_{2}}|>0,$ (4.127)
we follow the similar strategy as for the estimate of $F_{1,2;j}$ to see that
${\left\|F_{2;j_{1},j_{2}}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}\leqslant{}$ $C\lambda_{j_{1}-1}^{3-2\mu}\lambda_{j_{2}-1}^{3-2\mu}\sum_{\ell_{1}\in\Lambda_{j_{1}}}\sum_{\ell_{2}\in\Lambda_{j_{2}}}\sum_{k_{1},k_{2}\in\mathcal{K}}$ (4.128)
$\left(\frac{\lambda^{(\frac{n}{r}-2)\ell_{2}}}{\sqrt{\ell_{1}\ell_{2}}}|\lambda^{\ell_{1}-\ell_{2}}a_{k_{1}}\pm a_{k_{2}}|^{\frac{n}{r}-2}{\left\|\Gamma_{j_{1},k_{1}}[V_{j_{1}-1}]\Gamma_{j_{2},k_{2}}[V_{j_{2}-1}]\psi_{j_{1}}\right\|}_{L^{r}}\right$. (4.129)
$\left.\quad+\frac{\lambda^{-2\ell_{2}}}{\sqrt{\ell_{1}\ell_{2}}}|\lambda^{\ell_{1}-\ell_{2}}a_{k_{1}}\pm a_{k_{2}}|^{-2}{\left\|\nabla^{n}(\Gamma_{j_{1},k_{1}}[V_{j_{1}-1}]\Gamma_{j_{2},k_{2}}[V_{j_{2}-1}]\psi_{j_{1}})\right\|}_{L^{r}}\right)$ (4.130)
$\leqslant{}$ $C\lambda^{\frac{n}{r}j_{1}}\lambda_{j_{1}-1}^{3-2\mu}\lambda_{j_{2}-1}^{3-2\mu}\sum_{\ell_{1}\in\Lambda_{j_{1}}}\sum_{\ell_{2}:\ell_{2}\geqslant\ell_{1}}\lambda^{(\frac{n}{r}-2)\ell_{2}}$ (4.131)
$+C\lambda^{nj_{1}}\lambda_{j_{1}-1}^{3-2\mu}\lambda_{j_{2}-1}^{3-2\mu+2n}\sum_{\ell_{1}\in\Lambda_{j_{1}}}\sum_{\ell_{2}:\ell_{2}\geqslant\ell_{1}}\lambda^{-2\ell_{2}}$ (4.132)
$\leqslant{}$ $C\lambda^{\frac{n}{r}j_{1}}\lambda_{j_{1}-1}^{3-(4-\frac{n}{r})\mu}\lambda_{j_{2}-1}^{3-2\mu}+C\lambda_{j_{1}-1}^{3-4\mu}\lambda_{j_{2}-1}^{3-2\mu+2n}$ (4.133)
$\leqslant{}$ $C\lambda^{\frac{n}{r}j_{1}}\lambda_{j_{1}-1}^{3-(4-\frac{n}{r})\mu}\lambda_{j_{2}-1}^{3-2\mu}.$ (4.134)
Hence, we obtain
${\left\|F_{2}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-3}}$ $\leqslant C\sum_{j_{1}=1}^{\infty}\sum_{j_{2}=j_{1}+1}^{\infty}{\left\|F_{2;j_{1},j_{2}}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}$ (4.135)
$\leqslant C\sum_{j_{1}=1}^{\infty}\lambda^{\frac{n}{r}j_{1}}\lambda_{j_{1}-1}^{3-(4-\frac{n}{r})\mu}\sum_{j_{2}=j_{1}+1}^{\infty}\lambda_{j_{2}-1}^{3-2\mu}$ (4.136)
$\leqslant C\sum_{j_{1}=1}^{\infty}\lambda_{j_{1}-1}^{6-(6-\frac{n}{r})\mu}$ (4.137)
$\leqslant C\lambda^{6-(6-\frac{n}{r})\mu}\leqslant C\lambda^{-\beta}.$ (4.138)
Finally, we focus on the estimate of $F_{3}$. Let $n<p_{*}<2n$ satisfy (2.7) with $p=p_{*}$. It then follows from Lemma 2.1 that
${\left\|F_{3}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-3}}$ $\leqslant C\left({\left\|V^{({\rm p})}\right\|}_{\dot{B}_{p_{*},1}^{\frac{n}{p_{*}}-1}}+{\left\|V^{({\rm r})}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\right){\left\|V^{({\rm r})}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}$ (4.139)
$\leqslant C\left(\lambda^{3-(3-\frac{n}{p_{*}})\mu}+\lambda_{1}^{\frac{n}{r}-2}\lambda^{3-2\mu}\right)\lambda_{1}^{\frac{n}{r}-2}\lambda^{3-2\mu}$ (4.140)
$\leqslant C\lambda_{1}^{\frac{n}{r}-2}\lambda^{6-4\mu}\leqslant C\lambda^{-\beta}.$ (4.141)
Combining (4.121), (4.138), and (4.141), we obtain (4.3). Thus, we complete the proof.
5. Proof of Theorem 1.2
Now, we are ready to present the proof of Theorem 1.2.
Proof.
We separate the proof into two steps; the first step achieves the construction of the non-trivial solutions to the unforced stationary Navier–Stokes equations, and the second step establishes the desired non-stationary flow by the standard stability argument around the steady state provided in the previous step.
Step 1. Steady Navier–Stokes flow around $V_{\lambda}$. Fix a $n<p_{*}<2n$ and $n/2<r<n$ so that (2.7) with $p=p_{*}$. Let us construct the solution to (1.2) of the form $U_{\lambda}=V_{\lambda}+W_{\lambda}$. To this end, we consider the stationary perturbed system
$\begin{cases}-\Delta W_{\lambda}+\mathbb{P}\operatorname{div}\left(V_{\lambda}\otimes W_{\lambda}+W_{\lambda}\otimes V_{\lambda}+W_{\lambda}\otimes W_{\lambda}\right)=-F,\quad&x\in\mathbb{R}^{n},\\ \operatorname{div}W_{\lambda}=0,\quad&x\in\mathbb{R}^{n},\end{cases}$ (5.1)
and define an operator $\mathcal{X}$ on $\dot{B}_{r,1}^{n/r-1}$ by
$\mathcal{X}[W]:=-(-\Delta)^{-1}F-(-\Delta)^{-1}\mathbb{P}\operatorname{div}\left(V_{\lambda}\otimes W+W\otimes V_{\lambda}+W\otimes W\right).$ (5.2)
Then, it follows from Lemma 2.1 that
$\begin{aligned} {\left\|\mathcal{X}[W]\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}&\leqslant C{\left\|F\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-3}}+C\left({\left\|V_{\lambda}\right\|}_{\dot{B}_{p_{*},1}^{\frac{n}{p_{*}}-1}}+{\left\|W\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\right){\left\|W\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}},\\ &\leqslant C_{1}\lambda^{-\beta}+C_{1}\left(\lambda^{-\alpha(p_{*})}+{\left\|W\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\right){\left\|W\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}},\end{aligned}$ (5.3)
${\left\|\mathcal{X}[W_{1}]-\mathcal{X}[W_{2}]\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}$ (5.4)
$\quad\leqslant C\left({\left\|V_{\lambda}\right\|}_{\dot{B}_{p_{*},1}^{\frac{n}{p_{*}}-1}}+{\left\|W_{1}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}+{\left\|W_{2}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\right){\left\|W_{1}-W_{2}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}$ (5.5)
$\quad\leqslant C_{1}\left(\lambda^{-\alpha(p_{*})}+{\left\|W_{1}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}+{\left\|W_{2}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\right){\left\|W_{1}-W_{2}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}$ (5.6)
for all $W,W_{1},W_{2}\in\dot{B}_{r,1}^{n/r-1}(\mathbb{R}^{n})$ with some positive constant $C_{1}=C_{1}(n,p_{*},r)$. Set
$\mathscr{X}:=\left\{W\in\dot{B}_{r,1}^{n/r-1}(\mathbb{R}^{n})\ ;\ {\left\|W\right\|}_{\dot{B}_{r,1}^{n/r-1}}\leqslant 2C_{1}\lambda^{-\beta}\right\}.$ (5.7)
Then, choosing $\lambda$ so large that $2C_{1}\left(\lambda^{-\alpha(p_{*})}+4C_{1}\lambda^{-\beta}\right)\leqslant 1$, we have
${\left\|\mathcal{X}[W]\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\leqslant C_{1}\lambda^{-\beta},\qquad{\left\|\mathcal{X}[W_{1}]-\mathcal{X}[W_{2}]\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\leqslant\frac{1}{2}{\left\|W_{1}-W_{2}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}$ (5.8)
for all $W,W_{1},W_{2}\in\mathscr{X}$. Hence, the contraction mapping principle yields the existence of a unique solution $W_{\lambda}\in\mathscr{X}$ to (5.1).
Let $U_{\lambda}:=V_{\lambda}+W_{\lambda}$. Then, we see that $U_{\lambda}$ solves (1.2). Moreover, it follows from Theorem 4.1, $W_{\lambda}\in\mathscr{X}$, and $\alpha(p)<\beta$ that
${\left\|U_{\lambda}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\leqslant{\left\|V_{\lambda}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}+C{\left\|W_{\lambda}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\leqslant C\left(\lambda^{-\alpha(p)}+\lambda^{-\beta}\right)\leqslant C\lambda^{-\alpha(p)}$ (5.9)
and
${\left\|U_{\lambda}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\geqslant{\left\|V_{\lambda}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}-C{\left\|W_{\lambda}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\geqslant C^{-1}\lambda^{-\alpha(p)}-C\lambda^{-\beta}\geqslant C^{-1}\lambda^{-\alpha(p)}$ (5.10)
for sufficiently large $\lambda$. Thus, $U_{\lambda}$ is a desired steady state.
Step 2. Non-stationary Navier–Stokes flow around $U_{\lambda}$. Taking the smallness parameter $\eta=0$ for the case of $p\geqslant 2n$ or $q=\infty$, and recalling the continuous embedding $\dot{B}_{p_{1},q_{1}}^{n/p_{1}-1}(\mathbb{R}^{n})\hookrightarrow\dot{B}_{p_{2},q_{2}}^{n/p_{2}-1}(\mathbb{R}^{n})$ ($1\leqslant p_{1}\leqslant p_{2}\leqslant\infty$, $1\leqslant q_{1}\leqslant q_{2}\leqslant\infty$), we see that it suffices to consider only the case of $p<2n$ and $q<\infty$. Then, we may choose a $\theta$ satisfying (2.12). Let us consider the perturbed system of (1.1) around $U_{\lambda}$.
$\begin{cases}\partial_{t}w_{\lambda}-\Delta w_{\lambda}+\mathbb{P}\operatorname{div}(U_{\lambda}\otimes w_{\lambda}+U_{\lambda}\otimes w_{\lambda}+w_{\lambda}\otimes w_{\lambda})=0,\\ \operatorname{div}w_{\lambda}=0,\\ w_{\lambda}(0,x)=w_{0}:=u_{0}-U_{\lambda}.\end{cases}$ (5.11)
To construct the perturbed solution, we define a map
$\mathcal{Y}[w](t):=e^{t\Delta}w_{0}-\int_{0}^{t}e^{(t-\tau)\Delta}\mathbb{P}\operatorname{div}\left(U_{\lambda}\otimes w(\tau)+w(\tau)\otimes U_{\lambda}+w(\tau)\otimes w(\tau)\right)d\tau.$ (5.12)
Then, we see by Lemma 2.2 that
${\left\|\mathcal{Y}[w]\right\|}_{\widetilde{L^{\infty}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1})\cap\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}$ (5.13)
$\quad\leqslant{}C{\left\|w_{0}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}+C\left({\left\|w_{0}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}+{\left\|w\right\|}_{\widetilde{L^{\infty}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1})}\right){\left\|w\right\|}_{\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}$ (5.14)
$\quad\leqslant{}C_{2}{\left\|U_{\lambda}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}+C_{2}\left(\lambda^{-\alpha(p)}+{\left\|w\right\|}_{\widetilde{L^{\infty}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1})}\right){\left\|w\right\|}_{\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}$ (5.15)
and
${\left\|\mathcal{Y}[w_{1}]-\mathcal{Y}[w_{2}]\right\|}_{\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}$ (5.16)
$\quad\leqslant{}C\left({\left\|U_{\lambda}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}+\sum_{m=1}^{2}{\left\|w_{m}\right\|}_{\widetilde{L^{\infty}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1})}\right){\left\|w_{1}-w_{2}\right\|}_{\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}$ (5.17)
$\quad\leqslant{}C_{2}\left(\lambda^{-\alpha(p)}+\sum_{m=1}^{2}{\left\|w_{m}\right\|}_{\widetilde{L^{\infty}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1})}\right){\left\|w_{1}-w_{2}\right\|}_{\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}$ (5.18)
for all $w,w_{1},w_{2}\in\widetilde{C}([0,\infty);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))\cap\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{n/p-1+2/\theta}(\mathbb{R}^{n}))$ with some positive constant $C_{2}=C_{2}(n,p,r,\theta)$. We define the solution space $\mathscr{Y}$ by
$\mathscr{Y}:=\left\{\begin{aligned} w\in{}&\widetilde{C}([0,\infty);\dot{B}_{p,q}^{\frac{n}{p}-1}(\mathbb{R}^{n}))\cap\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}}(\mathbb{R}^{n}))\ ;\\ &{\left\|w\right\|}_{\widetilde{L^{\infty}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1})\cap\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}\leqslant 2C_{2}{\left\|w_{0}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\end{aligned}\right\}.$ (5.19)
Then, we see that $\mathcal{Y}$ is a contraction mapping on $\mathscr{Y}$, by choosing $\lambda$ and ${\left\|w_{0}\right\|}_{\dot{B}_{p,q}^{n/p-1}}$ sufficiently large and small, respectively. Hence, the Banach fixed point theorem implies that there exists a unique mild solution $w_{\lambda}\in\mathscr{Y}$ to the perturbed system (5.11). We then obtain the non-stationary solution to (1.1) by $u_{\lambda}:=U_{\lambda}+w_{\lambda}$.
Finally, we shall prove the asymptotic behavior $u_{\lambda}(t)-U_{\lambda}=w_{\lambda}(t)\to 0$ strongly in $\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ as $t\to\infty$. To see this, we fix a $T>0$ and rewrite the integral equation for $w_{\lambda}$ as
$w_{\lambda}(t)={}$ $e^{(t-T)\Delta}w_{\lambda}(T)$ (5.20)
$-\int_{T}^{t}e^{(t-\tau)\Delta}\mathbb{P}\operatorname{div}\left(U_{\lambda}\otimes w_{\lambda}(\tau)+w_{\lambda}(\tau)\otimes U_{\lambda}+w_{\lambda}(\tau)\otimes w_{\lambda}(\tau)\right)d\tau$ (5.21)
for $t>T$. Then, by Lemma 2.2, we have for $T^{\prime}>T$ that
${\left\|w_{\lambda}\right\|}_{L^{\infty}(T^{\prime},\infty;\dot{B}_{p,q}^{\frac{n}{p}-1})}\leqslant{}$ ${\left\|w_{\lambda}\right\|}_{\widetilde{L^{\infty}}(T^{\prime},\infty;\dot{B}_{p,q}^{\frac{n}{p}-1})}$ (5.22)
$\leqslant{}$ ${\left\|e^{(T^{\prime}-T)\Delta}w_{\lambda}(T)\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}$ (5.23)
$+C\left({\left\|U_{\lambda}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}+{\left\|w_{\lambda}\right\|}_{\widetilde{L^{\infty}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1})}\right){\left\|w_{\lambda}\right\|}_{\widetilde{L^{\theta}}(T,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}$ (5.24)
$\leqslant{}$ $C{\left\|e^{(T^{\prime}-T)\Delta}w_{\lambda}(T)\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}+C{\left\|w_{\lambda}\right\|}_{\widetilde{L^{\theta}}(T,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})},$ (5.25)
which implies
$\limsup_{T^{\prime}\to\infty}{\left\|w_{\lambda}\right\|}_{L^{\infty}(T^{\prime},\infty;\dot{B}_{p,q}^{\frac{n}{p}-1})}\leqslant C{\left\|w_{\lambda}\right\|}_{\widetilde{L^{\theta}}(T,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}.$ (5.26)
Hence, letting $T\to\infty$, we complete the proof. ∎
Acknowledgments.
The author was supported by JSPS KAKENHI, Grant Number JP25K17279.
Conflict of interest.
The author has declared no conflicts of interest.
Data availability.
Data sharing not applicable to this article as no dataset was generated or analyzed during the current study.

### v2

@@
-Sharp non-uniqueness for the Navier–Stokes equations in scaling critical spaces
+arXiv:2602.19846v2[math.AP] 02 Mar 2026
 Abstract.
@@
 where $u=u(t,x):[0,\infty)\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ is the unknown velocity field of the fluid, whereas $u_{0}=u_{0}(x):\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a given initial datum satisfying $\operatorname{div}u_{0}=0$.
-We denote by $\mathbb{P}={\rm Id}+\nabla(-\Delta)^{-1}\operatorname{div}$ the Helmholtz projection.
+We denote by $\mathbb{P}={\rm Id}+\nabla\operatorname{div}(-\Delta)^{-1}$ the Helmholtz projection onto divergence-free vector fields.
 In dimensions $n\geqslant 3$, Lions–Masmoudi [Lio-Mas] and Furioli–Rieusset–Terraneo [Fur-Rie-Ter-2000] proved that mild solutions to (1.1) are unique in the critical space $C([0,T);L^{n}(\mathbb{R}^{n}))$.
-The aim of this paper is to prove the sharpness of the choice of $L^{n}(\mathbb{R}^{n})$ in the sense that the uniqueness fails when one replaces $L^{n}(\mathbb{R}^{n})$ by a scaling critical space that is even slightly larger; for instance, relaxing $L^{n}(\mathbb{R}^{n})=\dot{F}_{n,2}^{0}(\mathbb{R}^{n})$ to a slightly larger Lizorkin–Triebel space $\dot{F}_{n,q}^{0}(\mathbb{R}^{n})$ with any $q>2$ leads to non-uniqueness of the mild solutions in $C([0,T);\dot{F}_{n,q}^{0}(\mathbb{R}^{n}))$.
+The aim of this paper is to prove the sharpness of the choice of $L^{n}(\mathbb{R}^{n})$ in the sense that the uniqueness fails as one replaces $L^{n}(\mathbb{R}^{n})$ by a scaling critical space that is even slightly larger; for instance, enlarging $L^{n}(\mathbb{R}^{n})=\dot{F}_{n,2}^{0}(\mathbb{R}^{n})$ to a slightly larger Lizorkin–Triebel space $\dot{F}_{n,q}^{0}(\mathbb{R}^{n})$ with any $q>2$ leads to non-uniqueness of the mild solutions in $C([0,T);\dot{F}_{n,q}^{0}(\mathbb{R}^{n}))$.
 To achieve this, we show that the necessary and sufficient condition for the uniqueness of mild solutions to (1.1) in the scaling critical Besov spaces is given by $1\leqslant p<n$ with $1\leqslant q\leqslant\infty$, or $p=n$ with $1\leqslant q\leqslant 2$; see Figure 1 below.
-Moreover, for the non-unique case $p=n$ with $2<q\leqslant\infty$, or $n<p\leqslant\infty$ with $1\leqslant q\leqslant\infty$, we prove that the zero initial datum $u_{0}=0$ generates infinitely many small global solutions $u\in C([0,\infty);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))$, which asymptotically converge to non-trivial small steady states $U\in\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ solving the unforced stationary Navier–Stokes equations:
+Moreover, for the non-unique case $p=n$ with $2<q\leqslant\infty$, or $n<p\leqslant\infty$ with $1\leqslant q\leqslant\infty$, we prove that the zero initial datum $u_{0}=0$ generates infinitely many small global solutions $u\in C([0,\infty);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))$, which asymptotically converge to a non-trivial flow $U\in\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ solving the unforced stationary Navier–Stokes equations:
 $\begin{cases}-\Delta U+\mathbb{P}\operatorname{div}(U\otimes U)=0,\qquad&x\in\mathbb{R}^{n},\\ \operatorname{div}U=0,\qquad&x\in\mathbb{R}^{n}.\end{cases}$ (1.2)
 1.1. Known results related to our work
-Before we state our main results precisely, we recall previous studies related to our work and mention the place of our work.
-By the Duhamel principle, we may formally rewrite (1.1) as the integral equation
-$u(t)=e^{t\Delta}u_{0}-\int_{0}^{t}e^{(t-\tau)\Delta}\mathbb{P}\operatorname{div}(u(\tau)\otimes u(\tau))d\tau,$ (1.3)
-A solution of (1.3) is called a mild solution.
-It is natural to construct mild solutions in scaling critical spaces.
+Before we state our main results precisely, we recall the previous studies related to our work. By the Duhamel principle, we may formally rewrite (1.1) as the integral equation
+$u(t)=e^{t\Delta}u_{0}-\int_{0}^{t}e^{(t-\tau)\Delta}\mathbb{P}\operatorname{div}(u(\tau)\otimes u(\tau))d\tau.$ (1.3)
+A solution to (1.3) is called a mild solution of (1.1).
+In view of the Fujita–Kato principle, it is natural to construct mild solutions in scaling critical spaces.
 Here, let us recall the notion of scaling criticality.
@@
 A Banach space $X$ is called a scaling critical space if ${\left\|u_{\lambda}(0,\cdot)\right\|}_{X}={\left\|u(0,\cdot)\right\|}_{X}$ for all $\lambda>0$.
-Typical examples of scaling critical spaces are as follows
-$\dot{H}^{\frac{n}{2}-1}(\mathbb{R}^{n})\hookrightarrow L^{n}(\mathbb{R}^{n})\hookrightarrow L^{n,\infty}(\mathbb{R}^{n}),$ (1.4)
-$\dot{B}_{p,q}^{\frac{n}{p}-1}(\mathbb{R}^{n})\hookrightarrow{\rm BMO}^{-1}(\mathbb{R}^{n})\qquad(1\leqslant p<\infty,\ 1\leqslant q\leqslant\infty).$ (1.5)
+Typical examples of scaling-critical spaces and their inclusion relations are as follows:
+$\dot{H}^{\frac{n}{2}-1}(\mathbb{R}^{n})\hookrightarrow L^{n}(\mathbb{R}^{n})$ $\hookrightarrow L^{n,\infty}(\mathbb{R}^{n})$ (1.4)
+$\hookrightarrow\dot{B}_{p,\infty}^{\frac{n}{p}-1}(\mathbb{R}^{n})\hookrightarrow{\rm BMO}^{-1}(\mathbb{R}^{n})\hookrightarrow\dot{B}_{\infty,\infty}^{-1}(\mathbb{R}^{n})$ (1.5)
+with $n<p<\infty$ and
+$\dot{B}_{p,q}^{\frac{n}{p}-1}(\mathbb{R}^{n})\hookrightarrow{\rm BMO}^{-1}(\mathbb{R}^{n})$ (1.6)
+with $1\leqslant p<\infty$ and $1\leqslant q\leqslant\infty$.
 This line of research originates from the pioneering work of Fujita–Kato [Fuj-Kato-64-ARMA], where they proved well-posedness in $\dot{H}^{n/2-1}(\mathbb{R}^{n})$, and was extended to $L^{n}(\mathbb{R}^{n})$ by Kato [Kato-84] and Giga–Miyakawa [Gig-Miy-85].
@@
 In the endpoint case $p=\infty$, Koch–Tataru [Koh-Tat-01] established well-posedness for small data in ${\rm BMO}^{-1}(\mathbb{R}^{n})$, however it was shown by [Bou-Pav-08, Yon-10, Wan-15] that (1.1) is ill-posed in $\dot{B}_{\infty,q}^{-1}(\mathbb{R}^{n})$ ($1\leqslant q\leqslant\infty$).
+See [IN] for more precise analysis on the well-posedness of (1.1) in critical Besov or Lizorkin–Triebel spaces.
 We refer to [Kozono-Shimizu-2018, Takeuchi-2025-SIMA, Yam-2000, arXiv:2509.21272v2] for the forced Navier–Stokes flow in critical spaces.
 Next, we address the issue of uniqueness for the Navier–Stokes equations.
-This topic was brought to the forefront by Leray’s pioneering work [Leray-1934-Acta], which produced weak solutions for any $L^{2}$ initial datum but did not settle their uniqueness.
+This topic was brought to the forefront by Leray’s pioneering work [Leray-1934-Acta], which produced weak solutions for any divergence-free initial datum in $L^{2}(\mathbb{R}^{3})$ but did not reveal their uniqueness.
 It is well-known that the Leray–Hopf weak solutions are unique if they belong to the Ladyzhenskaya–Prodi–Serrin class $L^{p}(0,T;L^{q}(\mathbb{R}^{n}))$ with $n/q+2/p=1$; see [Serrin-1963, Masuda-1984, Escauriaza-Seregin-Sverak-2003, Kozono-Sohr-1996] for more detail.
-For the mild solutions, Lions–Masmoudi [Lio-Mas] and Furioli–Rieusset–Terraneo [Fur-Rie-Ter-2000] proved the uniqueness in the scaling critical class $C([0,T);L^{n}(\mathbb{R}^{n}))$ with $n\geqslant 3$, which corresponds to an endpoint $(p,q)=(\infty,n)$ of the Ladyzhenskaya–Prodi–Serrin class.
-Note that [Fur-Rie-Ter-2000] considered not only in $L^{n}(\mathbb{R}^{n})$ but also other critical spaces such as Morrey spaces.
+For the mild solutions, Lions–Masmoudi [Lio-Mas] and Furioli–Rieusset–Terraneo [Fur-Rie-Ter-2000] proved uniqueness in the scaling critical class $C([0,T);L^{n}(\mathbb{R}^{n}))$ with $n\geqslant 3$, which corresponds to an endpoint $(p,q)=(\infty,n)$ of the Ladyzhenskaya–Prodi–Serrin class.
 See also [Yam-2000] for the uniqueness of small mild solutions in $L^{\infty}(0,T;L^{n,\infty}(\mathbb{R}^{n}))$ and [Miu-2005] for the uniqueness in $C([0,T);bmo^{-1}(\mathbb{R}^{n}))\cap L^{\infty}_{\rm loc}((0,T);L^{\infty}(\mathbb{R}^{n}))$.
-The recent progress of this direction is given by Zhan [arXiv:2402.01174v2], where he considered the uniqueness of mild solutions in $C([0,T);\widetilde{L}^{n,\infty}(\mathbb{R}^{n}))$ with a new space $\widetilde{L}^{n,\infty}(\mathbb{R}^{n})$ which is larger than $L^{n}(\mathbb{R}^{n})$.
-In terms of the Besov spaces, the following proposition holds.
+The recent progress of this direction is given by Zhan [arXiv:2402.01174v2], where he considered the uniqueness of mild solutions in $C([0,T);\widetilde{L}^{n,\infty}(\mathbb{R}^{n}))$ with a new space $\widetilde{L}^{n,\infty}(\mathbb{R}^{n})=\overline{L^{n,\infty}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})}^{L^{n,\infty}}$ which is strictly larger than the completion of $C_{c}^{\infty}(\mathbb{R}^{n})$ in $L^{n,\infty}(\mathbb{R}^{n})$.
+In terms of the scaling critical Besov spaces, the following proposition holds.
 Proposition 1.1.
@@
 Then, the uniqueness of mild solutions to (1.1) holds in $C([0,T);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))$, provided a smallness condition on solutions only when $q=\infty$.
-The proof of Proposition 1.1 follows immediately by adopting the argument in [Fur-Rie-Ter-2000, Fuj-AIPHC, Iwabuchi-Okazaki-2025] in the case of (U1) and by the result of [Fur-Rie-Ter-2000] with the continuous embedding $\dot{B}_{n,q}^{0}(\mathbb{R}^{n})\hookrightarrow L^{n}(\mathbb{R}^{n})$ for the case of (U2).
+The proof of Proposition 1.1 follows immediately by adopting the argument in [Fur-Rie-Ter-2000, Fuj-AIPHC, Iwabuchi-Okazaki-2025] in the case of (U1) and by the result of [Fur-Rie-Ter-2000] with the continuous embedding $\dot{B}_{n,q}^{0}(\mathbb{R}^{n})\hookrightarrow L^{n}(\mathbb{R}^{n})$ holds for the case of (U2).
 Note that Proposition 1.1 still holds if one replaces Besov spaces by Lizorkin–Triebel spaces as the continuous embedding $\dot{F}_{p,q}^{n/p-1}(\mathbb{R}^{n})\hookrightarrow L^{n}(\mathbb{R}^{n})$ for all $(p,q)$ satisfying (U1) or (U2).
 Finally, we briefly review works on non-uniqueness.
-One major direction originates from applying Nash’s convex integration to the 3D Euler equations on the torus (see, [DeLellis-Szekelyhidi-09, Isett-18, Daneri-Szekelyhidi-2017-ARMA, Lellis-Szekelyhidi-2013-Invent, Buckmaster-Lellis-Isett-Szekelyhidi-2015-AnnMath] for instance).
-Using an iterative scheme based on the Euler–Reynolds system, these works constructed Euler flows violating the energy conservation law and made substantial progress toward the Onsager conjecture.
+One major direction originates from applying the Nash convex integration to the Euler equations on the torus (see, [DeLellis-Szekelyhidi-09, Isett-18, Daneri-Szekelyhidi-2017-ARMA, Lellis-Szekelyhidi-2013-Invent, Buckmaster-Lellis-Isett-Szekelyhidi-2015-AnnMath] for instance).
+Using an iterative scheme based on the Euler–Reynolds system via the geometric lemma, these works constructed Euler flows violating the energy conservation law and made substantial progress toward the Onsager conjecture.
 Although incorporating the viscous term was long thought to be fundamentally difficult, Buckmaster–Vicol [Buc-Vic-19] introduced the concept of intermittent convex integration and proved that uniqueness of weak solutions to the Navier–Stokes equations on $\mathbb{T}^{3}$ breaks down in the supercritical class $C([0,T);H^{\beta}(\mathbb{T}^{3}))$ for some $0<\beta\ll 1/2$.
-Cheskidov–Luo [Che-Luo-21] showed that an endpoint Ladyzhenskaya–Prodi–Serrin class $L^{2}(0,T;L^{\infty}(\mathbb{T}^{n}))$, is optimal with respect to the time integrability exponent: they constructed non-unique solutions in $L^{p}(0,T;L^{\infty}(\mathbb{T}^{3}))$ for $p<2$.
-They also considered another endpoint phenomenon in two dimensions and proved non-uniqueness in $C([0,T);L^{p}(\mathbb{T}^{2}))$ for $p<2$ (see [Che-Luo-22-AnnPDE]).
-Furthermore, Cheskidov–Zeng–Zhang [arXiv:2503.05692] proved that every initial datum in $H^{1/2}(\mathbb{T}^{3})$ generates infinitely many Navier–Stokes flows whose kinetic energy is continuous and monotonically decreasing in time.
+Cheskidov–Luo [Che-Luo-21] showed that an endpoint Ladyzhenskaya–Prodi–Serrin class $L^{2}(0,T;L^{\infty}(\mathbb{T}^{n}))$, is optimal with respect to the time integrability exponent; they constructed non-unique solutions in $L^{p}(0,T;L^{\infty}(\mathbb{T}^{3}))$ for $p<2$.
+In [Che-Luo-22-AnnPDE], they also considered another endpoint in two dimensions and proved non-uniqueness in $C([0,T);L^{p}(\mathbb{T}^{2}))$ for $p<2$.
+Furthermore, Cheskidov–Zeng–Zhang [arXiv:2503.05692] claimed that every initial datum in $H^{1/2}(\mathbb{T}^{3})$ generates infinitely many Navier–Stokes flows whose kinetic energy is continuous and monotonically decreasing in time.
 Remarkably, Coiculescu–Palasek [Coi-Pal-25] showed that even in ${\rm BMO}^{-1}(\mathbb{T}^{3})$, where well-posedness for small data is known, a certain large initial datum generates two smooth solutions.
 Their argument builds on ideas developed by Palasek [Pal-25-IMRN] for the dyadic Navier–Stokes model.
-Pushing this direction further, Cheskidov–Dai–Palasek [arXiv:2511.09556] observed that non-uniqueness in the form of instantaneous Type I blow-up may be generated from arbitrary $C^{\infty}$ solenoidal initial data.
-Another strategy to prove non-uniqueness, developed by Jia–Sverák [Jia-Sve-Invent, Jia-Sve-JFA], is to exploit the instability of certain self-similar solutions.
+Pushing this direction further, Cheskidov–Dai–Palasek [arXiv:2511.09556] observed that non-uniqueness, in the form of instantaneous Type I blow-up, may be generated from arbitrary $C^{\infty}$ solenoidal initial data.
+Another strategy to prove non-uniqueness is developed by Jia–Sverák [Jia-Sve-Invent, Jia-Sve-JFA], and is to exploit the instability of certain self-similar solutions.
 A notable result in this direction is due to Albritton–Brué–Colombo [Alb-Bru-Col-AnnMath], who proved non-uniqueness of Leray–Hopf weak solutions to the forced Navier–Stokes equations.
 See Aoki–Maekawa [Aok-Mae] for the corresponding result on a two-dimensional half space.
-More recently, Hou–Wang–Yang [2509.25116v1] obtained an analogous statement in the unforced case $f\equiv 0$ by a computer-assisted proof.
+More recently, Hou–Wang–Yang [2509.25116v1] announced an analogous statement in three-dimension with the unforced case $f\equiv 0$ by a computer-assisted proof.
 1.2. Main results
-The purpose of this paper is to show non-uniqueness of mild solutions to (1.1) in $C([0,T);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))$ for all $(p,q)$ satisfying neither (U1) nor (U2).
+The purpose of this paper is to show non-uniqueness of mild solutions to (1.1) in $C([0,T);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))$ for all $(p,q)$ satisfying neither (U1) nor (U2) in Proposition 1.1.
 Now, the precise statement of our main theorem reads as follows.
@@
   • (N1)
-    $p=n$ and $2<q\leqslant\infty$
+    $p=n$ and $2<q\leqslant\infty$,
   • (N2)
     $n<p\leqslant\infty$ and $1\leqslant q\leqslant\infty$.
-Then, there exist constants $\varepsilon_{0}=\varepsilon_{0}(n,p,q)>0$ and $\eta=\eta(n,p,q)\geqslant 0$ such that for any $0<\varepsilon\leqslant\varepsilon_{0}$ and solenoidal initial datum $u_{0}\in\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ with ${\left\|u_{0}\right\|}_{\dot{B}_{p,q}^{n/p-1}}\leqslant\eta$, (1.1) possesses a global mild solution $u_{\varepsilon}\in C([0,\infty);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))$ satisfying
-$\lim_{t\to\infty}{\left\|u_{\varepsilon}(t)-U_{\varepsilon}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}=0,$ (1.6)
-where $U_{\varepsilon}\in\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ solves (1.2) and admits the estimate
-$C^{-1}\varepsilon\leqslant{\left\|U_{\varepsilon}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\leqslant C\varepsilon$ (1.7)
-with some positive constant $C=C(n,p,q)$ independent of $u_{0}$ and $\varepsilon$.
+Then, there exist a monotone decreasing sequence $\{\varepsilon_{m}\}_{m\in\mathbb{N}}\subset(0,1)$ with $\varepsilon_{m}\to 0$ and a constant $\eta=\eta(n,p,q)\geqslant 0$ such that for any $m\in\mathbb{N}$ and initial datum $u_{0}\in\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ with
+${\left\|u_{0}\right\|}_{\dot{B}_{p,q}^{n/p-1}}\leqslant\eta,\qquad\operatorname{div}u_{0}=0,$ (1.7)
+there exist a non-stationary global mild solution $u_{m}\in C([0,\infty);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))$ to (1.1) and a stationary solution $U_{m}\in\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ to (1.2) satisfying
+$\lim_{t\to\infty}{\left\|u_{m}(t)-U_{m}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}=0,\qquad C^{-1}\varepsilon_{m}\leqslant{\left\|U_{m}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\leqslant C\varepsilon_{m}$ (1.8)
+with some constant $C=C(n,p,q)>0$ independent of $u_{0}$ and $m$.
 Remark 1.3.
@@
     $C([0,T);\dot{B}_{p,q}^{\frac{n}{p}-1}(\mathbb{R}^{n}))\cap\widetilde{L^{1}}(0,T;\dot{B}_{p,q}^{\frac{n}{p}+1}(\mathbb{R}^{n})).$ (1.9)
-    whereas our non-unique solutions do not satisfy the auxiliary dissipative condition $\widetilde{L^{1}}(0,T;\dot{B}_{p,q}^{n/p+1}(\mathbb{R}^{n}))$.
+    whereas our non-unique solutions do not belong to the auxiliary dissipative space $\widetilde{L^{1}}(0,T;\dot{B}_{p,q}^{n/p+1}(\mathbb{R}^{n}))$.
+[^1] See Section 2 for the definition of Chemin–Lerner spaces $\widetilde{L^{\theta}}(0,T;\dot{B}_{p,q}^{s}(\mathbb{R}^{n}))$.
   2. 2.
-    The large-time behavior of global solutions to (1.1) with small initial data has been studied extensively; see [Pla-98, Nakasato-25-JDE, Gallagher-Iftimie-Planchon-2002, Gallagher-Iftimie-Planchon-2003, Takeuchi-2025, Watanabe-2025, Benameu-2015-JMAA] for instance.
+    The large-time behavior of global solutions to (1.1) with small initial data in critical Besov spaces has been studied extensively; see [Pla-98, Nakasato-25-JDE, Gallagher-Iftimie-Planchon-2002, Gallagher-Iftimie-Planchon-2003, Takeuchi-2025, Watanabe-2025, Benameu-2015-JMAA] for instance.
 However, all of these results exploit the viscous dissipation and address only solutions that decay and converge asymptotically to zero.
-In contrast, the non-decaying global solutions $u_{\varepsilon}$ constructed in this paper are the first examples [^1] In the framework of weak solutions, the existence of non-dissipative solutions is widely known; see [Buc-Vic-19].
-of non-dissipative unforced Navier–Stokes flows in critical spaces where the Stokes semigroup is strongly continuous.
-[^2] If this is not the case, solutions may not decay even at the linear level.
+In contrast, the non-decaying global solutions $u_{\varepsilon}$ constructed in this paper are the first examples [^2] In the framework of weak solutions, the existence of non-dissipative solutions is widely known; see [Buc-Vic-19] for instance.
+of non-dissipative unforced Navier–Stokes flows in critical spaces in which the Stokes semigroup decays at temporal infinity.
   3. 3.
@@
     hold with some $\delta=\delta(p,q)\in(0,1)$.
-Thus, non-uniqueness for $u_{0}=0$ remains valid after replacing the Besov spaces by the corresponding Lizorkin–Triebel spaces.
-In particular, we see that enlarging $L^{n}(\mathbb{R}^{n})=\dot{F}_{n,2}^{0}(\mathbb{R}^{n})$ in [Lio-Mas, Fur-Rie-Ter-2000] even slightly to $\dot{F}_{n,q}^{0}(\mathbb{R}^{n})$ with $q>2$ instantly leads to non-uniqueness.
-Moreover, we obtain the breakdown of unconditional uniqueness for small solutions in $L^{\infty}(0,T;{\rm BMO}^{-1}(\mathbb{R}^{n}))$ since ${\rm BMO}^{-1}(\mathbb{R}^{n})=\dot{F}_{\infty,2}^{-1}(\mathbb{R}^{n})$.
+Thus, non-uniqueness for $u_{0}=0$ remains valid if one replaces the Besov spaces by the corresponding Lizorkin–Triebel spaces.
+In particular, we see that enlarging $L^{n}(\mathbb{R}^{n})=\dot{F}_{n,2}^{0}(\mathbb{R}^{n})$ in [Lio-Mas, Fur-Rie-Ter-2000] even slightly to $\dot{F}_{n,q}^{0}(\mathbb{R}^{n})$ with $q>2$ leads to non-uniqueness.
+Moreover, we obtain the breakdown of unconditional uniqueness for small mild solutions in $L^{\infty}(0,T;{\rm BMO}^{-1}(\mathbb{R}^{n}))$ since ${\rm BMO}^{-1}(\mathbb{R}^{n})=\dot{F}_{\infty,2}^{-1}(\mathbb{R}^{n})$.
 This is different from the phenomenon in [Coi-Pal-25], which concerns non-uniqueness of smooth solutions for large data in ${\rm BMO}^{-1}(\mathbb{T}^{3})$.
-See Figure 1.
+See Figure 1 below.
   4. 4.
@@
   5. 5.
-    Concerning non-uniqueness for steady flows, Luo [Luo-ARMA-2019] constructed nontrivial solutions belonging to $L^{2}(\mathbb{T}^{n})$ with $n\geqslant 4$.
+    Concerning non-uniqueness for steady Navier–Stokes flows, Luo [Luo-ARMA-2019] constructed nontrivial solutions to (1.2) belonging to $L^{2}(\mathbb{T}^{n})$ with $n\geqslant 4$.
 See [Rieusset-25-JFA] for two-dimensional case in $H^{-1}(\mathbb{T}^{2})$.
 Note that both spaces are supercritical and the three-dimensional case has been open.
-Our result appears to be the first to establish non-uniqueness of steady solutions in scaling critical spaces including three-dimensional case.
+Our result appears to be the first to establish non-uniqueness of steady solutions in scaling critical spaces.
+Moreover our result includes the three-dimensional case.
 Meanwhile, our method cannot be implemented in two dimensions.
@@
   6. 6.
-    As will be explained later, our construction of solutions relies on the structure provided by the classical Nash lemma, whereas we do not employ any of the periodic building blocks commonly used in convex-integration approaches to fluid equations, such as Beltrami or Mikado flows.
+    As will be explained later, our construction of solutions relies on the structure provided by the classical Nash lemma, whereas we do not employ any of the periodic building blocks commonly used in convex integration approaches to fluid equations, such as Beltrami or Mikado flows.
+  7. 7.
+    Our proof works in essentially the same way if $\mathbb{R}^{n}$ is replaced by $\mathbb{T}^{n}$; see Remark 4.5.
 […]
@@
 1.3. Outline of the proof
-Let us explain the strategy of the proof of Theorem 1.2.
-The strategy is that we first build a dominant part of the solution and then solve the remaining perturbative problem by a fixed point argument in a function space where well-posedness is available.
-More precisely, we construct an almost frequency-localized bulding blocks $\{V_{j}\}_{j\geqslant 0}$ of stationary flows, each essentially supported on the $j$th frequency shell, and design the principal profile
-$V=\sum_{j=0}^{\infty}V_{j}\neq 0\qquad{\rm in}\quad\dot{B}_{p,q}^{\frac{n}{p}-1}(\mathbb{R}^{n})$ (1.10)
+Let us explain the idea of the proof of Theorem 1.2.
+The strategy is that we first establish a dominant part of the solution which lies in $\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ with $(p,q)$ satisfying (N1) or (N2), and then solve the remaining perturbative problem by a fixed point argument in a function space where well-posedness is available.
+More precisely, we construct an almost frequency-localized building blocks $\{V_{j}\}_{j\geqslant 0}$ of stationary flows, each of which are essentially built on the $j$th frequency shell, and design the principal profile
+$V=\sum_{j=0}^{\infty}V_{j}\in\dot{B}_{p,q}^{\frac{n}{p}-1}(\mathbb{R}^{n})\setminus\{0\}$ (1.11)
 so that the associated residual forcing
-$F:=-\Delta V+\mathbb{P}\operatorname{div}(V\otimes V),$ (1.11)
+$F:=-\Delta V+\mathbb{P}\operatorname{div}(V\otimes V)$ (1.12)
 belongs to well-posedness class $\dot{B}_{r,1}^{n/r-3}(\mathbb{R}^{n})$ ($r<n$) and admits the following smallness hierarchy:
@@
 $W$ driven by $F$ can be constructed in $\dot{B}_{r,1}^{n/r-1}(\mathbb{R}^{n})$ by well-posedness theory, and we see that
-${\left\|W\right\|}_{\dot{B}_{r,1}^{n/r-1}}\lesssim{\left\|F\right\|}_{\dot{B}_{r,1}^{n/r-3}}\ll{\left\|V\right\|}_{\dot{B}_{p,q}^{n/p-1}}.$ (1.13)
-Hence, the desired steady state is obtained in the form $U=V+W$.
-Note that ${\left\|U\right\|}_{\dot{B}_{p,q}^{n/p-1}}\sim{\left\|V\right\|}_{\dot{B}_{p,q}^{n/p-1}}\neq 0$ holds.
+${\left\|W\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\leqslant C{\left\|F\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-3}}\ll{\left\|V\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}.$ (1.14)
+Hence, the desired steady state is obtained in the form $U=V+W$. Note that $U\neq 0$ since
+${\left\|U\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\geqslant{\left\|V\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}-C{\left\|W\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\geqslant\frac{1}{2}{\left\|V\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}>0.$ (1.15)
 Furthermore, to construct a non-stationary solution $u$ whose asymptotic profile is $U$, we decompose $u=U+w$.
@@
 However, such a single-scale construction is too rigid to capture the delicate dependence on the Besov summability index $q$, which is essential for observing the subtle non-uniqueness phenomena addressed in (N1).
-Instead, we distribute the $j$th building block over many internal scales.
-To this end, for $\lambda\gg 1$ (which corresponds to $\varepsilon^{-1/\alpha}$ with some $\alpha$) and a fixed large number $\mu\in\mathbb{N}$, we introduce a tetration-like sequence [^4] This is defined by $\lambda_{0}=\lambda$ and $\lambda_{j}=\lambda^{2^{\lambda_{j-1}^{4\mu}}}$; see (4.9).
+Instead, we distribute the $j$th building block over many frequency scales.
+To this end, for $\lambda\gg 1$ (which is a parameter corresponding to $\varepsilon^{-1/\alpha}$ with some $\alpha$) and a fixed large number $\mu\in\mathbb{N}$, we introduce a tetration-like sequence [^4] This is defined by $\lambda_{0}=\lambda$ and $\lambda_{j}=\lambda^{2^{\lambda_{j-1}^{4\mu}}}$; see (4.9).
 $\{\lambda_{j}\}_{j\geqslant 0}$ and define the $j$th level frequency index set by
 $\Lambda_{j}\coloneqq\Bigl\{\ell\in\mathbb{N}\ ;\ \mu\log_{\lambda}\lambda_{j-1}\leqslant\ell\leqslant\log_{\lambda}\lambda_{j}\Bigr\},$
-and take $V_{j}$ in the schematic form
+and take $V_{j}$ in the schematic form [^5] In the actual construction of $V_{j}$, we apply a mollifier and introduce several additional modifications to enforce the divergence-free condition, so that its form becomes somewhat more involved.
 $V_{j}(x)=\sum_{k\in\mathcal{K}}f_{k,j}(x)\sum_{\ell\in\Lambda_{j}}\frac{1}{\sqrt{h_{j}\ell}}\cos\left(\lambda^{\ell}a_{k}\cdot x\right)k,\qquad h_{j}:=\sum_{\ell\in\Lambda_{j}}\frac{1}{\ell}.$ (1.16)
-Here, the choice of $1/\sqrt{h_{j}\ell}$ is inspired by [Tsu-19-ARMA]. With this multi-scale design, the oscillation calculus in Section 3 below yields
+Here, the choice of $1/\sqrt{\ell}$ is inspired by [Yon-10, Tsu-19-ARMA]. With this multi-scale design, the oscillation calculus in Section 3 below yields
 ${\left\|V_{j}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\sim\sum_{k\in\mathcal{K}}\frac{{\left\|f_{k,j}\right\|}_{L^{p}}}{\sqrt{h_{j}}}\left(\sum_{j\in\Lambda_{j}}\frac{\lambda^{(\frac{n}{p}-1)q\ell}}{\ell^{\frac{q}{2}}}\right)^{\frac{1}{q}},$ (1.17)
@@
 $F=\sum_{j=1}^{\infty}\Bigl(-\Delta V_{j-1}+\mathbb{P}\operatorname{div}(V_{j}\otimes V_{j})\Bigr)+\sum_{j_{1}\neq j_{2}}\mathbb{P}\operatorname{div}(V_{j_{1}}\otimes V_{j_{2}})=:F_{1}+F_{2}.$ (1.18)
-The cross-shell interaction term $F_{2}$ is favorable, as it consists of products of flows concentrated on distinct frequency shells.
+The cross-shell interaction term $F_{2}$ is favorable, as it remains in high-frequency region.
 By contrast, $F_{1}$ exhibits a more delicate structure.
 Indeed, within the nonlinear interaction $\operatorname{div}(V_{j}\otimes V_{j})$, resonances among modes inside the frequency shell $\Lambda_{j}$ generate an inverse-cascade transfer toward low frequencies, which exhibit the worst regularity.
-By selecting the directional set $\mathcal{K}$ and the amplitudes $f_{k,j}$ appropriately through Nash’s geometric lemma, one can arrange for this low-frequency contribution to cancel $\Delta V_{j-1}$.
+By selecting the directional set $\mathcal{K}$ and the amplitudes $f_{k,j}$ appropriately through the Nash geometric lemma, one may arrange for this low-frequency contribution to cancel the linear part $\Delta V_{j-1}$.
 Consequently, the forcing term $F$ attains improved regularity and smaller magnitude relative to $V$.
@@
 For $\mathbb{M}\subset\mathbb{Z}$, we define $2^{\mathbb{M}}:=\{2^{m}\ ;\ m\in\mathbb{M}\}$.
+For two quantities $A_{+}$ and $A_{-}$, we use the notation $\sum_{\pm}A_{\pm}:=A_{+}+A_{-}$.
 The identity matrix is denoted by $\rm Id$}.
-For a metric space $(X,d)$, $x_{0}\in X$, and a positive number $r>0$, set $B_{X}(x_{0},r):=\{x\in X\ ;\ d(x,x_{0})\leqslant r\}$.
-Let $\mathscr{S}(\mathbb{R}^{n})$ be the set of all Schwartz functions on $\mathbb{R}^{n}$, and we denote by $\mathscr{S}^{\prime}(\mathbb{R}^{n})$ the set of all tempered distributions on $\mathbb{R}^{n}$.
+For a metric space $(X,d)$, $x_{0}\in X$, and $r>0$, set $B_{X}(x_{0},r):=\{x\in X\ ;\ d(x,x_{0})\leqslant r\}$.
+Let $\mathscr{S}(\mathbb{R}^{n})$ be the set of all Schwartz functions on $\mathbb{R}^{n}$, and $\mathscr{S}^{\prime}(\mathbb{R}^{n})$ denotes the set of all tempered distributions on $\mathbb{R}^{n}$.
 Next, we recall the definitions of Besov and Chemin–Lerner spaces. Let $\varphi_{0}\in\mathscr{S}(\mathbb{R}^{n})$ satisfy
@@
 where we have set $\varphi_{j}(x):=2^{nj}\varphi_{0}(2^{j}x)$. Using this $\{\varphi_{j}\}_{j\in\mathbb{Z}}$, we may define the Littlewood–Paley decomposition as
-$\Delta_{j}f:=\mathscr{F}^{-1}\left[\widehat{\varphi_{j}}(\xi)\widehat{f}(\xi)\right],\qquad f\in\mathscr{S}^{\prime}(\mathbb{R}^{n}),\quad j\in\mathbb{Z}.$ (2.3)
-For $1\leqslant p,q\leqslant\infty$ and $s\in\mathbb{R}$, the Besov space $\dot{B}_{p,q}^{s}(\mathbb{R}^{n})$ is defined as
+$\Delta_{j}f:=\mathscr{F}^{-1}\left[\widehat{\varphi_{j}}(\xi)\widehat{f}(\xi)\right]$ (2.3)
+for $f\in\mathscr{S}^{\prime}(\mathbb{R}^{n})$ and $j\in\mathbb{Z}$. For $1\leqslant p,q\leqslant\infty$ and $s\in\mathbb{R}$, the Besov space $\dot{B}_{p,q}^{s}(\mathbb{R}^{n})$ is defined as
 $\dot{B}_{p,q}^{s}(\mathbb{R}^{n}):={}$ $\left\{f\in\mathscr{S}^{\prime}(\mathbb{R}^{n})/\mathscr{P}(\mathbb{R}^{n})\ ;\ {\left\|f\right\|}_{\dot{B}_{p,q}^{s}}<\infty\right\},$ (2.4)
@@
 $\frac{n}{p}+\frac{n}{r}>2,\qquad 0\leqslant\frac{n}{r}-\frac{n}{p}<1,$ (2.7)
-there exists a positive constant $C=C(n,p,q,r)$ such that
+there exists a constant $C=C(n,p,q,r)>0$ such that
 ${\left\|(-\Delta)^{-1}\mathbb{P}\operatorname{div}(V\otimes W)\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\leqslant C{\left\|V\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}{\left\|W\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}$ (2.8)
@@
 $\widetilde{C}(I;\dot{B}_{p,q}^{s}(\mathbb{R}^{n})):=C(I;\dot{B}_{p,q}^{s}(\mathbb{R}^{n}))\cap\widetilde{L^{\infty}}(I;\dot{B}_{p,q}^{s}(\mathbb{R}^{n})).$ (2.11)
-The Chemin–Lerner spaces are well-known to be particularly convenient when exploiting the maximal regularity of the heat kernel, and we take advantage of this framework through the following lemma, which will be used in our stability analysis for non-stationary flows around the steady state.
+The Chemin–Lerner spaces are well-known to be particularly convenient when exploiting the maximal regularity of the heat kernel.
+We take advantage of this framework through the following lemma, which will be used in our stability analysis for non-stationary flows around the steady state.
 Lemma 2.2 ([Fuj-AIPHC]).
-Let $n\geqslant 2$. Let $1\leqslant p,q,\theta\leqslant\infty$ satisfy
-$2<\theta<\infty,\qquad\min\left\{n,\frac{2n}{p}\right\}-2+\frac{2}{\theta}>0.$ (2.12)
-Then, there exists a positive constant $C=C(n,p,q,\theta)$ such that
+Let $n\geqslant 2$. Let $p$, $q$, $\theta$ satisfy
+$2\leqslant p<2n,\qquad 1\leqslant q\leqslant\infty,\qquad 2<\theta<\infty,\qquad\frac{n}{p}-1+\frac{1}{\theta}>0.$ (2.12)
+Then, there exists a constant $C=C(n,p,q,\theta)>0$ such that
 ${\left\|\int_{t_{0}}^{t}e^{(t-\tau)\Delta}\mathbb{P}\operatorname{div}(v(\tau)\otimes w(\tau))d\tau\right\|}_{\widetilde{L^{\infty}}(I;\dot{B}_{p,q}^{\frac{n}{p}-1})\cap\widetilde{L^{\theta}}(I;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}$ (2.13)
 $\quad\leqslant C{\left\|v\right\|}_{\widetilde{L^{\infty}}(I;\dot{B}_{p,q}^{\frac{n}{p}-1})}{\left\|w\right\|}_{\widetilde{L^{\theta}}(I;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}$ (2.14)
-for all $I=(t_{0},t_{1})\subset\mathbb{R}$, $v\in\widetilde{L^{\infty}}(I;\dot{B}_{p,q}^{\frac{n}{p}-1}(\mathbb{R}^{n}))$, and $w\in\widetilde{L^{\theta}}(I;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}}(\mathbb{R}^{n}))$.
+for all $I=(t_{0},t_{1})\subset\mathbb{R}$, $v\in\widetilde{L^{\infty}}(I;\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))$, and $w\in\widetilde{L^{\theta}}(I;\dot{B}_{p,q}^{n/p-1+2/\theta}(\mathbb{R}^{n}))$.
 3. Highly oscillating functions in Besov spaces
-In the proof of our main result, we frequently use the sharp Besov estimates for finite superpositions of functions obtained by modulating a smooth amplitude with a plane wave of the form (3.4) below.
-If each Fourier transform $\widehat{f_{k}}$ of the amplitudes were compactly supported in a low frequency region, then the Fourier support of each $g_{\lambda,\ell,k}$ is concentrated around $\xi\sim\lambda^{\ell}a_{k}$, and we have
-${\left\|g_{\lambda,\ell,k}\right\|}_{\dot{B}_{p,q}^{s}}\sim|b_{\ell}|(\lambda^{\ell}|a_{k}|)^{s}{\left\|f_{k}\right\|}_{L^{p}}$ (3.1)
+In our analysis, we frequently use the sharp Besov estimates for finite superpositions of functions obtained by modulating a smooth amplitude with a plane wave of the form (3.4) below.
+If each Fourier transform $\widehat{f_{k}}$ of the amplitudes were compactly supported in a low frequency region, then the Fourier support of each $g_{\lambda,k,\ell}$ is concentrated around $\xi\sim\lambda^{\ell}a_{k}$, and we have
+${\left\|g_{\lambda,k,\ell}\right\|}_{\dot{B}_{p,q}^{s}}\sim|b_{\ell}|\lambda^{s\ell}{\left\|f_{k}\right\|}_{L^{p}}$ (3.1)
 For sufficiently large $\lambda$, these supports are disjoint, which implies
-$\|F_{\lambda,L}\|_{\dot{B}^{s}_{p,q}}\sim\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}\sum_{k=1}^{K}|a_{k}|^{s}{\left\|f_{k}\right\|}_{L^{p}}.$ (3.2)
+$\|F_{\lambda,L}\|_{\dot{B}^{s}_{p,q}}\sim\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}\sum_{k=1}^{K}{\left\|f_{k}\right\|}_{L^{p}}.$ (3.2)
 However, in our construction of the non-unique steady Navier–Stokes flow, the amplitudes $f_{1},\dots,f_{K}$ necessarily belong to $C_{c}^{\infty}(\mathbb{R}^{n})$, so their Fourier transforms decay rapidly but are not compactly supported.
@@
 Let $K,L\in\mathbb{N}$.
-Let $a_{1},\dots,a_{K}\subset\mathbb{R}^{n}\setminus\{0\}$ satisfy $|a_{k}|\in 2^{\mathbb{N}}$ $(k=1,\dots,K)$ and assume that $a_{k}$, $a_{k^{\prime}}$ do not point in the same direction unless $k=k^{\prime}$.
-For $\lambda\in 2^{\mathbb{N}}$, $b_{1},\dots,b_{L}\in\mathbb{R}$, and $f_{1},\dots,f_{K}\in\mathscr{S}(\mathbb{R}^{n})$, set
-$F_{\lambda,L}(x):=\sum_{k=1}^{K}\sum_{\ell=1}^{L}g_{\lambda,\ell,k}(x),\qquad g_{\lambda,\ell,k}(x):=b_{\ell}f_{k}(x)e^{i\lambda^{\ell}a_{k}\cdot x}.$ (3.4)
-Then, for any $M\in\mathbb{N}$ with $M>s$, there exists a positive constant $C$, depending only on $n$, $p$, $q$, $s$, $K$, and $M$ such that
-${\left\|F_{\lambda,L}\right\|}_{\dot{B}_{p,q}^{s}}\leqslant{}$ $C\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}\sum_{k=1}^{K}|a_{k}|^{s}{\left\|f_{k}\right\|}_{L^{p}}$ (3.5)
-$+C\sum_{k=1}^{K}\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s+n(1-\frac{1}{p})-M}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}\cap L^{1}}$ (3.6)
+Let $a_{1},\dots,a_{K}\subset\mathbb{S}^{n-1}$ satisfy $a_{k}\neq a_{k^{\prime}}$ unless $k=k^{\prime}$.
+For $\lambda\in 2^{\mathbb{N}}$, $b_{1},\dots,b_{L}\in\mathbb{C}$, and $f_{1},\dots,f_{K}\in\mathscr{S}(\mathbb{R}^{n})$, set
+$F_{\lambda,L}(x):=\sum_{k=1}^{K}\sum_{\ell=1}^{L}g_{\lambda,k,\ell}(x),\qquad g_{\lambda,k,\ell}(x):=b_{\ell}f_{k}(x)e^{i\lambda^{\ell}a_{k}\cdot x}.$ (3.4)
+Then, for any $M\in\mathbb{N}$ with $M>s$, there exists a constant $C>0$, depending only on $n$, $p$, $q$, $s$, $K$, $a_{1},\dots,a_{K}$, and $M$ such that
+${\left\|F_{\lambda,L}\right\|}_{\dot{B}_{p,q}^{s}}\leqslant{}$ $C\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}\sum_{k=1}^{K}{\left\|f_{k}\right\|}_{L^{p}}$ (3.5)
+$+C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\lambda^{(s+n(1-\frac{1}{p})-M)\ell}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}\cap L^{1}}$ (3.6)
 and
-${\left\|F_{\lambda,L}\right\|}_{\dot{B}_{p,q}^{s}}\geqslant{}$ $C^{-1}\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}\max_{1\leqslant k\leqslant K}|a_{k}|^{s}{\left\|f_{k}\right\|}_{L^{p}}$ (3.7)
-$-C\max_{1\leqslant k\leqslant K}\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s-1}|b_{\ell}|{\left\|\nabla f_{k}\right\|}_{L^{p}}$ (3.8)
-$-C\max_{1\leqslant k\leqslant K}\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s+n(1-\frac{1}{p})-M}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}\cap L^{1}}$ (3.9)
+${\left\|F_{\lambda,L}\right\|}_{\dot{B}_{p,q}^{s}}\geqslant{}$ $C^{-1}\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}\max_{1\leqslant k\leqslant K}{\left\|f_{k}\right\|}_{L^{p}}$ (3.7)
+$-C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\lambda^{(s-1)\ell}|b_{\ell}|{\left\|\nabla f_{k}\right\|}_{L^{p}}$ (3.8)
+$-C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\lambda^{(s+n(1-\frac{1}{p})-M)\ell}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}\cap L^{1}}$ (3.9)
 for all $L\in\mathbb{N}$ and $\lambda\in 2^{\mathbb{N}}$ with $\lambda\geqslant 2^{100}$.
 Remark 3.2.
-To obtain (3.6) with $q=1$, we may remove the directional assumption on $a_{1},\dots,a_{K}$. Indeed, applying (3.6) with $K=L=1$ to each $f_{k}(x)e^{i\lambda^{\ell}a_{k}\cdot x}$ gives
+For (3.6) with $q=1$, we may remove the directional assumption on $a_{1},\dots,a_{K}$. Indeed, applying (3.6) with $K=L=1$ to each $g_{k,\ell}$ gives
 ${\left\|F_{\lambda,L}\right\|}_{\dot{B}_{p,1}^{s}}\leqslant{}$ $\sum_{k=1}^{K}\sum_{\ell=1}^{L}|b_{\ell}|{\left\|f_{k}(x)e^{i\lambda^{\ell}a_{k}\cdot x}\right\|}_{\dot{B}_{p,1}^{s}}$ (3.10)
-$\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s}|b_{\ell}|{\left\|f_{k}\right\|}_{L^{p}}$ (3.11)
-$+C\sum_{k=1}^{K}\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s+n(1-\frac{1}{p})-M}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}\cap L^{1}}.$ (3.12)
+$\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\lambda^{s\ell}|b_{\ell}|{\left\|f_{k}\right\|}_{L^{p}}$ (3.11)
+$+C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\lambda^{(s+n(1-\frac{1}{p})-M)\ell}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}\cap L^{1}}.$ (3.12)
 To prove Proposition 3.1, we introduce operators that localize frequencies in the angular direction.
@@
 for $k=0,\dots,K$ and $f\in\mathscr{S}^{\prime}(\mathbb{R}^{n})$ with $0\notin\operatorname{supp}\widehat{f}$. We remark that there holds
-$P_{m}\Delta_{j}g_{\lambda,\ell,k}$ $=e^{i\lambda^{\ell}a_{k}\cdot x}\mathscr{F}^{-1}\left[\widehat{\chi_{m,j}}(\xi+\lambda^{\ell}a_{k})\widehat{f_{k}}(\xi)\right]$ (3.16)
-$=e^{i\lambda^{\ell}a_{k}\cdot x}\left(\chi_{m,j,\ell,k}*f_{k}\right)(x),$ (3.17)
+$P_{m}\Delta_{j}g_{\lambda,k,\ell}$ $=e^{i\lambda^{\ell}a_{k}\cdot x}\mathscr{F}^{-1}\left[\widehat{\chi_{m,j}}(\xi+\lambda^{\ell}a_{k})\widehat{f_{k}}(\xi)\right]$ (3.16)
+$=e^{i\lambda^{\ell}a_{k}\cdot x}\left(\chi_{m,j,k,\ell}*f_{k}\right)(x),$ (3.17)
 where we have defined
-$\chi_{m,j}:=\mathscr{F}^{-1}[\phi_{m}\widehat{\varphi_{j}}],\qquad\chi_{m,j,\ell,k}(x):=e^{-i\lambda^{\ell}a_{k}\cdot x}\chi_{m,j}(x).$ (3.18)
-For these kernels, there holds ${\left\|\chi_{m,j,\ell,k}\right\|}_{L^{\rho}}={\left\|\chi_{m,j}\right\|}_{L^{\rho}}=2^{n(1-\frac{1}{\rho})j}{\left\|\chi_{m,0}\right\|}_{L^{\rho}}$ holds for $1\leqslant\rho\leqslant\infty$.
+$\chi_{m,j}:=\mathscr{F}^{-1}[\phi_{m}\widehat{\varphi_{j}}],\qquad\chi_{m,j,k,\ell}(x):=e^{-i\lambda^{\ell}a_{k}\cdot x}\chi_{m,j}(x).$ (3.18)
+For these kernels, there holds ${\left\|\chi_{m,j,k,\ell}\right\|}_{L^{\rho}}={\left\|\chi_{m,j}\right\|}_{L^{\rho}}=2^{n(1-\frac{1}{\rho})j}{\left\|\chi_{m,0}\right\|}_{L^{\rho}}$ holds for $1\leqslant\rho\leqslant\infty$.
 Moreover, we prepare some notations; for every $k\in\{1,\dots,K\}$ and $\ell\in\{1,\dots,L\}$, set
-$j_{\ell,k}:=\log_{2}\left(\lambda^{\ell}|a_{k}|\right),$ (3.19)
-$R_{\ell,k}:=\{(m,j)\in\{0,\dots,K\}\times\mathbb{Z}\ ;\ (m,j)\neq(k,j_{\ell,k}),(k,j_{\ell,k}\pm 1)\}.$ (3.20)
+$j_{\ell}:=\ell\log_{2}\lambda,$ (3.19)
+$R_{k,\ell}:=\{(m,j)\in\{0,\dots,K\}\times\mathbb{Z}\ ;\ (m,j)\neq(k,j_{\ell}),(k,j_{\ell}\pm 1)\}.$ (3.20)
 For the proof of Proposition 3.1, we prepare a lemma.
 Lemma 3.3.
-There exists a constant $c_{*}=c_{*}(a_{1},\dots,a_{K})\in 2^{\mathbb{Z}}$ such that for any $\ell\in\{1,\dots,L\}$, $k\in\{1,\dots,K\}$, $(m,j)\in R_{\ell,k}$, and $\lambda\in 2^{\mathbb{N}}$, there holds
-$\operatorname{supp}\left[\widehat{\chi_{m,j}}(\cdot+\lambda^{\ell}a_{k})\right]\subset\left\{\xi\in\mathbb{R}^{n}\ ;\ |\xi|\geqslant c_{*}\max\{2^{j},\lambda^{\ell}|a_{k}|\}\right\}.$ (3.21)
+There exists a constant $c_{*}=c_{*}(K,a_{1},\dots,a_{K})\in 2^{\mathbb{Z}}$ such that for any $\ell\in\{1,\dots,L\}$, $k\in\{1,\dots,K\}$, $(m,j)\in R_{k,\ell}$, and $\lambda\in 2^{\mathbb{N}}$, there holds
+$\operatorname{supp}\mathscr{F}\left[\chi_{m,j,k,\ell}\right]\subset\left\{\xi\in\mathbb{R}^{n}\ ;\ |\xi|\geqslant c_{*}\max\{2^{j},\lambda^{\ell}\}\right\}.$ (3.21)
 Proof.
-Let $\xi\in\mathbb{R}^{n}$ satisfy $\widehat{\chi_{m,j}}(\xi+\lambda^{\ell}a_{k})\neq 0$.
+Let $\xi\in\mathbb{R}^{n}$ satisfy $\mathscr{F}\left[\chi_{m,j,k,\ell}\right](\xi)=\widehat{\chi_{m,j}}(\xi+\lambda^{\ell}a_{k})\neq 0$.
 We first consider the case of $m=k$.
-Then, by $2^{j-1}\leqslant|\xi+\lambda^{\ell}a_{k}|\leqslant 2^{j+1}$ and $|\lambda^{\ell}a_{k}|=2^{j_{\ell,k}}$, it holds
+Then, by $2^{j-1}\leqslant|\xi+\lambda^{\ell}a_{k}|\leqslant 2^{j+1}$ and $\lambda^{\ell}=2^{j_{\ell}}$, it holds
 $|\xi|$ $=|\xi+\lambda^{\ell}a_{k}-\lambda^{\ell}a_{k}|$ (3.22)
-$\geqslant\begin{cases}|\xi+\lambda^{\ell}a_{k}|-|\lambda^{\ell}a_{k}|\geqslant 2^{j-1}-2^{j-2}&(j\geqslant j_{\ell,k}+2),\\ |\lambda^{\ell}a_{k}|-|\xi+\lambda^{\ell}a_{k}|\geqslant 2^{j_{\ell,k}}-2^{j_{\ell,k}-1}&(j\leqslant j_{\ell,k}-2),\end{cases}$ (3.23)
-$\geqslant\frac{1}{4}\max\left\{2^{j},|\lambda^{\ell}a_{k}|\right\}.$ (3.24)
+$\geqslant\begin{cases}|\xi+\lambda^{\ell}a_{k}|-\lambda^{\ell}\geqslant 2^{j-1}-2^{j-2}&(j\geqslant j_{\ell}+2),\\ \lambda^{\ell}-|\xi+\lambda^{\ell}a_{k}|\geqslant 2^{j_{\ell}}-2^{j_{\ell}-1}&(j\leqslant j_{\ell}-2),\end{cases}$ (3.23)
+$\geqslant c\max\left\{2^{j},\lambda^{\ell}\right\}.$ (3.24)
 For the case of $m\neq k$, we see that
-$|\xi|^{2}={}$ $|\xi+\lambda^{\ell}a_{k}|^{2}+|\lambda^{\ell}a_{k}|^{2}-2|\xi+\lambda^{\ell}a_{k}||\lambda^{\ell}a_{k}|\cos\vartheta(\xi+\lambda^{\ell}a_{k},\lambda^{\ell}a_{k})$ (3.25)
-$={}$ $\frac{1}{2}\left(|\xi+\lambda^{\ell}a_{k}|-|\lambda^{\ell}a_{k}|\cos\vartheta(\xi+\lambda^{\ell}a_{k},\lambda^{\ell}a_{k})\right)^{2}$ (3.26)
-$+\frac{1}{2}\left(|\lambda^{\ell}a_{k}|-|\xi+\lambda^{\ell}a_{k}|\cos\vartheta(\xi+\lambda^{\ell}a_{k},\lambda^{\ell}a_{k})\right)^{2}$ (3.27)
-$+\frac{1}{2}\sin^{2}\vartheta(\xi+\lambda^{\ell}a_{k},\lambda^{\ell}a_{k})\left(|\lambda^{\ell}a_{k}|^{2}+|\xi+\lambda^{\ell}a_{k}|^{2}\right)$ (3.28)
-$\geqslant{}$ $\frac{1}{2}\sin^{2}(\vartheta_{*}/6)\left(|\lambda^{\ell}a_{k}|^{2}+|\xi+\lambda^{\ell}a_{k}|^{2}\right),$ (3.29)
-which yields $|\xi|\geqslant c\sin(\vartheta_{*}/6)\max\left\{2^{j},|\lambda^{\ell}a_{k}|\right\}$. Thus, we complete the proof. ∎
+$|\xi|^{2}={}$ $|\xi+\lambda^{\ell}a_{k}|^{2}+\lambda^{2\ell}-2\lambda^{\ell}|\xi+\lambda^{\ell}a_{k}|\cos\vartheta(\xi+\lambda^{\ell}a_{k},\lambda^{\ell}a_{k})$ (3.25)
+$={}$ $\frac{1}{2}\left(|\xi+\lambda^{\ell}a_{k}|-\lambda^{\ell}\cos\vartheta(\xi+\lambda^{\ell}a_{k},\lambda^{\ell}a_{k})\right)^{2}$ (3.26)
+$+\frac{1}{2}\left(\lambda^{\ell}-|\xi+\lambda^{\ell}a_{k}|\cos\vartheta(\xi+\lambda^{\ell}a_{k},\lambda^{\ell}a_{k})\right)^{2}$ (3.27)
+$+\frac{1}{2}\sin^{2}\vartheta(\xi+\lambda^{\ell}a_{k},\lambda^{\ell}a_{k})\left(\lambda^{2\ell}+|\xi+\lambda^{\ell}a_{k}|^{2}\right)$ (3.28)
+$\geqslant{}$ $\frac{1}{2}\sin^{2}(\vartheta_{*}/6)\left(\lambda^{2\ell}+|\xi+\lambda^{\ell}a_{k}|^{2}\right),$ (3.29)
+which yields $|\xi|\geqslant c\max\left\{2^{j},\lambda^{\ell}\right\}$. Thus, we complete the proof. ∎
 Let us now provide the proof of Proposition 3.1
@@
 Step 1. The upper bound estimate. We may decompose $F_{\lambda,L}$ as
-$F_{\lambda,L}={}\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{k^{\prime}=0}^{K}\sum_{j\in\mathbb{Z}}P_{k^{\prime}}\Delta_{j}g_{\lambda,\ell,k}={}F_{\lambda,L}^{({\rm p})}+F_{\lambda,L}^{({\rm r})},$ (3.30)
+$F_{\lambda,L}={}\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{m=0}^{K}\sum_{j\in\mathbb{Z}}P_{m}\Delta_{j}g_{\lambda,k,\ell}={}F_{\lambda,L}^{({\rm p})}+F_{\lambda,L}^{({\rm r})},$ (3.30)
 where the principal part $F_{\lambda,L}^{({\rm p})}$ and the remainder part $F_{\lambda,L}^{({\rm r})}$ are defined by
-$F_{\lambda,L}^{({\rm p})}:=\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{|j-j_{\ell,k}|\leqslant 2}P_{k}\Delta_{j}g_{\lambda,\ell,k},$ (3.31)
-$\begin{aligned} F_{\lambda,L}^{({\rm r})}:={}&\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{\begin{subarray}{c}0\leqslant k^{\prime}\leqslant K\\ k^{\prime}\neq k\end{subarray}}\sum_{|j-j_{\ell,k}|\leqslant 2}P_{k^{\prime}}\Delta_{j}g_{\lambda,\ell,k}\\ &+\sum_{k^{\prime}=0}^{K}\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{|j-j_{\ell,k}|\geqslant 3}P_{k^{\prime}}\Delta_{j}g_{\lambda,\ell,k}.\end{aligned}$ (3.32)
+$F_{\lambda,L}^{({\rm p})}:=\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{|j-j_{\ell}|\leqslant 2}P_{k}\Delta_{j}g_{\lambda,k,\ell},$ (3.31)
+$\begin{aligned} F_{\lambda,L}^{({\rm r})}:={}&\sum_{m=0}^{K}\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{(m,j)\in R_{k,\ell}}P_{m}\Delta_{j}g_{\lambda,k,\ell}.\end{aligned}$ (3.32)
 For the estimate of the principal part, we see that
-${\left\|F_{\lambda,L}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{s}}$ $\leqslant\sum_{k=1}^{K}\left\{\sum_{j^{\prime}\in\mathbb{Z}}\left(2^{sj}{\left\|\sum_{\ell=1}^{L}\sum_{|j-j_{\ell,k}|\leqslant 2}P_{k}\Delta_{j}\Delta_{j^{\prime}}g_{\lambda,\ell,k}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}$ (3.33)
-$=\sum_{k=1}^{K}\left\{\sum_{\ell=1}^{L}\sum_{|j^{\prime}-j_{\ell,k}|\leqslant 3}\left(2^{sj}{\left\|\sum_{|j-j_{\ell,k}|\leqslant 2}P_{k}\Delta_{j}\Delta_{j^{\prime}}g_{\lambda,\ell,k}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}$ (3.34)
-$\leqslant C\sum_{k=1}^{K}\left\{\sum_{\ell=1}^{L}\sum_{|j^{\prime}-j_{\ell,k}|\leqslant 3}\left(2^{sj}{\left\|g_{\lambda,\ell,k}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}$ (3.35)
-$=C\sum_{k=1}^{K}\left\{\sum_{\ell=1}^{L}\sum_{|j^{\prime}-j_{\ell,k}|\leqslant 3}\left(2^{sj}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}{\left\|f_{k}\right\|}_{L^{p}}$ (3.36)
-$\leqslant C\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}\sum_{k=1}^{K}|a_{k}|^{s}{\left\|f_{k}\right\|}_{L^{p}}.$ (3.37)
+${\left\|F_{\lambda,L}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{s}}$ $\leqslant\sum_{k=1}^{K}\left\{\sum_{j^{\prime}\in\mathbb{Z}}\left(2^{sj^{\prime}}{\left\|\sum_{\ell=1}^{L}\sum_{|j-j_{\ell}|\leqslant 1}P_{k}\Delta_{j}\Delta_{j^{\prime}}g_{\lambda,k,\ell}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}$ (3.33)
+$=\sum_{k=1}^{K}\left\{\sum_{\ell=1}^{L}\sum_{|j^{\prime}-j_{\ell}|\leqslant 2}\left(2^{sj}{\left\|\sum_{|j-j_{\ell}|\leqslant 1}P_{k}\Delta_{j}\Delta_{j^{\prime}}g_{\lambda,k,\ell}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}$ (3.34)
+$\leqslant C\sum_{k=1}^{K}\left\{\sum_{\ell=1}^{L}\sum_{|j^{\prime}-j_{\ell}|\leqslant 2}\left(2^{sj}{\left\|g_{\lambda,k,\ell}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}$ (3.35)
+$=C\sum_{k=1}^{K}\left\{\sum_{\ell=1}^{L}\sum_{|j^{\prime}-j_{\ell}|\leqslant 2}\left(2^{sj}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}{\left\|f_{k}\right\|}_{L^{p}}$ (3.36)
+$\leqslant C\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}\sum_{k=1}^{K}{\left\|f_{k}\right\|}_{L^{p}}.$ (3.37)
 For the remainder terms, there holds
-${\left\|F_{\lambda,L}^{({\rm r})}\right\|}_{\dot{B}_{p,1}^{s}}$ $\leqslant\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{\begin{subarray}{c}0\leqslant k^{\prime}\leqslant K\\ k^{\prime}\neq k\end{subarray}}\sum_{|j-j_{\ell,k}|\leqslant 2}\sum_{|j^{\prime}-j|\leqslant 1}2^{sj^{\prime}}{\left\|P_{k^{\prime}}\Delta_{j}\Delta_{j^{\prime}}g_{\lambda,\ell,k}\right\|}_{L^{p}}$ (3.38)
-$\qquad+\sum_{k=1}^{K}\sum_{k^{\prime}=0}^{K}\sum_{\ell=1}^{L}\sum_{|j-j_{\ell,k}|\geqslant 3}\sum_{|j^{\prime}-j|\leqslant 1}2^{sj^{\prime}}{\left\|P_{k^{\prime}}\Delta_{j}\Delta_{j^{\prime}}g_{\lambda,\ell,k}\right\|}_{L^{p}}$ (3.39)
-$\leqslant C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{(k^{\prime},j)\in R_{\ell,k}}2^{sj}{\left\|P_{k^{\prime}}\Delta_{j}g_{\lambda,\ell,k}\right\|}_{L^{p}}.$ (3.40)
+${\left\|F_{\lambda,L}^{({\rm r})}\right\|}_{\dot{B}_{p,1}^{s}}$ $\leqslant\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{(m,j)\in R_{k,\ell}}2^{sj^{\prime}}{\left\|P_{k^{\prime}}\Delta_{j}\Delta_{j^{\prime}}g_{\lambda,k,\ell}\right\|}_{L^{p}}$ (3.38)
+$\leqslant C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{(k^{\prime},j)\in R_{k,\ell}}2^{sj}{\left\|P_{k^{\prime}}\Delta_{j}g_{\lambda,k,\ell}\right\|}_{L^{p}}.$ (3.39)
 Here, Lemma 3.3 yields
-${\left\|P_{m}\Delta_{j}g_{\lambda,\ell,k}\right\|}_{L^{p}}$ $\leqslant|b_{\ell}|\sum_{j^{\prime}\geqslant\max\{j,j_{\ell,k}\}+j_{*}}{\left\|\Delta_{j^{\prime}}\chi_{m,j,\ell,k}*f_{k}\right\|}_{L^{p}}$ (3.41)
-$\leqslant C|b_{\ell}|\sum_{j^{\prime}\geqslant\max\{j,j_{\ell,k}\}+j_{*}}2^{-Mj^{\prime}}{\left\|\chi_{m,j,\ell,k}*\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.42)
-$=C|b_{\ell}|\left(\max\{2^{j},\lambda^{\ell}|a_{k}|\}\right)^{-M}{\left\|\chi_{m,j,\ell,k}*\nabla^{M}f_{k}\right\|}_{L^{p}},$ (3.43)
-where we have put $j_{*}=\log_{2}(2c_{*})$. Then, we obtain
-${\left\|F_{\lambda,L}^{({\rm r})}\right\|}_{\dot{B}_{p,1}^{s}}\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{(k^{\prime},j)\in R_{\ell,k}}|b_{\ell}|2^{sj}\left(\max\{2^{j},\lambda^{\ell}|a_{k}|\}\right)^{-M}{\left\|\chi_{k^{\prime},j,\ell,k}*\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.44)
-$\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{k^{\prime}=0}^{K}\sum_{j\geqslant j_{\ell,k}+1}|b_{\ell}|2^{(s-M)j}{\left\|\chi_{k^{\prime},j,\ell,k}\right\|}_{L^{1}}{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.45)
-$+C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{k^{\prime}=0}^{K}\sum_{j\leqslant j_{\ell,k}}|b_{\ell}|2^{sj}(\lambda^{\ell}|a_{k}|)^{-M}{\left\|\chi_{k^{\prime},j,\ell,k}\right\|}_{L^{p}}{\left\|\nabla^{M}f_{k}\right\|}_{L^{1}}$ (3.46)
-$\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{j\geqslant j_{\ell,k}+1}|b_{\ell}|2^{(s-M)j}{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.47)
-$+C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{j\leqslant j_{\ell,k}}|b_{\ell}|2^{\{s+n(1-\frac{1}{p})\}j}(\lambda^{\ell}|a_{k}|)^{-M}{\left\|\nabla^{M}f_{k}\right\|}_{L^{1}}$ (3.48)
-$\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s-M}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.49)
-$+C\sum_{k=1}^{K}\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s+n(1-\frac{1}{p})-M}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{1}}.$ (3.50)
+${\left\|P_{m}\Delta_{j}g_{\lambda,k,\ell}\right\|}_{L^{p}}$ $\leqslant|b_{\ell}|\sum_{j^{\prime}\geqslant\max\{j,j_{\ell}\}+j_{*}}{\left\|\Delta_{j^{\prime}}\chi_{m,j,k,\ell}*f_{k}\right\|}_{L^{p}}$ (3.40)
+$\leqslant C|b_{\ell}|\sum_{j^{\prime}\geqslant\max\{j,j_{\ell}\}+j_{*}}2^{-Mj^{\prime}}{\left\|\chi_{m,j,k,\ell}*\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.41)
+$=C|b_{\ell}|\left(\max\{2^{j},\lambda^{\ell}\}\right)^{-M}{\left\|\chi_{m,j,k,\ell}*\nabla^{M}f_{k}\right\|}_{L^{p}},$ (3.42)
+where we have put $j_{*}=\log_{2}(2c_{*})$. Thus, we obtain
+${\left\|F_{\lambda,L}^{({\rm r})}\right\|}_{\dot{B}_{p,1}^{s}}\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{(m,j)\in R_{k,\ell}}|b_{\ell}|2^{sj}\left(\max\{2^{j},\lambda^{\ell}\}\right)^{-M}{\left\|\chi_{m,j,k,\ell}*\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.43)
+$\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{m=0}^{K}\sum_{j\geqslant j_{\ell}+1}|b_{\ell}|2^{(s-M)j}{\left\|\chi_{m,j,k,\ell}\right\|}_{L^{1}}{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.44)
+$+C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{m=0}^{K}\sum_{j\leqslant j_{\ell}}|b_{\ell}|2^{sj}(\lambda^{-M\ell}{\left\|\chi_{m,j,k,\ell}\right\|}_{L^{p}}{\left\|\nabla^{M}f_{k}\right\|}_{L^{1}}$ (3.45)
+$\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{j\geqslant j_{\ell}+1}|b_{\ell}|2^{(s-M)j}{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.46)
+$+C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{j\leqslant j_{\ell}}|b_{\ell}|2^{\{s+n(1-\frac{1}{p})\}j}\lambda^{-M\ell}{\left\|\nabla^{M}f_{k}\right\|}_{L^{1}}$ (3.47)
+$\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\lambda^{(s-M)\ell}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.48)
+$+C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\lambda^{(s+n(1-\frac{1}{p})-M)\ell}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{1}}.$ (3.49)
 Combining (3.37) and (3.49), we obtain the desired estimate.
 Step 2. The lower bound estimate. Since $P_{0},P_{1},\dots,P_{K}$ are bounded in homogeneous Besov spaces, there holds
-${\left\|F_{\lambda,L}\right\|}_{\dot{B}_{p,q}^{s}}\geqslant c\max_{1\leqslant k\leqslant K}{\left\|P_{k}F_{\lambda,L}\right\|}_{\dot{B}_{p,q}^{s}}.$ (3.51)
-Fix a $k\in\{1,\dots,K\}$ and decompose $P_{k}F_{\lambda,L}$ as
-$P_{k}F_{\lambda,L}=F^{({\rm p})}_{\lambda,k,L}+F^{({\rm r})}_{\lambda,k,L}$; (3.52)
-$\quad F^{({\rm p})}_{\lambda,k,L}:=\sum_{\ell=1}^{L}\sum_{|j-j_{\ell,k}|\leqslant 2}P_{k}\Delta_{j}g_{\lambda,\ell,k},$ (3.53)
-$\quad F^{({\rm r})}_{\lambda,k,L}:=\sum_{\ell=1}^{L}\sum_{|j-j_{\ell,k}|\geqslant 3}P_{k}\Delta_{j}g_{\lambda,\ell,k}+\sum_{\begin{subarray}{c}1\leqslant k^{\prime}\leqslant K\\ k^{\prime}\neq k\end{subarray}}\sum_{\ell=1}^{L}P_{k}g_{\lambda,\ell,k^{\prime}}.$ (3.54)
-For the lower bound estimate of the principal part $F^{({\rm p})}_{\lambda,k,L}$, we see that
-${\left\|F_{\lambda,k,L}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{s}}$ $=\sum_{k=1}^{K}\left\{\sum_{\ell=1}^{L}\sum_{|j^{\prime}-j_{\ell,k}|\leqslant 3}\left(2^{sj^{\prime}}{\left\|\sum_{|j-j_{\ell,k}|\leqslant 2}P_{k}\Delta_{j}\Delta_{j^{\prime}}g_{\lambda,\ell,k}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}$ (3.55)
-$\geqslant\sum_{k=1}^{K}\left\{\sum_{\ell=1}^{L}\left(2^{sj_{\ell,k}}{\left\|\sum_{|j-j_{\ell,k}|\leqslant 2}P_{k}\Delta_{j}\Delta_{j_{\ell,k}}g_{\lambda,\ell,k}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}$ (3.56)
-$=\sum_{k=1}^{K}\left\{\sum_{\ell=1}^{L}\left(2^{sj_{\ell,k}}{\left\|P_{k}\Delta_{j_{\ell,k}}g_{\lambda,\ell,k}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}.$ (3.57)
+${\left\|F_{\lambda,L}\right\|}_{\dot{B}_{p,q}^{s}}\geqslant c\max_{1\leqslant m\leqslant K}{\left\|P_{m}F_{\lambda,L}\right\|}_{\dot{B}_{p,q}^{s}}.$ (3.50)
+Fix a $m\in\{1,\dots,K\}$ and decompose $P_{m}F_{\lambda,L}$ as
+$P_{m}F_{\lambda,L}=F^{({\rm p})}_{\lambda,m,L}+F^{({\rm r})}_{\lambda,m,L}$; (3.51)
+$\quad F^{({\rm p})}_{\lambda,m,L}:=\sum_{\ell=1}^{L}\sum_{|j-j_{\ell}|\leqslant 1}P_{m}\Delta_{j}g_{\lambda,m,\ell},$ (3.52)
+$\quad F^{({\rm r})}_{\lambda,m,L}:=\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{j:(m,j)\in R_{k,\ell}}P_{m}\Delta_{j}g_{\lambda,k,\ell}.$ (3.53)
+For the lower bound estimate of the principal part $F^{({\rm p})}_{\lambda,m,L}$, we see that
+${\left\|F_{\lambda,m,L}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{s}}$ $=\left\{\sum_{j^{\prime}\in\mathbb{Z}}\left(2^{sj^{\prime}}{\left\|\sum_{\ell=1}^{L}\sum_{|j-j_{\ell}|\leqslant 1}P_{m}\Delta_{j}\Delta_{j^{\prime}}g_{\lambda,m,\ell}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}$ (3.54)
+$=\left\{\sum_{\ell=1}^{L}\sum_{|j^{\prime}-j_{\ell}|\leqslant 2}\left(2^{sj^{\prime}}{\left\|\sum_{|j-j_{\ell}|\leqslant 1}P_{m}\Delta_{j}\Delta_{j^{\prime}}g_{\lambda,m,\ell}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}$ (3.55)
+$\geqslant\left\{\sum_{\ell=1}^{L}\left(2^{sj_{\ell}}{\left\|\sum_{|j-j_{\ell}|\leqslant 1}P_{m}\Delta_{j}\Delta_{j_{\ell}}g_{\lambda,m,\ell}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}$ (3.56)
+$=\left\{\sum_{\ell=1}^{L}\left(2^{sj_{\ell}}{\left\|P_{m}\Delta_{j_{\ell}}g_{\lambda,m,\ell}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}.$ (3.57)
 Here, there holds
-$P_{k}\Delta_{j_{\ell,k}}g_{\lambda,\ell,k}(x)$ $=b_{\ell}\int_{\mathbb{R}^{n}}\chi_{k,j_{\ell,k}}(y)f_{k}(x-y)e^{i\lambda^{\ell}a_{k}\cdot(x-y)}dy$ (3.58)
-$=b_{\ell}e^{i\lambda^{\ell}a_{k}\cdot x}f_{k}(x)\int_{\mathbb{R}^{n}}\chi_{k,j_{\ell,k}}(y)e^{-i\lambda^{\ell}a_{k}\cdot y}dy$ (3.59)
-$\qquad+b_{\ell}e^{i\lambda^{\ell}a_{k}\cdot x}\int_{\mathbb{R}^{n}}(f_{k}(x-y)-f_{k}(x))\chi_{k,j_{\ell,k}}(y)e^{-i\lambda^{\ell}a_{k}\cdot y}dy$ (3.60)
-$=:b_{\ell}e^{i\lambda^{\ell}a_{k}\cdot x}f_{k}(x)+R_{\lambda,\ell,k}(x),$ (3.61)
-where we have used $\widehat{\chi_{k,j_{\ell,k}}}(\lambda^{\ell}a_{k})=1$. The remainder term $R_{\lambda,\ell,k}$ is bounded as
-${\left\|R_{\lambda,\ell,k}\right\|}_{L^{p}}$ $\leqslant|b_{\ell}|{\left\|\nabla f_{k}\right\|}_{L^{p}}\int_{\mathbb{R}^{n}}|y||\chi_{k,j_{\ell,k}}(y)|dy$ (3.62)
-$\leqslant C|b_{\ell}|2^{-j_{\ell,k}}{\left\|\nabla f_{k}\right\|}_{L^{p}}$ (3.63)
-$\leqslant C|b_{\ell}|(\lambda^{\ell}|a_{k}|)^{-1}{\left\|\nabla f_{k}\right\|}_{L^{p}}.$ (3.64)
+$P_{m}\Delta_{j_{\ell}}g_{\lambda,m,\ell}(x)$ $=b_{\ell}\int_{\mathbb{R}^{n}}\chi_{m,j_{\ell}}(y)f_{m}(x-y)e^{i\lambda^{\ell}a_{m}\cdot(x-y)}dy$ (3.58)
+$=b_{\ell}e^{i\lambda^{\ell}a_{m}\cdot x}f_{m}(x)\int_{\mathbb{R}^{n}}\chi_{m,j_{\ell}}(y)e^{-i\lambda^{\ell}a_{m}\cdot y}dy$ (3.59)
+$\qquad+b_{\ell}e^{i\lambda^{\ell}a_{k}\cdot x}\int_{\mathbb{R}^{n}}(f_{m}(x-y)-f_{m}(x))\chi_{m,j_{\ell}}(y)e^{-i\lambda^{\ell}a_{m}\cdot y}dy$ (3.60)
+$=:b_{\ell}e^{i\lambda^{\ell}a_{m}\cdot x}f_{m}(x)+R_{\lambda,m,\ell}(x),$ (3.61)
+where we have used $\widehat{\chi_{m,j_{\ell}}}(\lambda^{\ell}a_{m})=1$. The remainder term $R_{\lambda,m,\ell}$ is bounded as
+${\left\|R_{\lambda,m,\ell}\right\|}_{L^{p}}$ $\leqslant|b_{\ell}|{\left\|\nabla f_{m}\right\|}_{L^{p}}\int_{\mathbb{R}^{n}}|y||\chi_{m,j_{\ell}}(y)|dy$ (3.62)
+$\leqslant C|b_{\ell}|2^{-j_{\ell}}{\left\|\nabla f_{m}\right\|}_{L^{p}}$ (3.63)
+$=C|b_{\ell}|\lambda^{-\ell}{\left\|\nabla f_{m}\right\|}_{L^{p}}.$ (3.64)
 Thus, we see that
-${\left\|P_{k}\Delta_{j_{\ell,k}}g_{\lambda,\ell,k}\right\|}_{L^{p}}$ $\geqslant|b_{\ell}|{\left\|f_{k}\right\|}_{L^{p}}-C(\lambda^{\ell}|a_{k}|)^{-1}|b_{\ell}|{\left\|\nabla f_{k}\right\|}_{L^{p}},$ (3.65)
-which implies
-${\left\|F_{\lambda,k,L}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{s}}\geqslant{}$ $c\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}\max_{1\leqslant k\leqslant K}|a_{k}|^{s}{\left\|f_{k}\right\|}_{L^{p}}$ (3.66)
-$-C\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s-1}|b_{\ell}|{\left\|\nabla f_{k}\right\|}_{L^{p}}.$ (3.67)
-On the estimate of $F^{({\rm r})}_{\lambda,k,L}$, the similar strategy as for the estimate of $F^{({\rm r})}_{\lambda,k,L}$ yields
-${\left\|F^{({\rm r})}_{\lambda,k,L}\right\|}_{\dot{B}_{p,q}^{s}}\leqslant{}$ $\sum_{\ell=1}^{L}\sum_{|j-j_{\ell,k}|\geqslant 3}\sum_{|j^{\prime}-j|\leqslant 1}2^{sj^{\prime}}{\left\|P_{k}\Delta_{j}\Delta_{j^{\prime}}g_{\lambda,\ell,k}\right\|}_{L^{p}}$ (3.68)
-$+\sum_{\begin{subarray}{c}1\leqslant k^{\prime}\leqslant K\\ k^{\prime}\neq k\end{subarray}}\sum_{\ell=1}^{L}\sum_{j\in\mathbb{Z}}2^{sj}{\left\|P_{k}\Delta_{j}g_{\lambda,\ell,k}\right\|}_{L^{p}}$ (3.69)
-$\leqslant{}$ $C\max_{1\leqslant k\leqslant K}\sum_{\ell=1}^{L}\sum_{(k^{\prime},j)\in R_{\ell,k}}2^{sj}{\left\|P_{k^{\prime}}\Delta_{j}g_{\lambda,\ell,k}\right\|}_{L^{p}}$ (3.70)
-$\leqslant{}$ $C\max_{1\leqslant k\leqslant K}\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s-M}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.71)
-$+C\max_{1\leqslant k\leqslant K}\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s+n(1-\frac{1}{p})-M}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}\cap L^{1}}.$ (3.72)
-Gathering (3.67) and (3.72), we complete the proof. ∎
+${\left\|P_{m}\Delta_{j_{\ell}}g_{\lambda,m,\ell}\right\|}_{L^{p}}$ $\geqslant|b_{\ell}|{\left\|f_{m}\right\|}_{L^{p}}-C\lambda^{-\ell}|b_{\ell}|{\left\|\nabla f_{m}\right\|}_{L^{p}}.$ (3.65)
+By (3.57) and (3.65), it holds
+${\left\|F_{\lambda,m,L}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{s}}\geqslant{}$ $c\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}{\left\|f_{m}\right\|}_{L^{p}}-C\sum_{\ell=1}^{L}\lambda^{(s-1)\ell}|b_{\ell}|{\left\|\nabla f_{k}\right\|}_{L^{p}}.$ (3.66)
+On the estimate of $F^{({\rm r})}_{\lambda,m,L}$, the similar strategy as for the estimate of $F^{({\rm r})}_{\lambda,m,L}$ yields
+${\left\|F^{({\rm r})}_{\lambda,m,L}\right\|}_{\dot{B}_{p,1}^{s}}\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{j:(m,j)\in R_{k,\ell}}\sum_{|j^{\prime}-j|\leqslant 1}2^{sj^{\prime}}{\left\|P_{m}\Delta_{j^{\prime}}\Delta_{j}g_{\lambda,k,\ell}\right\|}_{L^{p}}$ (3.67)
+$\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{j:(m,j)\in R_{k,\ell}}2^{sj}{\left\|P_{m}\Delta_{j}g_{\lambda,k,\ell}\right\|}_{L^{p}}$ (3.68)
+$\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\lambda^{(s-M)\ell}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.69)
+$+C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\lambda^{(s+n(1-\frac{1}{p})-M)\ell}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}\cap L^{1}}.$ (3.70)
+Gathering (3.66) and (3.70), we obtain (3.9) and complete the proof. ∎
 4. Principal part of solutions and perturbed forces
-In this section, we construct the principal part of the non-unique steady flow outlined in Section 1.3, and we prove that the perturbative forcing term admits better regularity and smallness than the principal part.
-The central aim of this section is to establish the following theorem.
+In this section, we construct the principal part of the non-unique steady flow outlined in Section 1.3, and prove that the perturbative forcing term admits better regularity and smallness than the principal part.
+The central aim of this section is to show the following theorem.
 Theorem 4.1.
-For $n\geqslant 3$, there exists a positive constant $\lambda_{*}>0$ such that for every $\lambda\in 2^{\mathbb{N}}$ with $\lambda\geqslant\lambda_{*}$, there exists a divergence-free vector field $V_{\lambda}\in\mathscr{S}^{\prime}(\mathbb{R}^{n})$ such that the following (i) and (ii) holds:
+For $n\geqslant 3$, there exists a constant $\lambda_{*}>0$ such that for every $\lambda\in 2^{\mathbb{N}}$ with $\lambda\geqslant\lambda_{*}$, there exists a divergence-free vector field $V_{\lambda}\in\mathscr{S}^{\prime}(\mathbb{R}^{n})$ such that the following (i) and (ii) hold true.
   1. (i)
-    For any $p$ and $q$ satisfying (N1) or (N2) in Theorem 1.2, $V_{\lambda}$ belongs to $\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ and it holds
+    For any $(p,q)$ satisfying (N1) or (N2) in Theorem 1.2, $V_{\lambda}$ belongs to $\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ and it holds
     $C^{-1}\lambda^{-\alpha(p)}\leqslant{\left\|V_{\lambda}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\leqslant C\lambda^{-\alpha(p)}$ (4.1)
-    with $\alpha(p):=(3-n/p)(2n+4)-3$ and a positive constant $C=C(n,p,q)$ independent of $\lambda$.
+    with $\alpha(p):=(3-n/p)(4n+5)-4-n/p$ and some constant $C=C(n,p,q)>0$.
   2. (ii)
@@
     ${\left\|F_{\lambda}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-3}}\leqslant C\lambda^{-\beta}$ (4.3)
-    with $\beta:=6n+10$ and a positive constant $C=C(n,r)$ independent of $\lambda$.
+    with $\beta:=12n+12$ and some constant $C=C(n,r)>0$.
 Remark 4.2.
-Note that $\alpha(p)\leqslant\alpha(\infty)=6n+9<\beta$, which implies $F_{\lambda}$ is much smaller than $V_{\lambda}$.
+Since $\alpha(p)\leqslant\alpha(\infty)=12n+11<\beta$ and $\dot{B}_{r,1}^{n/r-1}(\mathbb{R}^{n})\hookrightarrow\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$, $F_{\lambda}$ is much smaller and more regular than $V_{\lambda}$.
 4.1. Notations and key lemmas
-Before proving Theorem 4.1, we collect several auxiliary results and notations. We begin by recalling the well-known geometric lemma known as the Nash lemma.
+Before proving Theorem 4.1, we collect several auxiliary results and notations. We begin by recalling the well-known geometric lemma.
 Lemma 4.3 ([Nash-54]).
@@
 For the readers’ convenience, we provide an outline of an elementary proof. Let us define $\mathcal{K}$ by
-$\mathcal{K}:=\left\{\mathbbm{e}_{i}\ ;\ i=1,\dots,n\right\}$ $\cup\left\{\mathbbm{e}_{i}+\mathbbm{e}_{j}\ ;\ i,j=1,\dots,n,\ i\neq j\right\}$ (4.5)
-$\cup\left\{\mathbbm{e}_{i}-\mathbbm{e}_{j}\ ;\ i,j=1,\dots,n,\ i\neq j\right\},$ (4.6)
+$\mathcal{K}:=\left\{\mathbbm{e}_{i}\ ;\ i=1,\dots,n\right\}$ $\cup\left\{\mathbbm{e}_{i}+\mathbbm{e}_{j}\ ;\ i,j=1,\dots,n,\ i<j\right\}$ (4.5)
+$\cup\left\{\mathbbm{e}_{i}-\mathbbm{e}_{j}\ ;\ i,j=1,\dots,n,\ i<j\right\},$ (4.6)
 where $\mathbbm{e}_{1},\dots,\mathbbm{e}_{n}\in\mathbb{R}^{n}$ stands for the $i$-th standard basis vector. We define $\{\Gamma_{k}\}_{k\in\mathcal{K}}$ by
-$\Gamma_{k}(M):=\begin{cases}\left(M_{ii}-\dfrac{1}{2}\right)^{\frac{1}{2}},&(k=\mathbbm{e}_{i},\ i=1,\dots,n),\\[7.0pt] \left(\dfrac{1}{4(n-1)}\pm\dfrac{M_{ij}}{2}\right)^{\frac{1}{2}},&(k=\mathbbm{e}_{i}\pm\mathbbm{e}_{j},\ i,j=1,\dots,n,\ i\neq j),\end{cases}$ (4.7)
+$\Gamma_{k}(M):=\begin{cases}\left(M_{ii}-\dfrac{1}{2}\right)^{\frac{1}{2}},&(k=\mathbbm{e}_{i},\ i=1,\dots,n),\\[7.0pt] \left(\dfrac{1}{4(n-1)}\pm\dfrac{M_{ij}}{2}\right)^{\frac{1}{2}},&(k=\mathbbm{e}_{i}\pm\mathbbm{e}_{j},\ i,j=1,\dots,n,\ i<j),\end{cases}$ (4.7)
 for all $M=(M_{ij})\in\mathcal{S}_{n}$ with $|M-{\rm Id}|\leqslant r_{0}:=(4(n-1))^{-1}$. Then, we may easily check that (4.4) holds true, and complete the proof. ∎
-Here, we introduce some notation that will be used throughout this section.
+Here, we introduce some notations that are to be used throughout this section.
   1. 1.
-    Let $\{a_{k}\}_{k\in\mathcal{K}}\subset\mathbb{R}^{n}\setminus\{0\}$ fulfill $k\cdot a_{k}=0$ and $|a_{k}|\in 2^{\mathbb{N}}$ for all $k\in\mathcal{K}$, and also satisfy that $a_{k}$ and $a_{k^{\prime}}$ are not parallel unless $k=k^{\prime}$.
-For instance, we may choose
-    $a_{k}:=\begin{cases}\mathbbm{e}_{i+1},&(k=\mathbbm{e}_{i},\ i=1,\dots,n-1),\\ \mathbbm{e}_{1},&(k=\mathbbm{e}_{n}),\\ \sqrt{2}\left(\mathbbm{e}_{i}\mp\mathbbm{e}_{j}\right),&(k=\mathbbm{e}_{i}\pm\mathbbm{e}_{j},\ i,j=1,\dots,n,\ i\neq j).\end{cases}$ (4.8)
+    Let $\{a_{k}\}_{k\in\mathcal{K}}\subset\mathbb{S}^{n-1}$ satisfy that $a_{k}$ and $a_{k^{\prime}}$ are not parallel unless $k=k^{\prime}$. For instance, we may choose
+    $a_{k}:=\begin{cases}\mathbbm{e}_{i+1},&(k=\mathbbm{e}_{i},\ i=1,\dots,n-1),\\ \mathbbm{e}_{1},&(k=\mathbbm{e}_{n}),\\ \sqrt{2}\left(\mathbbm{e}_{i}\mp\mathbbm{e}_{j}\right),&(k=\mathbbm{e}_{i}\pm\mathbbm{e}_{j},\ i,j=1,\dots,n,\ i<j).\end{cases}$ (4.8)
   2. 2.
-    Let $\mu:=2n+4$. For $\lambda\in 2^{\mathbb{N}}$, we define $\{\lambda_{j}\}_{j\in\mathbb{N}_{0}}$ inductively by
+    Let $\mu=4n+5$. For $\lambda\in 2^{\mathbb{N}}$, we define $\{\lambda_{j}\}_{j\in\mathbb{N}_{0}}$ inductively by
     $\lambda_{0}=\lambda,\qquad\lambda_{j}=\exp_{\lambda}\!\left(\exp_{2}\!\left(\lambda_{j-1}^{4\mu}\right)\right),$ (4.9)
@@
 We emphasize that the required lower bound on $\lambda$ is chosen independently of $j$.
-Moreover, almost all quantities introduced below depend on the parameter $\lambda$; however, we do not record this dependence explicitly in the notation so as to avoid unnecessarily complicating it.
+Moreover, almost all quantities introduced below depend on the parameter $\lambda$; however, we do not record this dependence explicitly in their notations so as to avoid unnecessarily complicating them.
   3. 3.
@@
     $\Lambda_{j}:=\left\{\ell\in\mathbb{N}\ ;\ \mu\log_{\lambda}\lambda_{j-1}\leqslant\ell\leqslant\log_{\lambda}\lambda_{j}\right\}.$ (4.10)
-    In what follows, we choose $\lambda$ sufficiently large so that $\max\Lambda_{j}<\min\Lambda_{j+1}$ holds for all $j\in\mathbb{N}_{0}$.
+    In what follows, we choose $\lambda$ sufficiently large so that $\Lambda_{j}\neq\varnothing$ for all $j\in\mathbb{N}$.
   4. 4.
@@
     $h_{j}:=\sum_{\ell\in\Lambda_{j}}\frac{1}{\ell}.$ (4.11)
-    Note that $h_{j}\sim\log_{2}\log_{\lambda}\lambda_{j}=\lambda_{j-1}^{4\mu}$.
+    Note that there holds
+    $h_{j}\sim{}$ $\log_{2}\frac{\log_{\lambda}\lambda_{j}}{\mu\log_{\lambda}\lambda_{j-1}}$ (4.12)
+    $={}$ $\begin{cases}\lambda_{j-1}^{4\mu}-\log_{2}\mu&(j=1),\\ \lambda_{j-1}^{4\mu}-\lambda_{j-2}^{4\mu}-\log_{2}\mu,\quad&(j=2,3,4,\dots),\end{cases}$ (4.13)
+    which yields $h_{j}\sim\lambda_{j-1}^{4\mu}$.
   5. 5.
-    We choose a $\rho\in C_{c}^{\infty}(\mathbb{R}^{n};[0,\infty))$ so that
+    We choose a function $\rho\in C_{c}^{\infty}(\mathbb{R}^{n};[0,\infty))$ so that
     $\operatorname{supp}\rho=B_{\mathbb{R}^{n}}(0,1),\qquad\int_{\mathbb{R}^{n}}\rho(y)dy=1.$ (4.14)
@@
     $\mathcal{D}v:=-\nabla v-(\nabla v)^{\top}+2(\operatorname{div}v){\rm Id}$ (4.17)
-    for smooth vector field $v:\mathbb{R}^{n}\to\mathbb{R}^{n}$. Note that $\mathcal{D}v$ is a real symmetric matrix and it holds
+    for all smooth vector field $v:\mathbb{R}^{n}\to\mathbb{R}^{n}$. Note that $\mathcal{D}v$ is a real symmetric matrix and it holds
     $\operatorname{div}\mathcal{D}v=-\Delta v+\nabla\operatorname{div}v=-\Delta\mathbb{P}v.$ (4.18)
-In order to define the elements of the desired $V_{\lambda}$, we introduce the following operators.
+In order to define the building blocks of the desired $V_{\lambda}$, we introduce the following operators.
 Definition 4.4.
-For $j\in\mathbb{N}$ and $v\in C_{c}^{\infty}(\mathbb{R}^{n})$ satisfying ${\left\|\mathcal{D}v\right\|}_{L^{\infty}}\leqslant r_{0}\lambda_{j-1}^{6}$, we define
-${\mathcal{V}}_{j}[v]:=\sqrt{2}\lambda_{j-1}^{3}\lambda_{j}^{-2}\sum_{k\in\mathcal{K}}{\rho}_{j}*\operatorname{div}\mathcal{D}\left({\Gamma}_{j,k}[v]{\Psi}_{j,k}k\right),$ (4.17)
+For $j\in\mathbb{N}$ and $v\in C_{c}^{\infty}(\mathbb{R}^{n})$ satisfying ${\left\|\mathcal{D}v\right\|}_{L^{\infty}}\leqslant r_{0}\lambda_{j-1}^{8}$, we define
+${\mathcal{V}}_{j}[v]:=\sqrt{2}\lambda_{j-1}^{4}\rho_{j}*\sum_{k\in\mathcal{K}}\operatorname{div}\mathcal{D}\left({\Gamma}_{j,k}[v]{\Psi}_{j,k}k\right),$ (4.19)
 where we have set
-${\Gamma}_{j,k}[v](x):=\Gamma_{k}\left({\rm Id}-\lambda_{j-1}^{-6}\mathcal{D}v(x)\right),$ (4.18)
-${\Psi}_{j,k}(x):=\frac{1}{\sqrt{h_{j}}}\sum_{\ell\in\Lambda_{j}}\frac{1}{\sqrt{\ell}}\psi_{j}(x)\cos\left(\lambda^{\ell}a_{k}\cdot x\right).$ (4.19)
-Let us establish elementary derivative estimates for each $\mathcal{V}_{j}[v]$.
+${\Gamma}_{j,k}[v](x):=\Gamma_{k}\left({\rm Id}-\lambda_{j-1}^{-8}\mathcal{D}v(x)\right),$ (4.20)
+${\Psi}_{j,k}(x):=\sum_{\ell\in\Lambda_{j}}\frac{1}{\lambda^{2\ell}\sqrt{h_{j}\ell}}\psi_{j}(x)\cos\left(\lambda^{\ell}a_{k}\cdot x\right).$ (4.21)
+Remark 4.5.
+When $\mathbb{R}^{n}$ is replaced by $\mathbb{T}^{n}$, it is not necessary to multiply by $\psi_{j}$ in the definition of $\Psi_{j,k}$.
+Let us establish elementary derivative estimates for each $\mathcal{V}_{j}[v]$, which rely on the mollifier.
 Although these bounds are far from optimal, they ensure the well-definedness of Definition 4.7 below, and they provide an essential foundation for obtaining the sharp estimates later on.
@@
 For every $m\in\mathbb{N}_{0}$, there exists a positive constant $C=C(m)$ such that
-${\left\|\nabla^{m}{\mathcal{V}}_{j}[v]\right\|}_{L^{p}}\leqslant C\lambda_{j}^{2m+3}.$ (4.20)
-for all $j\in\mathbb{N}$, $1\leqslant p\leqslant\infty$, and $v\in C_{c}^{\infty}(\mathbb{R}^{n})$ with ${\left\|\mathcal{D}v\right\|}_{L^{\infty}}\leqslant r_{0}\lambda_{j-1}^{6}$. Moreover, it holds
-${\left\|\mathcal{D}{\mathcal{V}}_{j}[v]\right\|}_{L^{\infty}}\leqslant r_{0}\lambda_{j}^{6}.$ (4.21)
+${\left\|\nabla^{m}{\mathcal{V}}_{j}[v]\right\|}_{L^{p}}\leqslant C\lambda_{j}^{2m+4}.$ (4.22)
+for all $j\in\mathbb{N}$, $1\leqslant p\leqslant\infty$, and $v\in C_{c}^{\infty}(\mathbb{R}^{n})$ with ${\left\|\mathcal{D}v\right\|}_{L^{\infty}}\leqslant r_{0}\lambda_{j-1}^{8}$. Moreover, it holds
+${\left\|\mathcal{D}{\mathcal{V}}_{j}[v]\right\|}_{L^{\infty}}\leqslant r_{0}\lambda_{j}^{8}.$ (4.23)
 Proof.
 From the Young inequality, we see that
-${\left\|\nabla^{m}{\mathcal{V}}_{j}[v]\right\|}_{L^{p}}$ $\leqslant C\lambda_{j-1}^{3}\lambda_{j}^{-2}{\left\|\nabla^{m+2}\rho_{k}\right\|}_{L^{1}}\sum_{k\in\mathcal{K}}{\left\|\Gamma_{j,k}[v]\Psi_{j,k}\right\|}_{L^{p}}$ (4.22)
-$\leqslant C\lambda_{j-1}^{3}\lambda_{j}^{2m+2}\sum_{k\in\mathcal{K}}{\left\|\Psi_{j,k}\right\|}_{L^{p}}.$ (4.23)
+${\left\|\nabla^{m}{\mathcal{V}}_{j}[v]\right\|}_{L^{p}}$ $\leqslant C\lambda_{j-1}^{4}{\left\|\nabla^{m+2}\rho_{k}\right\|}_{L^{1}}\sum_{k\in\mathcal{K}}{\left\|\Gamma_{j,k}[v]\Psi_{j,k}\right\|}_{L^{p}}$ (4.24)
+$\leqslant C\lambda_{j-1}^{4}\lambda_{j}^{2m+4}\sum_{k\in\mathcal{K}}{\left\|\Psi_{j,k}\right\|}_{L^{p}}.$ (4.25)
 Since $\Psi_{j,k}$ is bounded as
-${\left\|\Psi_{j,k}\right\|}_{L^{p}}$ $\leqslant\frac{C}{\sqrt{h_{j}}}\lambda^{\frac{n}{p}j}\sum_{\ell=1}^{\log_{\lambda}\lambda_{j}}\frac{1}{\sqrt{\ell}}$ (4.24)
-$\leqslant C\lambda^{-2\mu}_{j-1}\lambda^{\frac{n}{p}j}\sqrt{\log_{\lambda}\lambda_{j}}$ (4.25)
-$\leqslant C\lambda^{-2\mu}_{j-1}\lambda_{j},$ (4.26)
+${\left\|\Psi_{j,k}\right\|}_{L^{p}}\leqslant\sum_{\ell\in\Lambda_{j}}\frac{C\lambda^{\frac{n}{p}j}}{\lambda^{2\ell}\sqrt{h_{j}\ell}}\leqslant\frac{C\lambda^{\frac{n}{p}j}}{\lambda^{4\mu}_{j-1}\sqrt{\mu\log_{\lambda}\lambda_{j-1}}}\leqslant C\lambda^{-4\mu}_{j-1}$ (4.26)
 we obtain (4.22). For the estimate of (4.23), (4.22) with $p=\infty$ and $m=1$ yields
-${\left\|\mathcal{D}\mathcal{V}[v]\right\|}_{L^{\infty}}\leqslant C{\left\|\nabla\mathcal{V}[v]\right\|}_{L^{\infty}}\leqslant C\lambda_{j}^{5}\leqslant r_{0}\lambda_{j}^{6}$ (4.27)
+${\left\|\mathcal{D}\mathcal{V}[v]\right\|}_{L^{\infty}}\leqslant C{\left\|\nabla\mathcal{V}[v]\right\|}_{L^{\infty}}\leqslant C\lambda_{j}^{6}\leqslant r_{0}\lambda_{j}^{8}$ (4.27)
 provided that $\lambda$ is sufficiently large. Thus, we complete the proof. ∎
-We now define, by induction, a sequence of flows that will serve as building blocks for the velocity field required in Theorem 4.1.
+We now define the building blocks for the velocity field required in Theorem 4.1.
 Definition 4.7.
@@
   1. 1.
-    Note that (4.21) guarantees that $V_{j-1}$ belongs to the domain of the operator $\mathcal{V}_{j}[\cdot]$, and hence the sequence $\{V_{j}\}_{j\in\mathbb{N}_{0}}$ is well-defined.
+    Note that (4.23) guarantees that $V_{j-1}$ belongs to the domain of the operator $\mathcal{V}_{j}$, and hence the sequence $\{V_{j}\}_{j\in\mathbb{N}_{0}}$ is well-defined.
   2. 2.
-    By the definition of $\mathcal{V}_{j}$, we see that each $V_{j}$ is solenoidal.
+    By $\operatorname{div}\mathcal{D}=-\Delta\mathbb{P}$ in the definition of $\mathcal{V}_{j}$, we see that each $V_{j}$ is solenoidal.
   3. 3.
@@
 To investigate the precise structure of $\{V_{j}\}_{j\in\mathbb{N}_{0}}$, we decompose it into the principal and remainder part as
-$V_{j}=V^{(\rm p)}_{j}+V^{({\rm r})}_{j}$, (4.32)
+$V_{j}=V^{(\rm p)}_{j}+V^{({\rm r})}_{j},\qquad V^{(\rm r)}_{j}=V^{({\rm r},1)}_{j}+V^{({\rm r},2)}_{j}+V^{({\rm r},3)}_{j}$, (4.32)
 where we have set
-$V^{(\rm p)}_{j}(x):={}\sqrt{2}\lambda_{j-1}^{3}\sum_{k\in\mathcal{K}}\Gamma_{j,k}[V_{j-1}](x)\Psi_{j,k}(x)k$ (4.33)
+$V^{(\rm p)}_{j}(x):={}$ $\sqrt{2}\lambda_{j-1}^{4}\sum_{k\in\mathcal{K}}\sum_{\ell\in\Lambda_{j}}\frac{1}{\lambda^{2\ell}\sqrt{h_{j}\ell}}\Gamma_{j,k}[V_{j-1}](x)\psi_{j}(x)\operatorname{div}\mathcal{D}\left(k\cos\left(\lambda^{\ell}a_{k}\cdot x\right)\right)$ (4.33)
+$={}$ $\sqrt{2}\lambda_{j-1}^{4}\sum_{k\in\mathcal{K}}\sum_{\ell\in\Lambda_{j}}\frac{1}{\sqrt{h_{j}\ell}}\Gamma_{j,k}[V_{j-1}](x)\psi_{j}(x)k\cos\left(\lambda^{\ell}a_{k}\cdot x\right)$ (4.34)
 and
-$V^{(\rm r)}_{j}=V^{({\rm r},1)}_{j}+V^{({\rm r},2)}_{j}$, (4.34)
-$\quad V^{({\rm r},1)}_{j}(x):={}\rho_{j}*V^{(\rm p)}_{j}(x)-V^{(\rm p)}_{j}(x),$ (4.35)
-$\quad V^{({\rm r},2)}_{j}(x):={}\sqrt{2}\lambda_{j-1}^{3}\lambda_{j}^{-2}\rho_{j}*\sum_{k\in\mathcal{K}}\operatorname{div}\mathcal{D}\left(\Gamma_{j,k}[V_{j-1}](x)\psi_{j}(x)k\right)\cos\left(\lambda_{j}a_{k}\cdot x\right).$ (4.36)
+$V^{({\rm r},1)}_{j}(x):={}\rho_{j}*V^{(\rm p)}_{j}(x)-V^{(\rm p)}_{j}(x),$ (4.35)
+$\begin{aligned} V^{({\rm r},2)}_{j}(x):={}&\sqrt{2}\lambda_{j-1}^{4}\rho_{j}*\sum_{k\in\mathcal{K}}\sum_{\ell\in\Lambda_{j}}\frac{1}{\lambda^{\ell}\sqrt{h_{j}\ell}}\mathcal{D}\left(\Gamma_{j,k}[V_{j-1}]\psi_{j}k\right)\nabla\cos\left(\lambda^{\ell}a_{k}\cdot x\right)\\ ={}&-\sqrt{2}\lambda_{j-1}^{4}\rho_{j}*\sum_{k\in\mathcal{K}}\sum_{\ell\in\Lambda_{j}}\frac{1}{\lambda^{\ell}\sqrt{h_{j}\ell}}\mathcal{D}\left(\Gamma_{j,k}[V_{j-1}]\psi_{j}k\right)a_{k}\sin\left(\lambda^{\ell}a_{k}\cdot x\right),\end{aligned}$ (4.36)
+$V^{({\rm r},3)}_{j}(x):={}\sqrt{2}\lambda_{j-1}^{4}\rho_{j}*\sum_{k\in\mathcal{K}}\sum_{\ell\in\Lambda_{j}}\frac{1}{\lambda^{2\ell}\sqrt{h_{j}\ell}}\operatorname{div}\mathcal{D}\left(\Gamma_{j,k}[V_{j-1}]\psi_{j}k\right)\cos\left(\lambda^{\ell}a_{k}\cdot x\right).$ (4.37)
+We also introduce a function
+$\Phi_{j,k,\ell}(x):=\frac{1}{\sqrt{h_{j}\ell}}\Gamma_{j,k}[V_{j-1}](x)\psi_{j}(x)k\cos\left(\lambda^{\ell}a_{k}\cdot x\right)$ (4.38)
+so that $V^{({\rm p})}_{j}$ has the following simplified representation:
+$V^{(\rm p)}_{j}(x)=\sqrt{2}\lambda_{j-1}^{4}\sum_{k\in\mathcal{K}}\sum_{\ell\in\Lambda_{j}}\Phi_{j,k,\ell}(x).$ (4.39)
 Lemma 4.9.
-Let $1\leqslant p,q\leqslant\infty$ and $s\in\mathbb{R}$ satisfy
+Let $1\leqslant p,q\leqslant\infty$ and let $s\in\mathbb{R}$ satisfy
 $\frac{n}{p}-n<s<2n.$ (4.40)
 Then, there exists a positive constant $C=C(n,p,q,s)$ such that
-$C^{-1}\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q)\leqslant{\left\|V^{({\rm p})}_{j}\right\|}_{\dot{B}_{p,q}^{s}}\leqslant C\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q),$ (4.38)
+$C^{-1}\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-2\mu}\sigma_{j}(s,q)\leqslant{\left\|V^{({\rm p})}_{j}\right\|}_{\dot{B}_{p,q}^{s}}\leqslant C\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-2\mu}\sigma_{j}(s,q),$ (4.41)
 and
-${\left\|V^{({\rm r})}_{j}\right\|}_{\dot{B}_{p,q}^{s}}\leqslant C\lambda_{j}^{-1}\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q)$ (4.39)
+${\left\|V^{({\rm r})}_{j}\right\|}_{\dot{B}_{p,q}^{s}}\leqslant C\lambda^{nj}\lambda_{j-1}^{6-2\mu}\sigma_{j}(s-1,q)$ (4.42)
 for all $j\in\mathbb{N}$, where the sequence $\{\sigma_{j}(s,q)\}_{j\in\mathbb{N}}$ have been defined by
@@
 Since $\Gamma_{j,k}[V_{j-1}]$ is constant for $j=1$, we focus on the case of $j\geqslant 2$.
-Making use of (4.20) for $V_{j-1}=\mathcal{V}_{j-1}[V_{j-2}]$, it holds ${\left\|\nabla^{s}V_{j-1}\right\|}_{L^{\infty}}\leqslant C\lambda_{j-1}^{2s+3}$ ($s\in\mathbb{N}_{0}$).
+Making use of (4.22) for $V_{j-1}=\mathcal{V}_{j-1}[V_{j-2}]$, it holds ${\left\|\nabla^{s}V_{j-1}\right\|}_{L^{\infty}}\leqslant C\lambda_{j-1}^{2s+4}$ ($s\in\mathbb{N}_{0}$).
 Thus, from the Faà di Bruno formula, it follows that
-${\left\|\nabla^{m}\Gamma_{j,k}[V_{j-1}]\right\|}_{L^{\infty}}$ $\leqslant C\sum_{\ell=1}^{m}\sum_{s_{1}+\dots+s_{\ell}=m}\lambda_{j-1}^{-6\ell}{\left\|\nabla^{s_{1}}\mathcal{D}V_{j-1}\right\|}_{L^{\infty}}\cdots{\left\|\nabla^{s_{\ell}}\mathcal{D}V_{j-1}\right\|}_{L^{\infty}}$ (4.41)
-$\leqslant C\sum_{\ell=1}^{m}\sum_{s_{1}+\dots+s_{\ell}=m}\lambda_{j-1}^{-6\ell}\cdot{\lambda_{j-1}^{2(s_{1}+1)+3}}\cdots{\lambda_{j-1}^{2(s_{\ell}+1)+3}}$ (4.42)
-$=C\lambda_{j-1}^{2m}\sum_{\ell=1}^{m}\lambda_{j-1}^{-\ell}$ (4.43)
+${\left\|\nabla^{m}\Gamma_{j,k}[V_{j-1}]\right\|}_{L^{\infty}}$ $\leqslant C\sum_{\ell=1}^{m}\sum_{s_{1}+\dots+s_{\ell}=m}\lambda_{j-1}^{-8\ell}{\left\|\nabla^{s_{1}}\mathcal{D}V_{j-1}\right\|}_{L^{\infty}}\cdots{\left\|\nabla^{s_{\ell}}\mathcal{D}V_{j-1}\right\|}_{L^{\infty}}$ (4.44)
+$\leqslant C\sum_{\ell=1}^{m}\sum_{s_{1}+\dots+s_{\ell}=m}\lambda_{j-1}^{-8\ell}\cdot{\lambda_{j-1}^{2(s_{1}+1)+4}}\cdots{\lambda_{j-1}^{2(s_{\ell}+1)+4}}$ (4.45)
 $\leqslant C\lambda_{j-1}^{2m}.$ (4.46)
 Let us prove (4.41). It follows from Proposition 3.1 with $M=2n$ that
-${\left\|V^{({\rm p})}_{j}\right\|}_{\dot{B}_{p,q}^{s}}$ $\leqslant C\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q)\sum_{k\in\mathcal{K}}|a_{k}|^{s}{\left\|\Gamma_{j,k}[V_{j-1}]\psi_{j}\right\|}_{L^{p}}$ (4.45)
-$\quad+C\lambda_{j-1}^{3-2\mu}\sum_{\ell\in\Lambda_{j}}\frac{\lambda^{(s-n)\ell}}{\sqrt{\ell}}\sum_{k\in\mathcal{K}}|a_{k}|^{s-n}{\left\|\nabla^{2n}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}\cap L^{1}}$ (4.46)
-$\leqslant C\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q)\left(1+\lambda_{j-1}^{(4-\mu)n}\right)$ (4.47)
-$\leqslant C\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q),$ (4.48)
+${\left\|V^{({\rm p})}_{j}\right\|}_{\dot{B}_{p,q}^{s}}$ $\leqslant C\lambda_{j-1}^{4-2\mu}\sigma_{j}(s,q)\sum_{k\in\mathcal{K}}{\left\|\Gamma_{j,k}[V_{j-1}]\psi_{j}\right\|}_{L^{p}}$ (4.47)
+$\quad+C\lambda_{j-1}^{4-2\mu}\sum_{\ell\in\Lambda_{j}}\frac{\lambda^{(s-n)\ell}}{\sqrt{\ell}}\sum_{k\in\mathcal{K}}{\left\|\nabla^{2n}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}\cap L^{1}}$ (4.48)
+$\leqslant C\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-2\mu}\sigma_{j}(s,q)\left(1+\lambda^{n(1-\frac{n}{p})j}\lambda_{j-1}^{(4-\mu)n}\right)$ (4.49)
+$\leqslant C\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-2\mu}\sigma_{j}(s,q),$ (4.50)
 where we have used
@@
 and
-${\left\|\nabla^{2n}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}\cap L^{1}}\leqslant C\lambda_{j-1}^{4n}.$ (4.50)
+${\left\|\nabla^{2n}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}\cap L^{1}}\leqslant C\lambda^{nj}\lambda_{j-1}^{4n}.$ (4.52)
 For the estimate of $V_{j}^{({\rm p})}$ from below, it holds by Proposition 3.1 and similar calculations as above that
-${\left\|V^{({\rm p})}_{j}\right\|}_{\dot{B}_{p,q}^{s}}$ $\geqslant C^{-1}\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q)\max_{k\in\mathcal{K}}|a_{k}|^{s}{\left\|\Gamma_{j,k}[V_{j-1}]\psi_{j}\right\|}_{L^{p}}$ (4.51)
-$\quad-C\lambda_{j-1}^{3-2\mu}\sum_{\ell\in\Lambda_{j}}\frac{\lambda^{(s-1)\ell}}{\sqrt{\ell}}\max_{k\in\mathcal{K}}|a_{k}|^{s-1}{\left\|\nabla(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}}$ (4.52)
-$\quad-C\lambda_{j-1}^{3-2\mu}\sum_{\ell\in\Lambda_{j}}\frac{\lambda^{(s-n)\ell}}{\sqrt{\ell}}\max_{k\in\mathcal{K}}|a_{k}|^{s-n}{\left\|\nabla^{2n}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}\cap L^{1}}$ (4.53)
-$\geqslant C^{-1}\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q)\left(1-C^{2}\lambda_{j-1}^{2-\mu}-C^{2}\lambda_{j-1}^{(4-\mu)n}\right)$ (4.54)
-$\geqslant C^{-1}\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q).$ (4.55)
+${\left\|V^{({\rm p})}_{j}\right\|}_{\dot{B}_{p,q}^{s}}$ $\geqslant C^{-1}\lambda_{j-1}^{4-2\mu}\sigma_{j}(s,q)\max_{k\in\mathcal{K}}{\left\|\Gamma_{j,k}[V_{j-1}]\psi_{j}\right\|}_{L^{p}}$ (4.53)
+$\quad-C\lambda_{j-1}^{4-2\mu}\sum_{\ell\in\Lambda_{j}}\frac{\lambda^{(s-1)\ell}}{\sqrt{\ell}}\sum_{k\in\mathcal{K}}{\left\|\nabla(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}}$ (4.54)
+$\quad-C\lambda_{j-1}^{4-2\mu}\sum_{\ell\in\Lambda_{j}}\frac{\lambda^{(s-n)\ell}}{\sqrt{\ell}}\sum_{k\in\mathcal{K}}{\left\|\nabla^{2n}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}\cap L^{1}}$ (4.55)
+$\geqslant C^{-1}\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-2\mu}\sigma_{j}(s,q)\left(1-C^{2}\lambda^{n(1-\frac{1}{p})j}\lambda_{j-1}^{2-\mu}-C^{2}\lambda^{n(1-\frac{1}{p})j}\lambda_{j-1}^{(4-\mu)n}\right)$ (4.56)
+$\geqslant C^{-1}\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-2\mu}\sigma_{j}(s,q).$ (4.57)
 Next, we show (4.42). It follows from $\int_{\mathbb{R}^{n}}\rho_{j}(y)dy=1$ and the mean value formula that
@@
 ${\left\|V^{({\rm r},1)}_{j}\right\|}_{\dot{B}_{p,q}^{s}}$ $\leqslant C\lambda_{j}^{-2}{\left\|V^{({\rm p})}_{j}\right\|}_{\dot{B}_{p,q}^{s+1}}$ (4.61)
-$\leqslant C\lambda_{j}^{-2}\lambda_{j-1}^{3-2\mu}\sigma_{j}(s+1,q)$ (4.60)
-$\leqslant C\lambda_{j}^{-1}\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q).$ (4.61)
+$\leqslant C\lambda_{j}^{-2}\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-2\mu}\sigma_{j}(s+1,q)$ (4.62)
+$\leqslant C\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-2\mu}\sigma_{j}(s-1,q).$ (4.63)
 We follow the same argument as in the proof of (4.41) to obtain
-${\left\|V^{({\rm r},2)}_{j}\right\|}_{\dot{B}_{p,q}^{s}}$ $\leqslant C\lambda_{j}^{-2}\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q)\sum_{k\in\mathcal{K}}|a_{k}|^{s}{\left\|\nabla^{2}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}}$ (4.62)
-$\quad+C\lambda_{j}^{-2}\lambda_{j-1}^{3-2\mu}\sum_{\ell\in\Lambda_{j}}\frac{\lambda^{(s-n)\ell}}{\sqrt{\ell}}\sum_{k\in\mathcal{K}}|a_{k}|^{s-n}{\left\|\nabla^{2n+2}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}\cap L^{1}}$ (4.63)
-$\leqslant C\lambda_{j}^{-2}\lambda_{j-1}^{7-2\mu}\sigma_{j}(s,q)\left(1+\lambda_{j-1}^{(4-\mu)n}\right)$ (4.64)
-$\leqslant C\lambda_{j}^{-2}\lambda_{j-1}^{7-2\mu}\sigma_{j}(s,q)$ (4.65)
-$\leqslant C\lambda_{j}^{-1}\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q).$ (4.66)
-This completes the proof. ∎
+${\left\|V^{({\rm r},2)}_{j}\right\|}_{\dot{B}_{p,q}^{s}}$ $\leqslant C\lambda_{j-1}^{4-2\mu}\sigma_{j}(s-1,q)\sum_{k\in\mathcal{K}}{\left\|\nabla(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}}$ (4.64)
+$\quad+C\lambda_{j-1}^{4-2\mu}\sum_{\ell\in\Lambda_{j}}\frac{\lambda^{(s-1-n)\ell}}{\sqrt{\ell}}\sum_{k\in\mathcal{K}}{\left\|\nabla^{2n+1}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}\cap L^{1}}$ (4.65)
+$\leqslant C\lambda^{nj}\lambda_{j-1}^{6-2\mu}\sigma_{j}(s-1,q)\left(1+\lambda_{j-1}^{(4-\mu)n}\right)$ (4.66)
+$\leqslant C\lambda^{nj}\lambda_{j-1}^{6-2\mu}\sigma_{j}(s-1,q)$ (4.67)
+and
+${\left\|V^{({\rm r},3)}_{j}\right\|}_{\dot{B}_{p,q}^{s}}$ $\leqslant C\lambda_{j-1}^{4-2\mu}\sigma_{j}(s-2,q)\sum_{k\in\mathcal{K}}{\left\|\nabla^{2}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}}$ (4.68)
+$\quad+C\lambda_{j-1}^{4-2\mu}\sum_{\ell\in\Lambda_{j}}\frac{\lambda^{(s-2-n)\ell}}{\sqrt{\ell}}\sum_{k\in\mathcal{K}}{\left\|\nabla^{2n+2}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}\cap L^{1}}$ (4.69)
+$\leqslant C\lambda^{nj}\lambda_{j-1}^{8-2\mu}\sigma_{j}(s-2,q)\left(1+\lambda_{j-1}^{(4-\mu)n}\right)$ (4.70)
+$\leqslant C\lambda^{nj}\lambda_{j-1}^{8-3\mu}\sigma_{j}(s-1,q).$ (4.71)
+Gathering (4.63), (4.67), and (4.71), we obtain (4.42) and complete the proof. ∎
 4.2. Proof of Theorem 4.1
 Now, we are in a position to prove Theorem 4.1. We separate the proof into two steps.
-Step 1. Construction and estimates of the principal stationary flow. Let us define
-$\widetilde{\sigma}_{j}(s,q):={\left\|\left\{\sigma_{j}(s,q)\right\}_{j^{\prime}}\right\|}_{\ell^{q}(j^{\prime}\geqslant j)},\qquad j\in\mathbb{N}_{0}.$ (4.67)
-It follows from Lemmas 4.8 that
-$\sum_{j=1}^{\infty}{\left\|V_{j}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}$ $\leqslant C\sum_{j=1}^{\infty}\lambda_{j-1}^{3-2\mu}\sigma_{j}\left(\frac{n}{p}-1,q\right)$ (4.68)
-$\leqslant C\lambda^{3-2\mu}\widetilde{\sigma}_{1}\left(\frac{n}{p}-1,q\right)$ (4.69)
-$\leqslant C\lambda^{3-2\mu}\frac{\lambda^{(\frac{n}{p}-1)\mu}}{\sqrt{\mu}}=C\lambda^{3-(3-\frac{n}{p})\mu}.$ (4.70)
+Step 1. Construction and estimates of the principal stationary flow. It follows from Lemmas 4.9 that
+${\left\|V_{j}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}$ $\leqslant C\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-2\mu}\sigma_{j}\left(\frac{n}{p}-1,q\right)$ (4.72)
+$\leqslant C\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-2\mu}\frac{\lambda_{j-1}^{(\frac{n}{p}-1)\mu}}{\sqrt{\mu\log_{\lambda}\lambda_{j-1}}}\leqslant C\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-(3-\frac{n}{p})\mu}.$ (4.73)
 Hence, the series
@@
 absolutely converges in the strong topology of $\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ and we see that
-${\left\|V^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\leqslant\sum_{j=1}^{\infty}{\left\|V_{j}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\leqslant C\lambda^{3-(3-\frac{n}{p})\mu}.$ (4.72)
+${\left\|V^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\leqslant\sum_{j=1}^{\infty}{\left\|V_{j}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\leqslant C\lambda^{\frac{n}{p}+4-(3-\frac{n}{p})\mu}.$ (4.75)
 For the estimate of $V^{({\rm p})}$ from below, it holds
-${\left\|V^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\geqslant{\left\|V_{1}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}-\sum_{j=2}^{\infty}{\left\|V_{j}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}.$ (4.73)
-Using (4.38), we have
-${\left\|V_{1}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}$ $\geqslant C^{-1}\lambda^{3-2\mu}\sigma_{1}\left(\frac{n}{p}-1,q\right)$ (4.74)
-$\geqslant C^{-1}\lambda^{3-2\mu}\frac{\lambda^{(\frac{n}{p}-1)\mu}}{\sqrt{\mu}}=C^{-1}\lambda^{3-(3-\frac{n}{p})\mu}.$ (4.75)
-Similarly to (4.68), we see that
-$\sum_{j=2}^{\infty}{\left\|V_{j}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}$ $\leqslant C\sum_{j=2}^{\infty}\lambda_{j-1}^{3-2\mu}\sigma_{j}\left(\frac{n}{p}-1,q\right)$ (4.76)
-$\leqslant C\lambda_{1}^{3-2\mu}\widetilde{\sigma}_{2}\left(\frac{n}{p}-1,q\right)$ (4.77)
-$\leqslant C\lambda_{1}^{3-(3-\frac{n}{p})\mu}.$ (4.78)
-Hence, we obtain
-${\left\|V^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\geqslant C^{-1}\lambda^{3-(3-\frac{n}{p})\mu}\left\{1-C^{2}\left(\frac{\lambda}{\lambda_{1}}\right)^{3-(3-\frac{n}{p})\mu}\right\}\geqslant C^{-1}\lambda^{3-(3-\frac{n}{p})\mu}.$ (4.79)
+${\left\|V^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\geqslant{}$ ${\left\|V_{1}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}-\sum_{j=2}^{\infty}{\left\|V_{j}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}$ (4.76)
+$\geqslant{}$ $C^{-1}\lambda^{\frac{n}{p}}\lambda^{4-2\mu}\sigma_{1}\left(\frac{n}{p}-1,q\right)-C\sum_{j=2}^{\infty}\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-(3-\frac{n}{p})\mu}$ (4.77)
+$\geqslant{}$ $C^{-1}\lambda^{\frac{n}{p}}\lambda^{4-(3-\frac{n}{p})\mu}-C\lambda^{\frac{2n}{p}}\lambda_{1}^{4-(3-\frac{n}{p})\mu}$ (4.78)
+$\geqslant{}$ $C^{-1}\lambda^{\frac{n}{p}+4-(3-\frac{n}{p})\mu}$ (4.79)
 for sufficiently large $\lambda$.
-Next, we consider the remainder part. By Lemma 4.8, there holds
-$\sum_{j=1}^{\infty}{\left\|V_{j}^{({\rm r})}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}$ $\leqslant C\sum_{j=1}^{\infty}\lambda_{j}^{-1}\lambda_{j-1}^{3-2\mu}\sigma_{j}\left(\frac{n}{r}-1,1\right)$ (4.80)
-$\leqslant C\sum_{j=1}^{\infty}\lambda_{j}^{\frac{n}{r}-2}\lambda_{j-1}^{3-2\mu}$ (4.81)
-$\leqslant C\lambda_{1}^{\frac{n}{r}-2}\lambda^{3-2\mu}.$ (4.82)
-Thus, due to $r>n/2$, we see that the series
+Next, we consider the remainder part. Let $\rho>n/2$. Then, similarly as above, we see that the series
 $V^{({\rm r})}:=\sum_{j=1}^{\infty}V_{j}^{({\rm r})}$ (4.80)
-absolutely converges in the strong topology of $\dot{B}_{r,1}^{n/r-1}(\mathbb{R}^{n})$ and
-${\left\|V^{({\rm r})}\right\|}_{\dot{B}_{r,1}^{n/r-1}}$ $\leqslant\sum_{j=1}^{\infty}{\left\|V_{j}^{({\rm r})}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}$ (4.84)
-$\leqslant C\lambda_{1}^{\frac{n}{r}-2}\lambda^{3-2\mu}=C\lambda^{(1-\frac{n}{p})\mu}\lambda_{1}^{\frac{n}{r}-2}\lambda^{3-(3-\frac{n}{p})\mu}.$ (4.85)
+absolutely converges in the strong topology of $\dot{B}_{\rho,1}^{n/\rho-1}(\mathbb{R}^{n})$ and
+${\left\|V^{({\rm r})}\right\|}_{\dot{B}_{\rho,1}^{n/\rho-1}}$ $\leqslant\sum_{j=1}^{\infty}{\left\|V_{j}^{({\rm r})}\right\|}_{\dot{B}_{\rho,1}^{\frac{n}{\rho}-1}}$ (4.81)
+$\leqslant C\sum_{j=1}^{\infty}\lambda^{nj}\lambda_{j-1}^{6-2\mu}\sigma_{j}\left(\frac{n}{\rho}-2,1\right)$ (4.82)
+$\leqslant C\sum_{j=1}^{\infty}\lambda^{nj}\lambda_{j-1}^{6-(4-\frac{n}{\rho})\mu}$ (4.83)
+$\leqslant C\lambda^{n+6-(4-\frac{n}{\rho})\mu}$ (4.84)
+$=C\lambda^{n(1-\frac{1}{p})+2-(1+\frac{n}{p}-\frac{n}{\rho})\mu}\lambda^{\frac{n}{p}+4-(3-\frac{n}{p})\mu}.$ (4.85)
 Now, we define $V:=V^{({\rm p})}+V^{({\rm r})}$. Then, we see that
-$\begin{aligned} {\left\|V\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}&\leqslant{\left\|V^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}+C{\left\|V^{({\rm r})}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\\ &\leqslant C\left(1+\lambda^{(1-\frac{n}{p})\mu}\lambda_{1}^{\frac{n}{r}-2}\right)\lambda^{3-(3-\frac{n}{p})\mu}\\ &\leqslant C\lambda^{3-(3-\frac{n}{p})\mu},\end{aligned}$ (4.86)
-$\begin{aligned} {\left\|V\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}&\geqslant{\left\|V^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}-C{\left\|V^{({\rm r})}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\\ &\geqslant C^{-1}\left(1-C\lambda^{(1-\frac{n}{p})\mu}\lambda_{1}^{\frac{n}{r}-2}\right)\lambda^{3-(3-\frac{n}{p})\mu}\\ &\geqslant C^{-1}\lambda^{3-(3-\frac{n}{p})\mu},\end{aligned}$ (4.87)
-which implies the optimal estimate (4.1).
+${\left\|V\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}$ $\leqslant{\left\|V^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}+C{\left\|V^{({\rm r})}\right\|}_{\dot{B}_{p,1}^{\frac{n}{p}-1}}$ (4.86)
+$\leqslant C\left(1+\lambda^{n(1-\frac{1}{p})+2-\mu}\right)\lambda^{\frac{n}{p}+4-(3-\frac{n}{p})\mu}$ (4.87)
+$\leqslant C\lambda^{\frac{n}{p}+4-(3-\frac{n}{p})\mu}=C\lambda^{-\alpha(p)},$ (4.88)
+and
+${\left\|V\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}$ $\geqslant{\left\|V^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}-C{\left\|V^{({\rm r})}\right\|}_{\dot{B}_{p,1}^{\frac{n}{p}-1}}$ (4.89)
+$\geqslant C^{-1}\left(1-C^{2}\lambda^{n(1-\frac{1}{p})+2-\mu}\right)\lambda^{\frac{n}{p}+4-(3-\frac{n}{p})\mu}$ (4.90)
+$\geqslant C^{-1}\lambda^{\frac{n}{p}+4-(3-\frac{n}{p})\mu}=C^{-1}\lambda^{-\alpha(p)},$ (4.91)
+provided that $\lambda$ is chosen sufficiently large. Thus, we obtain (4.1).
 Step 2. Estimates of the perturbed force. In order to gain a better regularity and estimate for the force
 $F=-\Delta V+\mathbb{P}\operatorname{div}(V\otimes V),$ (4.92)
-we focus on its precise structure. To this end, we may decompose $F$ as
+we focus on its precise structure. To this end, we decompose $F$ as
 $F=F_{1}+F_{2}+F_{3};$ (4.93)
@@
 We first consider the estimate for $F_{1}$, which includes the most crucial cancellation in our analysis. Using the formula $2\cos^{2}\theta=1+\cos 2\theta$ and
-$\lambda_{j-1}^{6}\psi_{j}(x)^{2}\sum_{k\in\mathcal{K}}\Gamma_{k}\left({\rm Id}-\lambda_{j-1}^{-6}\mathcal{D}V_{j-1}\right)^{2}k\otimes k$ $=\lambda_{j-1}^{6}\psi_{j}(x)^{2}\left({\rm Id}-\lambda_{j-1}^{-6}\mathcal{D}V_{j-1}\right)$ (4.93)
-$=\lambda_{j-1}^{6}\psi_{j}^{2}(x){\rm Id}-\mathcal{D}V_{j-1},$ (4.94)
+$\lambda_{j-1}^{8}\psi_{j}^{2}\sum_{k\in\mathcal{K}}\Gamma_{k}\left({\rm Id}-\lambda_{j-1}^{-8}\mathcal{D}V_{j-1}\right)^{2}k\otimes k$ $=\lambda_{j-1}^{8}\psi_{j}^{2}\left({\rm Id}-\lambda_{j-1}^{-8}\mathcal{D}V_{j-1}\right)$ (4.97)
+$=\lambda_{j-1}^{8}\psi_{j}^{2}{\rm Id}-\mathcal{D}V_{j-1},$ (4.98)
 which is implied by Lemma 4.3 and $\psi_{j}(x)=1$ for $x\in\operatorname{supp}V_{j-1}$, we have
-$V^{({\rm p})}_{j}\otimes V^{({\rm p})}_{j}$={} $2\lambda_{j-1}^{6}\sum_{k_{1},k_{2}\in\mathcal{K}}\Gamma_{j,k_{1}}[V_{j-1}]\Gamma_{j,k_{2}}[V_{j-1}]\Psi_{j,k_{1}}\Psi_{j,k_{2}}k_{1}\otimes k_{2}$ (4.95)
-$={}$ $\lambda_{j-1}^{6}\psi_{j}^{2}\sum_{k\in\mathcal{K}}\Gamma_{k}\left({\rm Id}-\lambda_{j-1}^{6}\mathcal{D}V_{j-1}\right)^{2}k\otimes k$ (4.96)
-$+\frac{\lambda_{j-1}^{6}}{h_{j}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\Lambda_{j}}\frac{1}{\ell}\Gamma_{k}\left({\rm Id}-\lambda_{j-1}^{-6}\mathcal{D}V_{j-1}\right)^{2}\psi_{j}^{2}\cos\left(2\lambda^{\ell}a_{k}\cdot x\right)k\otimes k$ (4.97)
-$+2\lambda_{j-1}^{6}\sum_{k_{1}\neq k_{2}}\Gamma_{j,k_{1}}[V_{j-1}]\Gamma_{j,k_{2}}[V_{j-1}]\Psi_{j,k_{1}}\Psi_{j,k_{2}}k_{1}\otimes k_{2}$ (4.98)
-$={}$ $\lambda_{j-1}^{6}\psi_{j}^{2}{\rm Id}-\mathcal{D}V_{j-1}+F_{1,1;j}+F_{1,2;j},$ (4.99)
+$V^{({\rm p})}_{j}\otimes V^{({\rm p})}_{j}={}2\lambda_{j-1}^{8}\sum_{k_{1},k_{2}\in\mathcal{K}}\sum_{\ell_{1},\ell_{2}\in\Lambda_{j}}\Phi_{j,k_{1},\ell_{1}}\otimes\Phi_{j,k_{2},\ell_{2}}$ (4.99)
+$\quad={}\lambda_{j-1}^{8}\psi_{j}^{2}\sum_{k\in\mathcal{K}}\Gamma_{k}\left({\rm Id}-\lambda_{j-1}^{-8}\mathcal{D}V_{j-1}(x)\right)^{2}k\otimes k$ (4.100)
+$\qquad+\frac{\lambda_{j-1}^{8}}{h_{j}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\Lambda_{j}}\frac{1}{\ell}\Gamma_{k}\left({\rm Id}-\lambda_{j-1}^{-8}\mathcal{D}V_{j-1}(x)\right)^{2}\psi_{j}^{2}(x)\cos\left(2\lambda^{\ell}a_{k}\cdot x\right)k\otimes k$ (4.101)
+$\qquad+2\lambda_{j-1}^{8}\sum_{\begin{subarray}{c}k_{1},k_{2}\in\mathcal{K}\\ k_{1}\neq k_{2}\end{subarray}}\sum_{\ell_{1},\ell_{2}\in\Lambda_{j}}\Phi_{j,k_{1},\ell}(x)\Phi_{j,k_{2},\ell}(x)$ (4.102)
+$\quad={}\lambda_{j-1}^{8}\psi_{j}^{2}{\rm Id}-\mathcal{D}V_{j-1}+F_{1,1;j}+F_{1,2;j},$ (4.103)
 where we have set
-$F_{1,1;j}:=\frac{\lambda_{j-1}^{6}}{h_{j}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\Lambda_{j}}\frac{1}{\ell}\Gamma_{j,k}\left[V_{j-1}\right]^{2}\psi_{j}^{2}\cos\left(2\lambda^{\ell}a_{k}\cdot x\right)k\otimes k,$ (4.100)
-$F_{1,2;j}:=2\lambda_{j-1}^{6}\sum_{\begin{subarray}{c}k_{1},k_{2}\in\mathcal{K}\\ k_{1}\neq k_{2}\end{subarray}}\Gamma_{j,k_{1}}[V_{j-1}]\Gamma_{j,k_{2}}[V_{j-1}]\Psi_{j,k_{1}}\Psi_{j,k_{2}}k_{1}\otimes k_{2}.$ (4.101)
+$F_{1,1;j}:=\frac{\lambda_{j-1}^{8}}{h_{j}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\Lambda_{j}}\frac{1}{\ell}\Gamma_{j,k}\left[V_{j-1}\right]^{2}\psi_{j}^{2}\cos\left(2\lambda^{\ell}a_{k}\cdot x\right)k\otimes k,$ (4.104)
+$F_{1,2;j}:=2\lambda_{j-1}^{8}\sum_{\begin{subarray}{c}k_{1},k_{2}\in\mathcal{K}\\ k_{1}\neq k_{2}\end{subarray}}\sum_{\ell_{1},\ell_{2}\in\Lambda_{j}}\Phi_{j,k_{1},\ell_{1}}(x)\otimes\Phi_{j,k_{2},\ell_{2}}(x).$ (4.105)
 Then, from the cancellation property
-$-\Delta V_{j-1}+\mathbb{P}\operatorname{div}\left(\lambda_{j-1}^{6}\psi_{j}^{2}{\rm Id}-\mathcal{D}V_{j-1}\right)=-\Delta V_{j-1}+\mathbb{P}\nabla(\lambda_{j-1}^{6}\psi_{j}^{2})+\Delta V_{j-1}=0$ (4.102)
+$-\Delta V_{j-1}+\mathbb{P}\operatorname{div}\left(\lambda_{j-1}^{8}\psi_{j}^{2}{\rm Id}-\mathcal{D}V_{j-1}\right)=-\Delta V_{j-1}+\mathbb{P}\nabla(\lambda_{j-1}^{8}\psi_{j}^{2})+\Delta V_{j-1}=0$ (4.106)
 it follows that
-$F_{1}$ $=\sum_{j=1}^{\infty}\left(-\Delta V_{j-1}+\mathbb{P}\operatorname{div}\left(\lambda_{j-1}^{6}\psi_{j}^{2}{\rm Id}-\mathcal{D}V_{j-1}+F_{1,1;j}+F_{1,2;j}\right)\right)$ (4.103)
+$F_{1}$ $=\sum_{j=1}^{\infty}\left(-\Delta V_{j-1}+\mathbb{P}\operatorname{div}\left(\lambda_{j-1}^{8}\psi_{j}^{2}{\rm Id}-\mathcal{D}V_{j-1}+F_{1,1;j}+F_{1,2;j}\right)\right)$ (4.107)
 $=\sum_{j=1}^{\infty}\mathbb{P}\operatorname{div}F_{1,1;j}+\sum_{j=1}^{\infty}\mathbb{P}\operatorname{div}F_{1,2;j}$ (4.108)
@@
 To proceed with the estimate of $F_{1,1}$, we begin with the estimate of each element $F_{1,1;j}$. It follows from Proposition 3.1 with $M=n$ that
-${\left\|F_{1,1;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}$ $\leqslant C\lambda_{j-1}^{6-4\mu}\sum_{\ell\in\lambda_{j}}\frac{\lambda^{(\frac{n}{r}-2)\ell}}{\ell}\sum_{k\in\mathcal{K}}{\left\|\Gamma_{j,k}[V_{j-1}]\psi_{j}\right\|}_{L^{r}}$ (4.106)
-$\quad+C\lambda_{j-1}^{6-4\mu}\sum_{\ell\in\lambda_{j}}\frac{\lambda^{-2\ell}}{\ell}\sum_{k\in\mathcal{K}}{\left\|\nabla^{n}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{r}\cap L^{1}}$ (4.107)
-$\leqslant C\lambda_{j-1}^{6-4\mu}\cdot C\lambda_{j-1}^{(\frac{n}{r}-2)\mu}\cdot C\lambda^{\frac{n}{r}}+C\lambda_{j-1}^{6-4\mu}\cdot C\lambda_{j-1}^{-2\mu}\cdot C\lambda^{n}\lambda_{j-1}^{2n}$ (4.108)
-$\leqslant C\lambda_{j-1}^{2n+6-4\mu}.$ (4.109)
-For the estimate of $F_{1,2;j}$, since it is rewritten as
-$F_{1,2;j}(x)=\frac{\lambda_{j-1}^{6}}{h_{j}}\sum_{\pm}\sum_{\ell_{1},\ell_{2}\in\Lambda_{j}}\sum_{\begin{subarray}{c}k_{1}\neq k_{2}\end{subarray}}$ $\frac{1}{\sqrt{\ell_{1}\ell_{2}}}\Gamma_{j,k_{1}}[V_{j-1}](x)\Gamma_{j,k_{2}}[V_{j-1}](x)$ (4.110)
+${\left\|F_{1,1;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}$ $\leqslant C\lambda_{j-1}^{8-4\mu}\sum_{\ell\in\lambda_{j}}\frac{\lambda^{(\frac{n}{r}-2)\ell}}{\ell}\sum_{k\in\mathcal{K}}{\left\|\Gamma_{j,k}[V_{j-1}]\psi_{j}\right\|}_{L^{r}}$ (4.110)
+$\quad+C\lambda_{j-1}^{8-4\mu}\sum_{\ell\in\lambda_{j}}\frac{\lambda^{-2\ell}}{\ell}\sum_{k\in\mathcal{K}}{\left\|\nabla^{n}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{r}\cap L^{1}}$ (4.111)
+$\leqslant C\lambda_{j-1}^{8-4\mu}\cdot C\lambda_{j-1}^{(\frac{n}{r}-2)\mu}\cdot C\lambda^{\frac{n}{r}j}+C\lambda_{j-1}^{8-4\mu}\cdot C\lambda_{j-1}^{-2\mu}\cdot C\lambda^{nj}\lambda_{j-1}^{2n}$ (4.112)
+$\leqslant C\lambda_{j-1}^{4n+8-4\mu}.$ (4.113)
+For the estimate of $F_{1,2;j}$, since this part is rewritten as
+$F_{1,2;j}(x)=\frac{\lambda_{j-1}^{8}}{h_{j}}\sum_{\pm}\sum_{\ell_{1},\ell_{2}\in\Lambda_{j}}\sum_{\begin{subarray}{c}k_{1}\neq k_{2}\end{subarray}}$ $\frac{1}{\sqrt{\ell_{1}\ell_{2}}}\Gamma_{j,k_{1}}[V_{j-1}](x)\Gamma_{j,k_{2}}[V_{j-1}](x)$ (4.114)
 $\psi_{j}(x)^{2}\cos\left((\lambda^{\ell_{1}}a_{k_{1}}\pm\lambda^{\ell_{2}}a_{k_{2}})\cdot x\right)k_{1}\otimes k_{2},$ (4.115)
 we see that
-${\left\|F_{1,2;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}\leqslant{}$ $C\lambda_{j-1}^{6-2\mu}\sum_{\pm}\sum_{\begin{subarray}{c}\ell_{1}\in\Lambda_{j}\end{subarray}}\sum_{\ell_{2}:\ell_{2}\geqslant\ell_{1}}\sum_{\begin{subarray}{c}k_{1}\neq k_{2}\end{subarray}}$ (4.112)
+${\left\|F_{1,2;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}\leqslant{}$ $C\lambda_{j-1}^{8-4\mu}\sum_{\pm}\sum_{\begin{subarray}{c}\ell_{1}\in\Lambda_{j}\end{subarray}}\sum_{\ell_{2}:\ell_{2}\geqslant\ell_{1}}\sum_{\begin{subarray}{c}k_{1}\neq k_{2}\end{subarray}}$ (4.116)
 $\left(\frac{\lambda^{(\frac{n}{r}-2)\ell_{2}}}{\sqrt{\ell_{1}\ell_{2}}}|\lambda^{\ell_{1}-\ell_{2}}a_{k_{1}}\pm a_{k_{2}}|^{\frac{n}{r}-2}{\left\|\Gamma_{j,k_{1}}[V_{j-1}]\Gamma_{j,k_{2}}[V_{j-1}]\psi_{j}^{2}\right\|}_{L^{r}}\right$.
 (4.117)
-$\left.\quad+\frac{\lambda^{-2\ell_{2}}}{\sqrt{\ell_{1}\ell_{2}}}|\lambda^{\ell_{1}-\ell_{2}}a_{k_{1}}\pm a_{k_{2}}|^{-2}{\left\|\nabla^{n}(\Gamma_{j,k_{1}}[V_{j-1}]\Gamma_{j,k_{2}}[V_{j-1}]\psi_{j}^{2})\right\|}_{L^{r}}\right)$ (4.114)
-$\leqslant{}$ $C\lambda_{j-1}^{6-4\mu}\cdot C\lambda_{j-1}^{(\frac{n}{r}-2)\mu}\cdot C\lambda^{\frac{n}{r}}+C\lambda_{j-1}^{6-4\mu}\cdot C\lambda_{j-1}^{-2\mu}\cdot C\lambda^{n}\lambda_{j-1}^{2n}$ (4.115)
-$\leqslant{}$ $C\lambda_{j-1}^{2n+6-4\mu},$ (4.116)
+$\left.\quad+\frac{\lambda^{-2\ell_{2}}}{\sqrt{\ell_{1}\ell_{2}}}|\lambda^{\ell_{1}-\ell_{2}}a_{k_{1}}\pm a_{k_{2}}|^{-2}{\left\|\nabla^{n}(\Gamma_{j,k_{1}}[V_{j-1}]\Gamma_{j,k_{2}}[V_{j-1}]\psi_{j}^{2})\right\|}_{L^{r}\cap L^{1}}\right)$ (4.118)
+$\leqslant{}$ $C\lambda_{j-1}^{8-4\mu}\cdot C\lambda_{j-1}^{(\frac{n}{r}-2)\mu}\cdot C\lambda^{\frac{n}{r}}+C\lambda_{j-1}^{8-4\mu}\cdot C\lambda_{j-1}^{-2\mu}\cdot C\lambda^{n}\lambda_{j-1}^{2n}$ (4.119)
+$\leqslant{}$ $C\lambda_{j-1}^{4n+8-4\mu},$ (4.120)
 where we have used
-$|\lambda^{\ell_{1}-\ell_{2}}a_{k_{1}}\pm a_{k_{2}}|^{2}={}$ $\left(\lambda^{\ell_{1}-\ell_{2}}|a_{k_{1}}|\pm|a_{k_{2}}|\cos\vartheta(a_{k_{1}},a_{k_{2}})\right)^{2}+|a_{k_{2}}|^{2}\sin^{2}\vartheta(a_{k_{1}},a_{k_{2}})$ (4.117)
-$\geqslant{}$ $|a_{k_{2}}|^{2}\sin^{2}\vartheta(a_{k_{1}},a_{k_{2}})>0.$ (4.118)
+$|\lambda^{\ell_{1}-\ell_{2}}a_{k_{1}}\pm a_{k_{2}}|^{2}={}$ $\left(\lambda^{\ell_{1}-\ell_{2}}\pm\cos\vartheta(a_{k_{1}},a_{k_{2}})\right)^{2}+\sin^{2}\vartheta(a_{k_{1}},a_{k_{2}})$ (4.121)
+$\geqslant{}$ $\sin^{2}\vartheta(a_{k_{1}},a_{k_{2}})>0.$ (4.122)
 Here, one may recall (3.13) for the definition of $\vartheta(\cdot,\cdot)$. Hence, we obtain
 ${\left\|F_{1}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-3}}$ $\leqslant C\sum_{j=1}^{\infty}\left({\left\|F_{1,1;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}+{\left\|F_{1,2;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}\right)$ (4.123)
-$\leqslant C\sum_{j=1}^{\infty}\lambda_{j-1}^{2n+6-4\mu}$ (4.120)
-$\leqslant{}C\lambda^{2n+6-4\mu}=C\lambda^{-\beta}.$ (4.121)
+$\leqslant C\sum_{j=1}^{\infty}\lambda_{j-1}^{4n+8-4\mu}$ (4.124)
+$\leqslant{}C\lambda^{4n+8-4\mu}=C\lambda^{-\beta}.$ (4.125)
 Next, we consider the estimate on $F_{2}$. Let us rewrite $F_{2}$ as
@@
 where we have set
-$F_{2;j_{1},j_{2}}:={}V^{({\rm p})}_{j_{1}}\otimes V^{({\rm p})}_{j_{2}}$ (4.123)
-$\quad=2\lambda_{j_{1}-1}^{3}\lambda_{j_{2}-1}^{3}\sum_{k_{1},k_{2}\in\mathcal{K}}\Gamma_{j_{1},k_{1}}[V_{j_{1}-1}]\Gamma_{j_{2},k_{2}}[V_{j_{2}-1}]\Psi_{j_{1},k_{1}}\Psi_{j_{2},k_{2}}k_{1}\otimes k_{2}$ (4.124)
-$\quad=\frac{\lambda_{j_{1}-1}^{3}\lambda_{j_{2}-1}^{3}}{\sqrt{h_{j_{1}}h_{j_{2}}}}\sum_{\pm}\sum_{\ell_{1}\in\Lambda_{j_{1}}}\sum_{\ell_{2}\in\Lambda_{j_{2}}}\sum_{k_{1},k_{2}\in\mathcal{K}}\frac{1}{\sqrt{\ell_{1}\ell_{2}}}\Gamma_{j_{1},k_{1}}[V_{j_{1}-1}](x)\Gamma_{j_{2},k_{2}}[V_{j_{2}-1}](x)$ (4.125)
+$F_{2;j_{1},j_{2}}(x):={}V^{({\rm p})}_{j_{1}}\otimes V^{({\rm p})}_{j_{2}}(x)$ (4.127)
+$\quad=2\lambda_{j_{1}-1}^{4}\lambda_{j_{2}-1}^{4}\sum_{k_{1},k_{2}\in\mathcal{K}}\sum_{\ell_{1},\ell_{2}\in\Lambda_{j}}\Phi_{j_{1},k_{1},\ell_{1}}(x)\otimes\Phi_{j_{2},k_{2},\ell_{2}}(x)$ (4.128)
+$\quad=\frac{\lambda_{j_{1}-1}^{4}\lambda_{j_{2}-1}^{4}}{\sqrt{h_{j_{1}}h_{j_{2}}}}\sum_{\pm}\sum_{\ell_{1}\in\Lambda_{j_{1}}}\sum_{\ell_{2}\in\Lambda_{j_{2}}}\sum_{k_{1},k_{2}\in\mathcal{K}}\frac{1}{\sqrt{\ell_{1}\ell_{2}}}\Gamma_{j_{1},k_{1}}[V_{j_{1}-1}](x)\Gamma_{j_{2},k_{2}}[V_{j_{2}-1}](x)$ (4.129)
 $\qquad\qquad\qquad\qquad\qquad\qquad\psi_{j_{1}}(x)\psi_{j_{2}}(x)\cos\left((\lambda^{\ell_{1}}a_{k_{1}}\pm\lambda^{\ell_{2}}a_{k_{2}})\cdot x\right)k_{1}\otimes k_{2}.$ (4.130)
-As $F_{2;j_{1},j_{2}}=F_{2;j_{2},j_{1}}$, it suffices to consider the case $j_{1}<j_{2}$. Since it holds
-$|\lambda^{\ell_{1}-\ell_{2}}a_{k_{1}}\pm a_{k_{2}}|\geqslant|a_{k_{2}}|-\lambda^{\ell_{1}-\ell_{2}}|a_{k_{1}}|\geqslant\frac{1}{2}|a_{k_{2}}|>0,$ (4.127)
+As $F_{2;j_{1},j_{2}}=(F_{2;j_{2},j_{1}})^{\top}$, it suffices to consider the case $j_{1}<j_{2}$. Since it holds
+$|\lambda^{\ell_{1}-\ell_{2}}a_{k_{1}}\pm a_{k_{2}}|\geqslant 1-\lambda^{\ell_{1}-\ell_{2}}\geqslant\frac{1}{2}>0,$ (4.131)
 we follow the similar strategy as for the estimate of $F_{1,2;j}$ to see that
-${\left\|F_{2;j_{1},j_{2}}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}\leqslant{}$ $C\lambda_{j_{1}-1}^{3-2\mu}\lambda_{j_{2}-1}^{3-2\mu}\sum_{\ell_{1}\in\Lambda_{j_{1}}}\sum_{\ell_{2}\in\Lambda_{j_{2}}}\sum_{k_{1},k_{2}\in\mathcal{K}}$ (4.128)
+${\left\|F_{2;j_{1},j_{2}}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}\leqslant{}$ $C\lambda_{j_{1}-1}^{4-2\mu}\lambda_{j_{2}-1}^{4-2\mu}\sum_{\ell_{1}\in\Lambda_{j_{1}}}\sum_{\ell_{2}\in\Lambda_{j_{2}}}\sum_{k_{1},k_{2}\in\mathcal{K}}$ (4.132)
 $\left(\frac{\lambda^{(\frac{n}{r}-2)\ell_{2}}}{\sqrt{\ell_{1}\ell_{2}}}|\lambda^{\ell_{1}-\ell_{2}}a_{k_{1}}\pm a_{k_{2}}|^{\frac{n}{r}-2}{\left\|\Gamma_{j_{1},k_{1}}[V_{j_{1}-1}]\Gamma_{j_{2},k_{2}}[V_{j_{2}-1}]\psi_{j_{1}}\right\|}_{L^{r}}\right$.
 (4.133)
-$\left.\quad+\frac{\lambda^{-2\ell_{2}}}{\sqrt{\ell_{1}\ell_{2}}}|\lambda^{\ell_{1}-\ell_{2}}a_{k_{1}}\pm a_{k_{2}}|^{-2}{\left\|\nabla^{n}(\Gamma_{j_{1},k_{1}}[V_{j_{1}-1}]\Gamma_{j_{2},k_{2}}[V_{j_{2}-1}]\psi_{j_{1}})\right\|}_{L^{r}}\right)$ (4.130)
-$\leqslant{}$ $C\lambda^{\frac{n}{r}j_{1}}\lambda_{j_{1}-1}^{3-2\mu}\lambda_{j_{2}-1}^{3-2\mu}\sum_{\ell_{1}\in\Lambda_{j_{1}}}\sum_{\ell_{2}:\ell_{2}\geqslant\ell_{1}}\lambda^{(\frac{n}{r}-2)\ell_{2}}$ (4.131)
-$+C\lambda^{nj_{1}}\lambda_{j_{1}-1}^{3-2\mu}\lambda_{j_{2}-1}^{3-2\mu+2n}\sum_{\ell_{1}\in\Lambda_{j_{1}}}\sum_{\ell_{2}:\ell_{2}\geqslant\ell_{1}}\lambda^{-2\ell_{2}}$ (4.132)
-$\leqslant{}$ $C\lambda^{\frac{n}{r}j_{1}}\lambda_{j_{1}-1}^{3-(4-\frac{n}{r})\mu}\lambda_{j_{2}-1}^{3-2\mu}+C\lambda_{j_{1}-1}^{3-4\mu}\lambda_{j_{2}-1}^{3-2\mu+2n}$ (4.133)
-$\leqslant{}$ $C\lambda^{\frac{n}{r}j_{1}}\lambda_{j_{1}-1}^{3-(4-\frac{n}{r})\mu}\lambda_{j_{2}-1}^{3-2\mu}.$ (4.134)
+$\left.\quad+\frac{\lambda^{-2\ell_{2}}}{\sqrt{\ell_{1}\ell_{2}}}|\lambda^{\ell_{1}-\ell_{2}}a_{k_{1}}\pm a_{k_{2}}|^{-2}{\left\|\nabla^{n}(\Gamma_{j_{1},k_{1}}[V_{j_{1}-1}]\Gamma_{j_{2},k_{2}}[V_{j_{2}-1}]\psi_{j_{1}})\right\|}_{L^{r}\cap L^{1}}\right)$ (4.134)
+$\leqslant{}$ $C\lambda^{\frac{n}{r}j_{1}}\lambda_{j_{1}-1}^{4-2\mu}\lambda_{j_{2}-1}^{4-2\mu}\sum_{\ell_{1}\in\Lambda_{j_{1}}}\sum_{\ell_{2}:\ell_{2}\geqslant\ell_{1}}\lambda^{(\frac{n}{r}-2)\ell_{2}}$ (4.135)
+$+C\lambda^{nj_{1}}\lambda_{j_{1}-1}^{4-2\mu}\lambda_{j_{2}-1}^{4-2\mu+2n}\sum_{\ell_{1}\in\Lambda_{j_{1}}}\sum_{\ell_{2}:\ell_{2}\geqslant\ell_{1}}\lambda^{-2\ell_{2}}$ (4.136)
+$\leqslant{}$ $C\lambda^{\frac{n}{r}j_{1}}\lambda_{j_{1}-1}^{4-(4-\frac{n}{r})\mu}\lambda_{j_{2}-1}^{4-2\mu}+C\lambda^{nj_{1}}\lambda_{j_{1}-1}^{4-4\mu}\lambda_{j_{2}-1}^{4-2\mu+2n}$ (4.137)
+$\leqslant{}$ $C\lambda^{nj_{1}}\lambda_{j_{1}-1}^{4-(4-\frac{n}{r})\mu}\lambda_{j_{2}-1}^{4-2\mu}.$ (4.138)
 Hence, we obtain
 ${\left\|F_{2}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-3}}$ $\leqslant C\sum_{j_{1}=1}^{\infty}\sum_{j_{2}=j_{1}+1}^{\infty}{\left\|F_{2;j_{1},j_{2}}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}$ (4.139)
-$\leqslant C\sum_{j_{1}=1}^{\infty}\lambda^{\frac{n}{r}j_{1}}\lambda_{j_{1}-1}^{3-(4-\frac{n}{r})\mu}\sum_{j_{2}=j_{1}+1}^{\infty}\lambda_{j_{2}-1}^{3-2\mu}$ (4.136)
-$\leqslant C\sum_{j_{1}=1}^{\infty}\lambda_{j_{1}-1}^{6-(6-\frac{n}{r})\mu}$ (4.137)
-$\leqslant C\lambda^{6-(6-\frac{n}{r})\mu}\leqslant C\lambda^{-\beta}.$ (4.138)
+$\leqslant C\sum_{j_{1}=1}^{\infty}\lambda^{nj_{1}}\lambda_{j_{1}-1}^{4-(4-\frac{n}{r})\mu}\sum_{j_{2}=j_{1}+1}^{\infty}\lambda_{j_{2}-1}^{4-2\mu}$ (4.140)
+$\leqslant C\sum_{j_{1}=1}^{\infty}\lambda^{nj_{1}}\lambda_{j_{1}-1}^{8-(6-\frac{n}{r})\mu}$ (4.141)
+$\leqslant C\lambda^{8+n-(6-\frac{n}{r})\mu}\leqslant C\lambda^{4n+8-4\mu}=C\lambda^{-\beta}.$ (4.142)
 Finally, we focus on the estimate of $F_{3}$. Let $n<p_{*}<2n$ satisfy (2.7) with $p=p_{*}$. It then follows from Lemma 2.1 that
 ${\left\|F_{3}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-3}}$ $\leqslant C\left({\left\|V^{({\rm p})}\right\|}_{\dot{B}_{p_{*},1}^{\frac{n}{p_{*}}-1}}+{\left\|V^{({\rm r})}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\right){\left\|V^{({\rm r})}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}$ (4.143)
-$\leqslant C\left(\lambda^{3-(3-\frac{n}{p_{*}})\mu}+\lambda_{1}^{\frac{n}{r}-2}\lambda^{3-2\mu}\right)\lambda_{1}^{\frac{n}{r}-2}\lambda^{3-2\mu}$ (4.140)
-$\leqslant C\lambda_{1}^{\frac{n}{r}-2}\lambda^{6-4\mu}\leqslant C\lambda^{-\beta}.$ (4.141)
+$\leqslant C\left(\lambda^{\frac{n}{p_{*}}+4-(3-\frac{n}{p_{*}})\mu}+\lambda^{n+6-(4-\frac{n}{r})\mu}\right)\lambda^{n+6-(4-\frac{n}{r})\mu}$ (4.144)
+$\leqslant C\left(\lambda^{\frac{9}{2}-\frac{5}{2}\mu}+\lambda^{n+6-2\mu}\right)\lambda^{n+6-2\mu}$ (4.145)
+$\leqslant C\lambda^{2n+12-4\mu}<C\lambda^{4n+8-4\mu}=C\lambda^{-\beta}.$ (4.146)
 Combining (4.125), (4.142), and (4.146), we obtain (4.3). Thus, we complete the proof.
@@
 We separate the proof into two steps; the first step achieves the construction of the non-trivial solutions to the unforced stationary Navier–Stokes equations, and the second step establishes the desired non-stationary flow by the standard stability argument around the steady state provided in the previous step.
+Let $V_{\lambda}$ and $F_{\lambda}$ be vector fields constructed in Theorem 4.1.
 Step 1.
 Steady Navier–Stokes flow around $V_{\lambda}$.
-Fix a $n<p_{*}<2n$ and $n/2<r<n$ so that (2.7) with $p=p_{*}$.
+Fix a $n<p_{*}<2n$ and $n/2<r<n$ so that (2.7) holds with $p$ replaced by $p_{*}$.
 Let us construct the solution to (1.2) of the form $U_{\lambda}=V_{\lambda}+W_{\lambda}$.
@@
 $\begin{cases}-\Delta W_{\lambda}+\mathbb{P}\operatorname{div}\left(V_{\lambda}\otimes W_{\lambda}+W_{\lambda}\otimes V_{\lambda}+W_{\lambda}\otimes W_{\lambda}\right)=-F,\quad&x\in\mathbb{R}^{n},\\ \operatorname{div}W_{\lambda}=0,\quad&x\in\mathbb{R}^{n},\end{cases}$ (5.1)
-and define an operator $\mathcal{X}$ on $\dot{B}_{r,1}^{n/r-1}$ by
-$\mathcal{X}[W]:=-(-\Delta)^{-1}F-(-\Delta)^{-1}\mathbb{P}\operatorname{div}\left(V_{\lambda}\otimes W+W\otimes V_{\lambda}+W\otimes W\right).$ (5.2)
+and define an operator $\mathcal{X}$ on $\dot{B}_{r,1}^{n/r-1}(\mathbb{R}^{n})$ by
+$\mathcal{X}[W]:=-(-\Delta)^{-1}F_{\lambda}-(-\Delta)^{-1}\mathbb{P}\operatorname{div}\left(V_{\lambda}\otimes W+W\otimes V_{\lambda}+W\otimes W\right).$ (5.2)
 Then, it follows from Lemma 2.1 that
-$\begin{aligned} {\left\|\mathcal{X}[W]\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}&\leqslant C{\left\|F\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-3}}+C\left({\left\|V_{\lambda}\right\|}_{\dot{B}_{p_{*},1}^{\frac{n}{p_{*}}-1}}+{\left\|W\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\right){\left\|W\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}},\\ &\leqslant C_{1}\lambda^{-\beta}+C_{1}\left(\lambda^{-\alpha(p_{*})}+{\left\|W\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\right){\left\|W\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}},\end{aligned}$ (5.3)
+$\begin{aligned} {\left\|\mathcal{X}[W]\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}&\leqslant C{\left\|F_{\lambda}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-3}}+C\left({\left\|V_{\lambda}\right\|}_{\dot{B}_{p_{*},1}^{\frac{n}{p_{*}}-1}}+{\left\|W\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\right){\left\|W\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}},\\ &\leqslant C_{1}\lambda^{-\beta}+C_{1}\left(\lambda^{-\alpha(p_{*})}+{\left\|W\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\right){\left\|W\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}},\end{aligned}$ (5.3)
 ${\left\|\mathcal{X}[W_{1}]-\mathcal{X}[W_{2}]\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}$ (5.4)
@@
 Then, choosing $\lambda$ so large that $2C_{1}\left(\lambda^{-\alpha(p_{*})}+4C_{1}\lambda^{-\beta}\right)\leqslant 1$, we have
-${\left\|\mathcal{X}[W]\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\leqslant C_{1}\lambda^{-\beta},\qquad{\left\|\mathcal{X}[W_{1}]-\mathcal{X}[W_{2}]\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\leqslant\frac{1}{2}{\left\|W_{1}-W_{2}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}$ (5.8)
+${\left\|\mathcal{X}[W]\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\leqslant 2C_{1}\lambda^{-\beta},\qquad{\left\|\mathcal{X}[W_{1}]-\mathcal{X}[W_{2}]\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\leqslant\frac{1}{2}{\left\|W_{1}-W_{2}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}$ (5.8)
 for all $W,W_{1},W_{2}\in\mathscr{X}$. Hence, the contraction mapping principle yields the existence of a unique solution $W_{\lambda}\in\mathscr{X}$ to (5.1).
@@
 $\begin{cases}\partial_{t}w_{\lambda}-\Delta w_{\lambda}+\mathbb{P}\operatorname{div}(U_{\lambda}\otimes w_{\lambda}+U_{\lambda}\otimes w_{\lambda}+w_{\lambda}\otimes w_{\lambda})=0,\\ \operatorname{div}w_{\lambda}=0,\\ w_{\lambda}(0,x)=w_{0}:=u_{0}-U_{\lambda}.\end{cases}$ (5.11)
-To construct the perturbed solution, we define a map
+To construct the perturbed solution, we define a mapping
 $\mathcal{Y}[w](t):=e^{t\Delta}w_{0}-\int_{0}^{t}e^{(t-\tau)\Delta}\mathbb{P}\operatorname{div}\left(U_{\lambda}\otimes w(\tau)+w(\tau)\otimes U_{\lambda}+w(\tau)\otimes w(\tau)\right)d\tau.$ (5.12)
@@
 ${\left\|\mathcal{Y}[w]\right\|}_{\widetilde{L^{\infty}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1})\cap\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}$ (5.13)
-$\quad\leqslant{}C{\left\|w_{0}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}+C\left({\left\|w_{0}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}+{\left\|w\right\|}_{\widetilde{L^{\infty}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1})}\right){\left\|w\right\|}_{\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}$ (5.14)
-$\quad\leqslant{}C_{2}{\left\|U_{\lambda}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}+C_{2}\left(\lambda^{-\alpha(p)}+{\left\|w\right\|}_{\widetilde{L^{\infty}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1})}\right){\left\|w\right\|}_{\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}$ (5.15)
+$\quad\leqslant{}C{\left\|w_{0}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}+C\left({\left\|U_{\lambda}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}+{\left\|w\right\|}_{\widetilde{L^{\infty}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1})}\right){\left\|w\right\|}_{\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}$ (5.14)
+$\quad\leqslant{}C_{2}{\left\|w_{0}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}+C_{2}\left(\lambda^{-\alpha(p)}+{\left\|w\right\|}_{\widetilde{L^{\infty}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1})}\right){\left\|w\right\|}_{\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}$ (5.15)
 and
@@
 Finally, we shall prove the asymptotic behavior $u_{\lambda}(t)-U_{\lambda}=w_{\lambda}(t)\to 0$ strongly in $\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ as $t\to\infty$.
-To see this, we fix a $T>0$ and rewrite the integral equation for $w_{\lambda}$ as
+Let us fix a $T>0$ and rewrite the integral equation for $w_{\lambda}$ as
 $w_{\lambda}(t)={}$ $e^{(t-T)\Delta}w_{\lambda}(T)$ (5.20)

### v3

@@
-arXiv:2602.19846v2[math.AP] 02 Mar 2026
+Sharp non-uniqueness for the Navier–Stokes equations in scaling critical spaces
 Abstract.
@@
 We achieve this through a complete classification for every pair $(p,q)$ of whether uniqueness of mild solutions in the critical Besov class $C([0,T);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))$ holds or not.
-Our non-uniqueness mechanism produces infinitely many global solutions emanating even from zero initial state, whose large-time asymptotics are governed by non-trivial stationary flow.
+Our non-uniqueness mechanism produces infinitely many global solutions emanating even from zero initial state, whose large-time asymptotics are governed by non-trivial stationary flows.
 To the best of our knowledge, such non-unique solutions provide the first examples of non-dissipative unforced Navier–Stokes flow with critical regularity.
 Key words and phrases:
-Navier–Sotkes equations, non-uniqueness, scaling critical Besov spaces
+Navier–Stokes equations, non-uniqueness, scaling critical Besov spaces
 2020 Mathematics Subject Classification:
@@
 In dimensions $n\geqslant 3$, Lions–Masmoudi [Lio-Mas] and Furioli–Rieusset–Terraneo [Fur-Rie-Ter-2000] proved that mild solutions to (1.1) are unique in the critical space $C([0,T);L^{n}(\mathbb{R}^{n}))$.
-The aim of this paper is to prove the sharpness of the choice of $L^{n}(\mathbb{R}^{n})$ in the sense that the uniqueness fails as one replaces $L^{n}(\mathbb{R}^{n})$ by a scaling critical space that is even slightly larger; for instance, enlarging $L^{n}(\mathbb{R}^{n})=\dot{F}_{n,2}^{0}(\mathbb{R}^{n})$ to a slightly larger Lizorkin–Triebel space $\dot{F}_{n,q}^{0}(\mathbb{R}^{n})$ with any $q>2$ leads to non-uniqueness of the mild solutions in $C([0,T);\dot{F}_{n,q}^{0}(\mathbb{R}^{n}))$.
+The aim of this paper is to prove the sharpness of the choice of $L^{n}(\mathbb{R}^{n})$ in the sense that the uniqueness fails if one replaces $L^{n}(\mathbb{R}^{n})$ by a scaling critical space that is even slightly larger; for instance, enlarging $L^{n}(\mathbb{R}^{n})=\dot{F}_{n,2}^{0}(\mathbb{R}^{n})$ to a slightly larger Lizorkin–Triebel space $\dot{F}_{n,q}^{0}(\mathbb{R}^{n})$ with any $q>2$ leads to non-uniqueness of the mild solutions in $C([0,T);\dot{F}_{n,q}^{0}(\mathbb{R}^{n}))$.
 To achieve this, we show that the necessary and sufficient condition for the uniqueness of mild solutions to (1.1) in the scaling critical Besov spaces is given by $1\leqslant p<n$ with $1\leqslant q\leqslant\infty$, or $p=n$ with $1\leqslant q\leqslant 2$; see Figure 1 below.
@@
 Moreover, Kozono–Yamazaki [Koz-Yam-94], Cannone–Planchon [Can-Pla-96], and Planchon [Pla-98] proved well-posedness in $\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ for $1\leqslant p<\infty$ and $1\leqslant q\leqslant\infty$.
-In the endpoint case $p=\infty$, Koch–Tataru [Koh-Tat-01] established well-posedness for small data in ${\rm BMO}^{-1}(\mathbb{R}^{n})$, however it was shown by [Bou-Pav-08, Yon-10, Wan-15] that (1.1) is ill-posed in $\dot{B}_{\infty,q}^{-1}(\mathbb{R}^{n})$ ($1\leqslant q\leqslant\infty$).
+In the endpoint case $p=\infty$, Koch–Tataru [Koh-Tat-01] established well-posedness for small data in ${\rm BMO}^{-1}(\mathbb{R}^{n})$, however, it was shown by [Bou-Pav-08, Yon-10, Wan-15] that (1.1) is ill-posed in $\dot{B}_{\infty,q}^{-1}(\mathbb{R}^{n})$ ($1\leqslant q\leqslant\infty$).
 See [IN] for more precise analysis on the well-posedness of (1.1) in critical Besov or Lizorkin–Triebel spaces.
@@
 Finally, we briefly review works on non-uniqueness.
-One major direction originates from applying the Nash convex integration to the Euler equations on the torus (see, [DeLellis-Szekelyhidi-09, Isett-18, Daneri-Szekelyhidi-2017-ARMA, Lellis-Szekelyhidi-2013-Invent, Buckmaster-Lellis-Isett-Szekelyhidi-2015-AnnMath] for instance).
+One major direction originates from applying the Nash convex integration to the Euler equations on the torus (see for instance [DeLellis-Szekelyhidi-09, Isett-18, Daneri-Szekelyhidi-2017-ARMA, Lellis-Szekelyhidi-2013-Invent, Buckmaster-Lellis-Isett-Szekelyhidi-2015-AnnMath]).
 Using an iterative scheme based on the Euler–Reynolds system via the geometric lemma, these works constructed Euler flows violating the energy conservation law and made substantial progress toward the Onsager conjecture.
 Although incorporating the viscous term was long thought to be fundamentally difficult, Buckmaster–Vicol [Buc-Vic-19] introduced the concept of intermittent convex integration and proved that uniqueness of weak solutions to the Navier–Stokes equations on $\mathbb{T}^{3}$ breaks down in the supercritical class $C([0,T);H^{\beta}(\mathbb{T}^{3}))$ for some $0<\beta\ll 1/2$.
-Cheskidov–Luo [Che-Luo-21] showed that an endpoint Ladyzhenskaya–Prodi–Serrin class $L^{2}(0,T;L^{\infty}(\mathbb{T}^{n}))$, is optimal with respect to the time integrability exponent; they constructed non-unique solutions in $L^{p}(0,T;L^{\infty}(\mathbb{T}^{3}))$ for $p<2$.
+Cheskidov–Luo [Che-Luo-21] showed that an endpoint Ladyzhenskaya–Prodi–Serrin class $L^{2}(0,T;L^{\infty}(\mathbb{T}^{n}))$ is optimal with respect to the time integrability exponent; they constructed non-unique solutions in $L^{p}(0,T;L^{\infty}(\mathbb{T}^{3}))$ for $p<2$.
 In [Che-Luo-22-AnnPDE], they also considered another endpoint in two dimensions and proved non-uniqueness in $C([0,T);L^{p}(\mathbb{T}^{2}))$ for $p<2$.
@@
 See Aoki–Maekawa [Aok-Mae] for the corresponding result on a two-dimensional half space.
-More recently, Hou–Wang–Yang [2509.25116v1] announced an analogous statement in three-dimension with the unforced case $f\equiv 0$ by a computer-assisted proof.
+More recently, Hou–Wang–Yang [2509.25116v1] announced an analogous statement in three dimensions with the unforced case $f\equiv 0$ by a computer-assisted proof.
 1.2. Main results
@@
     $n<p\leqslant\infty$ and $1\leqslant q\leqslant\infty$.
-Then, there exist a monotone decreasing sequence $\{\varepsilon_{m}\}_{m\in\mathbb{N}}\subset(0,1)$ with $\varepsilon_{m}\to 0$ and a constant $\eta=\eta(n,p,q)\geqslant 0$ such that for any $m\in\mathbb{N}$ and initial datum $u_{0}\in\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ with
+Then, there exist a monotone decreasing sequence $\{\varepsilon_{m}\}_{m\in\mathbb{N}}\subset(0,\infty)$ with $\varepsilon_{m}\to 0$ as $m\to\infty$ and a constant [^1] We choose $\eta>0$ for $p<2n$ and $q<\infty$, while $\eta=0$ for the other cases.
+$\eta=\eta(n,p,q)\geqslant 0$ such that for every $m\in\mathbb{N}$ and any initial datum $u_{0}\in\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ with
 ${\left\|u_{0}\right\|}_{\dot{B}_{p,q}^{n/p-1}}\leqslant\eta,\qquad\operatorname{div}u_{0}=0,$ (1.7)
@@
     Chemin [Che-92] and Planchon [Pla-98] established well-posedness for (1.1) in critical Besov spaces for all $1\leqslant p<\infty$ and $1\leqslant q\leqslant\infty$.
-This does not contradict our result, since their solution space is based on
-    $C([0,T);\dot{B}_{p,q}^{\frac{n}{p}-1}(\mathbb{R}^{n}))\cap\widetilde{L^{1}}(0,T;\dot{B}_{p,q}^{\frac{n}{p}+1}(\mathbb{R}^{n})).$ (1.9)
+This does not contradict our result, since their solution space is
+    $C([0,T);\dot{B}_{p,q}^{\frac{n}{p}-1}(\mathbb{R}^{n}))\cap\widetilde{L^{1}}(0,T;\dot{B}_{p,q}^{\frac{n}{p}+1}(\mathbb{R}^{n})),$ (1.9)
     whereas our non-unique solutions do not belong to the auxiliary dissipative space $\widetilde{L^{1}}(0,T;\dot{B}_{p,q}^{n/p+1}(\mathbb{R}^{n}))$.
@@
 Our result appears to be the first to establish non-uniqueness of steady solutions in scaling critical spaces.
-Moreover our result includes the three-dimensional case.
+Moreover, our result includes the three-dimensional case.
 Meanwhile, our method cannot be implemented in two dimensions.
@@
 The strategy is that we first establish a dominant part of the solution which lies in $\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ with $(p,q)$ satisfying (N1) or (N2), and then solve the remaining perturbative problem by a fixed point argument in a function space where well-posedness is available.
-More precisely, we construct an almost frequency-localized building blocks $\{V_{j}\}_{j\geqslant 0}$ of stationary flows, each of which are essentially built on the $j$th frequency shell, and design the principal profile
+More precisely, we construct almost frequency-localized building blocks $\{V_{j}\}_{j\geqslant 0}$ of stationary flows, each of which are essentially built on the $j$th frequency shell, and design the principal profile
 $V=\sum_{j=0}^{\infty}V_{j}\in\dot{B}_{p,q}^{\frac{n}{p}-1}(\mathbb{R}^{n})\setminus\{0\}$ (1.11)
@@
 $F:=-\Delta V+\mathbb{P}\operatorname{div}(V\otimes V)$ (1.12)
-belongs to well-posedness class $\dot{B}_{r,1}^{n/r-3}(\mathbb{R}^{n})$ ($r<n$) and admits the following smallness hierarchy:
+belongs to well-posedness [^4] See [Kan-Koz-Shi-19] for well-posedness of the stationary Navier–Stokes equations in critical Besov spaces.
+class $\dot{B}_{r,1}^{n/r-3}(\mathbb{R}^{n})$ ($r<n$) and admits the following smallness hierarchy:
 $\|F\|_{\dot{B}_{r,1}^{\frac{n}{r}-3}}\ll\|V\|_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\ll 1.$ (1.13)
@@
 Then, we may follow the standard stability argument to construct a small dissipative perturbed flow $w\in C([0,\infty);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))\cap\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{n/p-1+2/\theta}(\mathbb{R}^{n}))$ with some $2<\theta<\infty$.
-Then, the auxilliary dissipative space enables us to show $\|w(t)\|_{\dot{B}_{p,q}^{n/p-1}}\to 0$ as $t\to\infty$.
+Then, the auxiliary dissipative norm enables us to show $\|w(t)\|_{\dot{B}_{p,q}^{n/p-1}}\to 0$ as $t\to\infty$.
 We now explain how to construct the leading profile $V$ satisfying (1.13).
@@
 Instead, we distribute the $j$th building block over many frequency scales.
-To this end, for $\lambda\gg 1$ (which is a parameter corresponding to $\varepsilon^{-1/\alpha}$ with some $\alpha$) and a fixed large number $\mu\in\mathbb{N}$, we introduce a tetration-like sequence [^4] This is defined by $\lambda_{0}=\lambda$ and $\lambda_{j}=\lambda^{2^{\lambda_{j-1}^{4\mu}}}$; see (4.9).
+To this end, for $\lambda\gg 1$ (which is a parameter corresponding to $\varepsilon_{m}^{-1/\alpha}$ with some $\alpha$) and a fixed large number $\mu\in\mathbb{N}$, we introduce a tetration-like sequence [^6] This is defined by $\lambda_{0}=\lambda$ and $\lambda_{j}=\lambda^{2^{\lambda_{j-1}^{4\mu}}}$; see (4.9).
 $\{\lambda_{j}\}_{j\geqslant 0}$ and define the $j$th level frequency index set by
@@
 Here, the choice of $1/\sqrt{\ell}$ is inspired by [Yon-10, Tsu-19-ARMA]. With this multi-scale design, the oscillation calculus in Section 3 below yields
-${\left\|V_{j}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\sim\sum_{k\in\mathcal{K}}\frac{{\left\|f_{k,j}\right\|}_{L^{p}}}{\sqrt{h_{j}}}\left(\sum_{j\in\Lambda_{j}}\frac{\lambda^{(\frac{n}{p}-1)q\ell}}{\ell^{\frac{q}{2}}}\right)^{\frac{1}{q}},$ (1.17)
+${\left\|V_{j}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\sim\sum_{k\in\mathcal{K}}\frac{{\left\|f_{k,j}\right\|}_{L^{p}}}{\sqrt{h_{j}}}\left(\sum_{\ell\in\Lambda_{j}}\frac{\lambda^{(\frac{n}{p}-1)q\ell}}{\ell^{\frac{q}{2}}}\right)^{\frac{1}{q}},$ (1.17)
 which may shrink as $j\to\infty$ if and only if $(p,q)$ satisfies (N1) or (N2). Thus, the series (1.11) converges and we may define $V$. For the analysis of $F$, we decompose $F$ as
 $F=\sum_{j=1}^{\infty}\Bigl(-\Delta V_{j-1}+\mathbb{P}\operatorname{div}(V_{j}\otimes V_{j})\Bigr)+\sum_{j_{1}\neq j_{2}}\mathbb{P}\operatorname{div}(V_{j_{1}}\otimes V_{j_{2}})=:F_{1}+F_{2}.$ (1.18)
-The cross-shell interaction term $F_{2}$ is favorable, as it remains in high-frequency region.
+The cross-shell interaction term $F_{2}$ is favorable, as it remains in the high-frequency region.
 By contrast, $F_{1}$ exhibits a more delicate structure.
@@
 In Section 3, we provide key Besov estimates for oscillating functions, which play a key role in constructing the leading part of the non-unique stationary solutions.
-In Section 4, we construct the principal part of the non-trivial steady state and calculate Besov norms of it and the perturbative force.
-In Section 5, we make use of special flow constructed in Section 4 to complete the proof of Theorem 1.2.
+In Section 4, we construct the principal part of the non-trivial steady state and calculate its Besov norms and the perturbative force.
+In Section 5, we make use of the special flow constructed in Section 4 to complete the proof of Theorem 1.2.
 2. Preliminaries
@@
 where the principal part $F_{\lambda,L}^{({\rm p})}$ and the remainder part $F_{\lambda,L}^{({\rm r})}$ are defined by
-$F_{\lambda,L}^{({\rm p})}:=\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{|j-j_{\ell}|\leqslant 2}P_{k}\Delta_{j}g_{\lambda,k,\ell},$ (3.31)
+$F_{\lambda,L}^{({\rm p})}:=\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{|j-j_{\ell}|\leqslant 1}P_{k}\Delta_{j}g_{\lambda,k,\ell},$ (3.31)
 $\begin{aligned} F_{\lambda,L}^{({\rm r})}:={}&\sum_{m=0}^{K}\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{(m,j)\in R_{k,\ell}}P_{m}\Delta_{j}g_{\lambda,k,\ell}.\end{aligned}$ (3.32)
@@
 $=b_{\ell}e^{i\lambda^{\ell}a_{m}\cdot x}f_{m}(x)\int_{\mathbb{R}^{n}}\chi_{m,j_{\ell}}(y)e^{-i\lambda^{\ell}a_{m}\cdot y}dy$ (3.59)
-$\qquad+b_{\ell}e^{i\lambda^{\ell}a_{k}\cdot x}\int_{\mathbb{R}^{n}}(f_{m}(x-y)-f_{m}(x))\chi_{m,j_{\ell}}(y)e^{-i\lambda^{\ell}a_{m}\cdot y}dy$ (3.60)
+$\qquad+b_{\ell}e^{i\lambda^{\ell}a_{m}\cdot x}\int_{\mathbb{R}^{n}}(f_{m}(x-y)-f_{m}(x))\chi_{m,j_{\ell}}(y)e^{-i\lambda^{\ell}a_{m}\cdot y}dy$ (3.60)
 $=:b_{\ell}e^{i\lambda^{\ell}a_{m}\cdot x}f_{m}(x)+R_{\lambda,m,\ell}(x),$ (3.61)
@@
 By (3.57) and (3.65), it holds
-${\left\|F_{\lambda,m,L}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{s}}\geqslant{}$ $c\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}{\left\|f_{m}\right\|}_{L^{p}}-C\sum_{\ell=1}^{L}\lambda^{(s-1)\ell}|b_{\ell}|{\left\|\nabla f_{k}\right\|}_{L^{p}}.$ (3.66)
+${\left\|F_{\lambda,m,L}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{s}}\geqslant{}$ $c\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}{\left\|f_{m}\right\|}_{L^{p}}-C\sum_{\ell=1}^{L}\lambda^{(s-1)\ell}|b_{\ell}|{\left\|\nabla f_{m}\right\|}_{L^{p}}.$ (3.66)
 On the estimate of $F^{({\rm r})}_{\lambda,m,L}$, the similar strategy as for the estimate of $F^{({\rm r})}_{\lambda,m,L}$ yields
@@
     Let $\{a_{k}\}_{k\in\mathcal{K}}\subset\mathbb{S}^{n-1}$ satisfy that $a_{k}$ and $a_{k^{\prime}}$ are not parallel unless $k=k^{\prime}$. For instance, we may choose
-    $a_{k}:=\begin{cases}\mathbbm{e}_{i+1},&(k=\mathbbm{e}_{i},\ i=1,\dots,n-1),\\ \mathbbm{e}_{1},&(k=\mathbbm{e}_{n}),\\ \sqrt{2}\left(\mathbbm{e}_{i}\mp\mathbbm{e}_{j}\right),&(k=\mathbbm{e}_{i}\pm\mathbbm{e}_{j},\ i,j=1,\dots,n,\ i<j).\end{cases}$ (4.8)
+    $a_{k}:=\begin{cases}\mathbbm{e}_{i+1},&(k=\mathbbm{e}_{i},\ i=1,\dots,n-1),\\ \mathbbm{e}_{1},&(k=\mathbbm{e}_{n}),\\ \dfrac{\mathbbm{e}_{i}\mp\mathbbm{e}_{j}}{\sqrt{2}},&(k=\mathbbm{e}_{i}\pm\mathbbm{e}_{j},\ i,j=1,\dots,n,\ i<j).\end{cases}$ (4.8)
   2. 2.
@@
     where we have used the notation $\exp_{a}(b)\coloneqq a^{b}$ for $a>0$ and $b\in\mathbb{R}$.
-    We note that, throughout this section, $\lambda$ may be replaced by a larger dyadic number whenever necessary, without further comment.
+    Throughout this section, $\lambda$ may be replaced by a larger dyadic number whenever necessary, without further comment.
 We emphasize that the required lower bound on $\lambda$ is chosen independently of $j$.
 Moreover, almost all quantities introduced below depend on the parameter $\lambda$; however, we do not record this dependence explicitly in their notations so as to avoid unnecessarily complicating them.
+We also note that the generic constants are always independent of $\lambda$ and $j$.
   3. 3.
@@
 From the Young inequality, we see that
-${\left\|\nabla^{m}{\mathcal{V}}_{j}[v]\right\|}_{L^{p}}$ $\leqslant C\lambda_{j-1}^{4}{\left\|\nabla^{m+2}\rho_{k}\right\|}_{L^{1}}\sum_{k\in\mathcal{K}}{\left\|\Gamma_{j,k}[v]\Psi_{j,k}\right\|}_{L^{p}}$ (4.24)
+${\left\|\nabla^{m}{\mathcal{V}}_{j}[v]\right\|}_{L^{p}}$ $\leqslant C\lambda_{j-1}^{4}{\left\|\nabla^{m+2}\rho_{j}\right\|}_{L^{1}}\sum_{k\in\mathcal{K}}{\left\|\Gamma_{j,k}[v]\Psi_{j,k}\right\|}_{L^{p}}$ (4.24)
 $\leqslant C\lambda_{j-1}^{4}\lambda_{j}^{2m+4}\sum_{k\in\mathcal{K}}{\left\|\Psi_{j,k}\right\|}_{L^{p}}.$ (4.25)
 Since $\Psi_{j,k}$ is bounded as
-${\left\|\Psi_{j,k}\right\|}_{L^{p}}\leqslant\sum_{\ell\in\Lambda_{j}}\frac{C\lambda^{\frac{n}{p}j}}{\lambda^{2\ell}\sqrt{h_{j}\ell}}\leqslant\frac{C\lambda^{\frac{n}{p}j}}{\lambda^{4\mu}_{j-1}\sqrt{\mu\log_{\lambda}\lambda_{j-1}}}\leqslant C\lambda^{-4\mu}_{j-1}$ (4.26)
+${\left\|\Psi_{j,k}\right\|}_{L^{p}}\leqslant\sum_{\ell\in\Lambda_{j}}\frac{C\lambda^{\frac{n}{p}j}}{\lambda^{2\ell}\sqrt{h_{j}\ell}}\leqslant\frac{C\lambda^{\frac{n}{p}j}}{\lambda^{4\mu}_{j-1}\sqrt{\mu\log_{\lambda}\lambda_{j-1}}}\leqslant C\lambda^{n-4\mu}_{j-1}$ (4.26)
 we obtain (4.22). For the estimate of (4.23), (4.22) with $p=\infty$ and $m=1$ yields
-${\left\|\mathcal{D}\mathcal{V}[v]\right\|}_{L^{\infty}}\leqslant C{\left\|\nabla\mathcal{V}[v]\right\|}_{L^{\infty}}\leqslant C\lambda_{j}^{6}\leqslant r_{0}\lambda_{j}^{8}$ (4.27)
+${\left\|\mathcal{D}\mathcal{V}_{j}[v]\right\|}_{L^{\infty}}\leqslant C{\left\|\nabla\mathcal{V}_{j}[v]\right\|}_{L^{\infty}}\leqslant C\lambda_{j}^{6}\leqslant r_{0}\lambda_{j}^{8}$ (4.27)
 provided that $\lambda$ is sufficiently large. Thus, we complete the proof. ∎
@@
 ${\left\|V_{j}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}$ $\leqslant C\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-2\mu}\sigma_{j}\left(\frac{n}{p}-1,q\right)$ (4.72)
-$\leqslant C\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-2\mu}\frac{\lambda_{j-1}^{(\frac{n}{p}-1)\mu}}{\sqrt{\mu\log_{\lambda}\lambda_{j-1}}}\leqslant C\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-(3-\frac{n}{p})\mu}.$ (4.73)
+$\leqslant\begin{cases}C\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-2\mu}\left(\mu\log_{\lambda}\lambda_{j-1}\right)^{\frac{1}{q}-\frac{1}{2}}&(p=n,\ q>2),\\[5.0pt] C\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-2\mu}\lambda_{j-1}^{(\frac{n}{p}-1)\mu}\left(\mu\log_{\lambda}\lambda_{j-1}\right)^{-\frac{1}{2}}&(p>n)\end{cases}$ (4.73)
+$\leqslant C\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-(3-\frac{n}{p})\mu}.$ (4.74)
 Hence, the series
@@
 which is implied by Lemma 4.3 and $\psi_{j}(x)=1$ for $x\in\operatorname{supp}V_{j-1}$, we have
-$V^{({\rm p})}_{j}\otimes V^{({\rm p})}_{j}={}2\lambda_{j-1}^{8}\sum_{k_{1},k_{2}\in\mathcal{K}}\sum_{\ell_{1},\ell_{2}\in\Lambda_{j}}\Phi_{j,k_{1},\ell_{1}}\otimes\Phi_{j,k_{2},\ell_{2}}$ (4.99)
+$V^{({\rm p})}_{j}(x)\otimes V^{({\rm p})}_{j}(x)={}2\lambda_{j-1}^{8}\sum_{k_{1},k_{2}\in\mathcal{K}}\sum_{\ell_{1},\ell_{2}\in\Lambda_{j}}\Phi_{j,k_{1},\ell_{1}}(x)\otimes\Phi_{j,k_{2},\ell_{2}}(x)$ (4.100)
 $\quad={}\lambda_{j-1}^{8}\psi_{j}^{2}\sum_{k\in\mathcal{K}}\Gamma_{k}\left({\rm Id}-\lambda_{j-1}^{-8}\mathcal{D}V_{j-1}(x)\right)^{2}k\otimes k$ (4.101)
 $\qquad+\frac{\lambda_{j-1}^{8}}{h_{j}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\Lambda_{j}}\frac{1}{\ell}\Gamma_{k}\left({\rm Id}-\lambda_{j-1}^{-8}\mathcal{D}V_{j-1}(x)\right)^{2}\psi_{j}^{2}(x)\cos\left(2\lambda^{\ell}a_{k}\cdot x\right)k\otimes k$ (4.102)
-$\qquad+2\lambda_{j-1}^{8}\sum_{\begin{subarray}{c}k_{1},k_{2}\in\mathcal{K}\\ k_{1}\neq k_{2}\end{subarray}}\sum_{\ell_{1},\ell_{2}\in\Lambda_{j}}\Phi_{j,k_{1},\ell}(x)\Phi_{j,k_{2},\ell}(x)$ (4.102)
-$\quad={}\lambda_{j-1}^{8}\psi_{j}^{2}{\rm Id}-\mathcal{D}V_{j-1}+F_{1,1;j}+F_{1,2;j},$ (4.103)
+$\qquad+2\lambda_{j-1}^{8}\sum_{\begin{subarray}{c}k_{1},k_{2}\in\mathcal{K}\\ k_{1}\neq k_{2}\end{subarray}}\sum_{\ell_{1},\ell_{2}\in\Lambda_{j}}\Phi_{j,k_{1},\ell}(x)\otimes\Phi_{j,k_{2},\ell}(x)$ (4.103)
+$\qquad+2\lambda_{j-1}^{8}\sum_{k\in\mathcal{K}}\sum_{\begin{subarray}{c}\ell_{1},\ell_{2}\in\Lambda_{j}\\ \ell_{1}\neq\ell_{2}\end{subarray}}\Phi_{j,k,\ell_{1}}(x)\otimes\Phi_{j,k,\ell_{2}}(x)$ (4.104)
+$\quad={}\lambda_{j-1}^{8}\psi_{j}(x)^{2}{\rm Id}-\mathcal{D}V_{j-1}(x)+F_{1,1;j}(x)+F_{1,2;j}(x)+F_{1,3;j}(x),$ (4.105)
 where we have set
-$F_{1,1;j}:=\frac{\lambda_{j-1}^{8}}{h_{j}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\Lambda_{j}}\frac{1}{\ell}\Gamma_{j,k}\left[V_{j-1}\right]^{2}\psi_{j}^{2}\cos\left(2\lambda^{\ell}a_{k}\cdot x\right)k\otimes k,$ (4.104)
-$F_{1,2;j}:=2\lambda_{j-1}^{8}\sum_{\begin{subarray}{c}k_{1},k_{2}\in\mathcal{K}\\ k_{1}\neq k_{2}\end{subarray}}\sum_{\ell_{1},\ell_{2}\in\Lambda_{j}}\Phi_{j,k_{1},\ell_{1}}(x)\otimes\Phi_{j,k_{2},\ell_{2}}(x).$ (4.105)
+$F_{1,1;j}(x):=\frac{\lambda_{j-1}^{8}}{h_{j}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\Lambda_{j}}\frac{1}{\ell}\Gamma_{j,k}\left[V_{j-1}\right](x)^{2}\psi_{j}^{2}\cos\left(2\lambda^{\ell}a_{k}\cdot x\right)k\otimes k,$ (4.106)
+$F_{1,2;j}(x):=2\lambda_{j-1}^{8}\sum_{\begin{subarray}{c}k_{1},k_{2}\in\mathcal{K}\\ k_{1}\neq k_{2}\end{subarray}}\sum_{\ell_{1},\ell_{2}\in\Lambda_{j}}\Phi_{j,k_{1},\ell_{1}}(x)\otimes\Phi_{j,k_{2},\ell_{2}}(x),$ (4.107)
+$F_{1,3;j}(x):=2\lambda_{j-1}^{8}\sum_{k\in\mathcal{K}}\sum_{\begin{subarray}{c}\ell_{1},\ell_{2}\in\Lambda_{j}\\ \ell_{1}\neq\ell_{2}\end{subarray}}\Phi_{j,k,\ell_{1}}(x)\otimes\Phi_{j,k,\ell_{2}}(x).$ (4.108)
 Then, from the cancellation property
@@
 it follows that
-$F_{1}$ $=\sum_{j=1}^{\infty}\left(-\Delta V_{j-1}+\mathbb{P}\operatorname{div}\left(\lambda_{j-1}^{8}\psi_{j}^{2}{\rm Id}-\mathcal{D}V_{j-1}+F_{1,1;j}+F_{1,2;j}\right)\right)$ (4.107)
-$=\sum_{j=1}^{\infty}\mathbb{P}\operatorname{div}F_{1,1;j}+\sum_{j=1}^{\infty}\mathbb{P}\operatorname{div}F_{1,2;j}$ (4.108)
-$=:F_{1,1}+F_{1,2}.$ (4.109)
+$F_{1}$ $=\sum_{j=1}^{\infty}\left(-\Delta V_{j-1}+\mathbb{P}\operatorname{div}\left(\lambda_{j-1}^{8}\psi_{j}^{2}{\rm Id}-\mathcal{D}V_{j-1}+F_{1,1;j}+F_{1,2;j}+F_{1,3;j}\right)\right)$ (4.110)
+$=\sum_{j=1}^{\infty}\mathbb{P}\operatorname{div}F_{1,1;j}+\sum_{j=1}^{\infty}\mathbb{P}\operatorname{div}F_{1,2;j}+\sum_{j=1}^{\infty}\mathbb{P}\operatorname{div}F_{1,3;j}$ (4.111)
+$=:F_{1,1}+F_{1,2}+F_{1,3}.$ (4.112)
 To proceed with the estimate of $F_{1,1}$, we begin with the estimate of each element $F_{1,1;j}$. It follows from Proposition 3.1 with $M=n$ that
@@
 we see that
-${\left\|F_{1,2;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}\leqslant{}$ $C\lambda_{j-1}^{8-4\mu}\sum_{\pm}\sum_{\begin{subarray}{c}\ell_{1}\in\Lambda_{j}\end{subarray}}\sum_{\ell_{2}:\ell_{2}\geqslant\ell_{1}}\sum_{\begin{subarray}{c}k_{1}\neq k_{2}\end{subarray}}$ (4.116)
+${\left\|F_{1,2;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}\leqslant{}$ $C\lambda_{j-1}^{8-4\mu}\sum_{\pm}\sum_{\begin{subarray}{c}\ell_{1}\in\Lambda_{j}\end{subarray}}\sum_{\ell_{2}=\ell_{1}}^{\infty}\sum_{\begin{subarray}{c}k_{1}\neq k_{2}\end{subarray}}$ (4.119)
 $\left(\frac{\lambda^{(\frac{n}{r}-2)\ell_{2}}}{\sqrt{\ell_{1}\ell_{2}}}|\lambda^{\ell_{1}-\ell_{2}}a_{k_{1}}\pm a_{k_{2}}|^{\frac{n}{r}-2}{\left\|\Gamma_{j,k_{1}}[V_{j-1}]\Gamma_{j,k_{2}}[V_{j-1}]\psi_{j}^{2}\right\|}_{L^{r}}\right$.
@@
 $\geqslant{}$ $\sin^{2}\vartheta(a_{k_{1}},a_{k_{2}})>0.$ (4.125)
-Here, one may recall (3.13) for the definition of $\vartheta(\cdot,\cdot)$. Hence, we obtain
-${\left\|F_{1}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-3}}$ $\leqslant C\sum_{j=1}^{\infty}\left({\left\|F_{1,1;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}+{\left\|F_{1,2;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}\right)$ (4.123)
+Here, one may recall (3.13) for the definition of $\vartheta(\cdot,\cdot)$. We may treat $F_{1,3;j}$ similarly. Indeed, as there holds
+$F_{1,3;j}(x)=\frac{\lambda_{j-1}^{8}}{h_{j}}\sum_{\pm}\sum_{\ell_{1}\neq\ell_{2}}\sum_{k\in\mathcal{K}}$ $\frac{1}{\sqrt{\ell_{1}\ell_{2}}}\Gamma_{j,k}[V_{j-1}](x)^{2}$ (4.126)
+$\psi_{j}(x)^{2}\cos\left((\lambda^{\ell_{1}}\pm\lambda^{\ell_{2}})a_{k}\cdot x\right)k\otimes k,$ (4.127)
+we have
+${\left\|F_{1,3;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}\leqslant{}$ $C\lambda_{j-1}^{8-4\mu}\sum_{\pm}\sum_{k\in\mathcal{K}}\sum_{\begin{subarray}{c}\ell_{1}\in\Lambda_{j}\end{subarray}}\sum_{\ell_{2}=\ell_{1}+1}^{\infty}$ (4.128)
+$\left(\frac{\lambda^{(\frac{n}{r}-2)\ell_{2}}}{\sqrt{\ell_{1}\ell_{2}}}|(1\pm\lambda^{\ell_{1}-\ell_{2}})a_{k}|^{\frac{n}{r}-2}{\left\|\Gamma_{j,k}[V_{j-1}]^{2}\psi_{j}^{2}\right\|}_{L^{r}}\right$.
+(4.129)
+$\left.\quad+\frac{\lambda^{-2\ell_{2}}}{\sqrt{\ell_{1}\ell_{2}}}|(1\pm\lambda^{\ell_{1}-\ell_{2}})a_{k}|^{-2}{\left\|\nabla^{n}(\Gamma_{j,k}[V_{j-1}]^{2}\psi_{j}^{2})\right\|}_{L^{r}\cap L^{1}}\right)$ (4.130)
+$\leqslant{}$ $C\lambda_{j-1}^{8-4\mu}\cdot C\lambda_{j-1}^{(\frac{n}{r}-2)\mu}\cdot C\lambda^{\frac{n}{r}}+C\lambda_{j-1}^{8-4\mu}\cdot C\lambda_{j-1}^{-2\mu}\cdot C\lambda^{n}\lambda_{j-1}^{2n}$ (4.131)
+$\leqslant{}$ $C\lambda_{j-1}^{4n+8-4\mu}.$ (4.132)
+Hence, we obtain
+${\left\|F_{1}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-3}}$ $\leqslant C\sum_{j=1}^{\infty}\left({\left\|F_{1,1;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}+{\left\|F_{1,2;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}+{\left\|F_{1,3;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}\right)$ (4.133)
 $\leqslant C\sum_{j=1}^{\infty}\lambda_{j-1}^{4n+8-4\mu}$ (4.134)
@@
 $\leqslant{}$ $C\lambda^{\frac{n}{r}j_{1}}\lambda_{j_{1}-1}^{4-(4-\frac{n}{r})\mu}\lambda_{j_{2}-1}^{4-2\mu}+C\lambda^{nj_{1}}\lambda_{j_{1}-1}^{4-4\mu}\lambda_{j_{2}-1}^{4-2\mu+2n}$ (4.147)
-$\leqslant{}$ $C\lambda^{nj_{1}}\lambda_{j_{1}-1}^{4-(4-\frac{n}{r})\mu}\lambda_{j_{2}-1}^{4-2\mu}.$ (4.138)
+$\leqslant{}$ $C\lambda^{nj_{1}}\lambda_{j_{1}-1}^{4-(4-\frac{n}{r})\mu}\lambda_{j_{2}-1}^{4-2\mu+2n}.$ (4.148)
 Hence, we obtain
 ${\left\|F_{2}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-3}}$ $\leqslant C\sum_{j_{1}=1}^{\infty}\sum_{j_{2}=j_{1}+1}^{\infty}{\left\|F_{2;j_{1},j_{2}}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}$ (4.149)
-$\leqslant C\sum_{j_{1}=1}^{\infty}\lambda^{nj_{1}}\lambda_{j_{1}-1}^{4-(4-\frac{n}{r})\mu}\sum_{j_{2}=j_{1}+1}^{\infty}\lambda_{j_{2}-1}^{4-2\mu}$ (4.140)
-$\leqslant C\sum_{j_{1}=1}^{\infty}\lambda^{nj_{1}}\lambda_{j_{1}-1}^{8-(6-\frac{n}{r})\mu}$ (4.141)
-$\leqslant C\lambda^{8+n-(6-\frac{n}{r})\mu}\leqslant C\lambda^{4n+8-4\mu}=C\lambda^{-\beta}.$ (4.142)
+$\leqslant C\sum_{j_{1}=1}^{\infty}\lambda^{nj_{1}}\lambda_{j_{1}-1}^{4-(4-\frac{n}{r})\mu}\sum_{j_{2}=j_{1}+1}^{\infty}\lambda_{j_{2}-1}^{4-2\mu+2n}$ (4.150)
+$\leqslant C\sum_{j_{1}=1}^{\infty}\lambda^{nj_{1}}\lambda_{j_{1}-1}^{8+2n-(6-\frac{n}{r})\mu}$ (4.151)
+$\leqslant C\lambda^{8+3n-(6-\frac{n}{r})\mu}\leqslant C\lambda^{4n+8-4\mu}=C\lambda^{-\beta}.$ (4.152)
 Finally, we focus on the estimate of $F_{3}$. Let $n<p_{*}<2n$ satisfy (2.7) with $p=p_{*}$. It then follows from Lemma 2.1 that
@@
 $\leqslant C\left(\lambda^{\frac{n}{p_{*}}+4-(3-\frac{n}{p_{*}})\mu}+\lambda^{n+6-(4-\frac{n}{r})\mu}\right)\lambda^{n+6-(4-\frac{n}{r})\mu}$ (4.154)
-$\leqslant C\left(\lambda^{\frac{9}{2}-\frac{5}{2}\mu}+\lambda^{n+6-2\mu}\right)\lambda^{n+6-2\mu}$ (4.145)
+$\leqslant C\left(\lambda^{5-2\mu}+\lambda^{n+6-2\mu}\right)\lambda^{n+6-2\mu}$ (4.155)
 $\leqslant C\lambda^{2n+12-4\mu}<C\lambda^{4n+8-4\mu}=C\lambda^{-\beta}.$ (4.156)
@@
 To this end, we consider the stationary perturbed system
-$\begin{cases}-\Delta W_{\lambda}+\mathbb{P}\operatorname{div}\left(V_{\lambda}\otimes W_{\lambda}+W_{\lambda}\otimes V_{\lambda}+W_{\lambda}\otimes W_{\lambda}\right)=-F,\quad&x\in\mathbb{R}^{n},\\ \operatorname{div}W_{\lambda}=0,\quad&x\in\mathbb{R}^{n},\end{cases}$ (5.1)
+$\begin{cases}-\Delta W_{\lambda}+\mathbb{P}\operatorname{div}\left(V_{\lambda}\otimes W_{\lambda}+W_{\lambda}\otimes V_{\lambda}+W_{\lambda}\otimes W_{\lambda}\right)=-F_{\lambda},\quad&x\in\mathbb{R}^{n},\\ \operatorname{div}W_{\lambda}=0,\quad&x\in\mathbb{R}^{n},\end{cases}$ (5.1)
 and define an operator $\mathcal{X}$ on $\dot{B}_{r,1}^{n/r-1}(\mathbb{R}^{n})$ by
@@
 Then, choosing $\lambda$ so large that $2C_{1}\left(\lambda^{-\alpha(p_{*})}+4C_{1}\lambda^{-\beta}\right)\leqslant 1$, we have
-${\left\|\mathcal{X}[W]\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\leqslant 2C_{1}\lambda^{-\beta},\qquad{\left\|\mathcal{X}[W_{1}]-\mathcal{X}[W_{2}]\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\leqslant\frac{1}{2}{\left\|W_{1}-W_{2}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}$ (5.8)
+${\left\|\mathcal{X}[W]\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\leqslant\frac{3}{2}C_{1}\lambda^{-\beta},\qquad{\left\|\mathcal{X}[W_{1}]-\mathcal{X}[W_{2}]\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\leqslant\frac{1}{2}{\left\|W_{1}-W_{2}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}$ (5.8)
 for all $W,W_{1},W_{2}\in\mathscr{X}$. Hence, the contraction mapping principle yields the existence of a unique solution $W_{\lambda}\in\mathscr{X}$ to (5.1).
@@
 Let us consider the perturbed system of (1.1) around $U_{\lambda}$.
-$\begin{cases}\partial_{t}w_{\lambda}-\Delta w_{\lambda}+\mathbb{P}\operatorname{div}(U_{\lambda}\otimes w_{\lambda}+U_{\lambda}\otimes w_{\lambda}+w_{\lambda}\otimes w_{\lambda})=0,\\ \operatorname{div}w_{\lambda}=0,\\ w_{\lambda}(0,x)=w_{0}:=u_{0}-U_{\lambda}.\end{cases}$ (5.11)
+$\begin{cases}\partial_{t}w_{\lambda}-\Delta w_{\lambda}+\mathbb{P}\operatorname{div}(U_{\lambda}\otimes w_{\lambda}+w_{\lambda}\otimes U_{\lambda}+w_{\lambda}\otimes w_{\lambda})=0,\\ \operatorname{div}w_{\lambda}=0,\\ w_{\lambda}(0,x)=w_{0}:=u_{0}-U_{\lambda}.\end{cases}$ (5.11)
 To construct the perturbed solution, we define a mapping
@@
 $\quad\leqslant{}C_{2}\left(\lambda^{-\alpha(p)}+\sum_{m=1}^{2}{\left\|w_{m}\right\|}_{\widetilde{L^{\infty}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1})}\right){\left\|w_{1}-w_{2}\right\|}_{\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}$ (5.18)
-for all $w,w_{1},w_{2}\in\widetilde{C}([0,\infty);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))\cap\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{n/p-1+2/\theta}(\mathbb{R}^{n}))$ with some positive constant $C_{2}=C_{2}(n,p,r,\theta)$.
+for all $w,w_{1},w_{2}\in\widetilde{C}([0,\infty);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))\cap\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{n/p-1+2/\theta}(\mathbb{R}^{n}))$ with some positive constant $C_{2}=C_{2}(n,p,q,\theta)$.
 We define the solution space $\mathscr{Y}$ by
 $\mathscr{Y}:=\left\{\begin{aligned} w\in{}&\widetilde{C}([0,\infty);\dot{B}_{p,q}^{\frac{n}{p}-1}(\mathbb{R}^{n}))\cap\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}}(\mathbb{R}^{n}))\ ;\\ &{\left\|w\right\|}_{\widetilde{L^{\infty}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1})\cap\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}\leqslant 2C_{2}{\left\|w_{0}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\end{aligned}\right\}.$ (5.19)
-Then, we see that $\mathcal{Y}$ is a contraction mapping on $\mathscr{Y}$, by choosing $\lambda$ and ${\left\|w_{0}\right\|}_{\dot{B}_{p,q}^{n/p-1}}$ sufficiently large and small, respectively.
+Choosing $\lambda$ so large and ${\left\|w_{0}\right\|}_{\dot{B}_{p,q}^{n/p-1}}$ so small that
+$2C_{2}\left(\lambda^{-\alpha(p)}+4C_{2}{\left\|w_{0}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\right)\leqslant 1,$ (5.20)
+we have
+${\left\|\mathcal{Y}[w]\right\|}_{\widetilde{L^{\infty}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1})\cap\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}\leqslant\frac{3}{2}C_{2}{\left\|w_{0}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}},$ (5.21)
+${\left\|\mathcal{Y}[w_{1}]-\mathcal{Y}[w_{2}]\right\|}_{\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}\leqslant\frac{1}{2}{\left\|w_{1}-w_{2}\right\|}_{\widetilde{L^{\theta}}(0,\infty;\dot{B}_{p,q}^{\frac{n}{p}-1+\frac{2}{\theta}})}$ (5.22)
+for all $w,w_{1},w_{2}\in\mathscr{Y}$, which implies that $\mathcal{Y}$ is a contraction mapping on $\mathscr{Y}$.
 Hence, the Banach fixed point theorem implies that there exists a unique mild solution $w_{\lambda}\in\mathscr{Y}$ to the perturbed system (5.11).